### v1

# Predictively Oriented Posteriors

Abstract
We advocate for a new statistical principle that combines the most desirable aspects of both parameter inference and density estimation. This leads us to the predictively oriented (PrO) posterior, which expresses uncertainty as a consequence of predictive ability. Doing so leads to inferences which predictively dominate both classical and generalised Bayes posterior predictive distributions: up to logarithmic factors, PrO posteriors converge to the predictively optimal model average at rate $n^{-1/2}$. Whereas classical and generalised Bayes posteriors only achieve this rate if the model can recover the data-generating process, PrO posteriors adapt to the level of model misspecification. This means that they concentrate around the true model at rate $n^{1/2}$ in the same way as Bayes and Gibbs posteriors if the model can recover the data-generating distribution, but do not concentrate in the presence of non-trivial forms of model misspecification. Instead, they stabilise towards a predictively optimal posterior whose degree of irreducible uncertainty admits an interpretation as the degree of model misspecification—a sharp contrast to how Bayesian uncertainty and its existing extensions behave. Lastly, we show that PrO posteriors can be sampled from by evolving particles based on mean field Langevin dynamics, and verify the practical significance of our theoretical developments on a number of numerical examples.
[…]
Often motivated through its favourable decision-theoretic properties, the primary goal of Bayesian statistical inference is to quantify uncertainty about a parameter of interest given a set of prior beliefs (e.g. Robert et al., 2007). All other inferential tasks are derivatives of this primary directive. In the current era of black-box prediction models and the wide-spread adoption of machine learning techniques across science and society, this principle leads to a fundamental tension with Bayesian statistical methods: in many modern problems, prediction is the primary goal of interest, with parameter inference being only a secondary concern. Because much of Bayesian methodology is centered on parameter inference, its predictive performance quickly falters outside of the narrow conditions that make Bayesian belief updates optimal (see Zellner, 1988; Aitchison, 1975). In other words, the practice of Bayesian statistics is overwhelmingly focused on parameter inference, not prediction.
One of the clearest ways in which this manifests is posterior collapse: as data accumulate, the Bayes posterior generally concentrates around a single parameter value (see e.g van der Vaart, 1998, Section 10.2). This is sensible for parameter inference when
uncertainty in a parameter of interest arises due to missing observations. If a complete population is observed, then the parameter of interest can be assumed to be known precisely.
(Fong et al., 2023)
In the context of prediction however, posterior collapse is generally not desirable: it implies that as the sample size increases, marginalising over the Bayes posterior ceases to function as a genuine combination of individual predictive distributions. Instead of averaging over a model class, such predictive distributions essentially predict from a single model. Crucially, this behaviour persists even for generalised Bayesian methods that replace a likelihood-based approach with a loss-based belief update (Bissiri et al., 2016; Jiang and Tanner, 2008; Zhang, 2006; Knoblauch et al., 2022; Alquier, 2024; Syring and Martin, 2019; Martin and Syring, 2022) As a result, the posterior predictives obtained from both classical and generalised Bayes approaches quickly resemble those obtained from a point estimator as sample size grows (McLatchie et al., 2025).
To better understand the impact of this phenomenon on prediction, it is useful to distinguish between two regimes: when a statistical model is well-specified so that the data-generating process is recoverable by a singular parameter value, posterior concentration is the desired outcome: as more data accumulates, uncertainty contracts around the true parameter, and the posterior predictive distribution is well-calibrated. In practice however, statistical analysis is almost always subject to some degree of model misspecification. In this regime, no single parameter value is capable of describing the data generating process, so that the posterior predictives based on standard or generalised Bayesian posteriors become poorly calibrated (Syring and Martin, 2019). Importantly, this issue is not caused by obtaining predictive distributions through the marginalisation of a posterior, but by marginalising over a posterior which collapses to a point mass. This in turn provides strong motivation for methods whose parameter uncertainty can be guaranteed not to collapse to a point mass when this
is appropriate and desirable, since […] some uncertainty always remains about the true (idealised) distribution, no matter how much data are observed.
(Miller and Dunson, 2019)
In this paper, we synthesise the above discussion into a new statistical principle for inference that results in a predictively oriented (PrO) posterior. The resulting approach is distinct from existing inference frameworks in two main ways: first, it treats prediction as its main goal without requiring a black box approach, while naturally inducing uncertainty quantification for a parameter of interest as a secondary inferential outcome. Second, it is adaptive: the PrO posterior collapses to a point mass only in the appropriate setting where the model is well-specified. When the model is misspecified however, PrO posteriors acknowledge that no single parameter is sufficient for capturing the true data-generating process, and yield an irreducible level of uncertainty, even with increasing sample size.
Conceptually, PrO posteriors represent a useful intermediary between existing statistical paradigms and compromise between the two tenants of data modelling and algorithmic modelling sub-cultures in statistics that were famously identified by Breiman (2001). In practical terms, this means that PrO posteriors reconcile parameter inference and predictive inference in situations where the choice between either is difficult. Intriguingly, they do so through a posterior that expresses epistemological uncertainty in a way that is meaningfully distinct from standard and generalised Bayesian approaches.
We strengthen this intuitive appeal by demonstrating that PrO posteriors possess strong theoretical guarantees and an algorithmically viable approach for computation. In particular, we formally establish that PrO posteriors predictively dominate classical and generalised Bayesian posteriors. Further, we characterise the type of parameter uncertainty they confer, and show that PrO posteriors are adaptive to model misspecification. Lastly, we introduce a sampling algorithm which can be used to sample from any such posterior, and illustrate its performance across a range of examples.
In the remainder, we motivate the need for PrO posteriors in more detail (Section 1), develop new mathematical tools for our theoretical analysis (Section 2), and then present our main findings (Section 3). A full discussion of all results and the assumptions we use is deferred to the end of the paper (Section 7). We then introduce computationally feasible algorithms for PrO posteriors (Section 4) and by illustrating their theoretical performance on a number of numerical examples (Section 5) before concluding with a detailed discussion of our findings (Section 6).
1 Motivation
Throughout, we consider data ${x}_{1:n}=(x_{1},\dots,x_{n})$ for $x_{i}\in\mathcal{X}$ and $i=1,\dots,n$. We assume that each $x_{i}$ was generated independently from an unknown distribution $P_{0}$, and denote expectations of a function $f:\mathcal{X}\to\mathbb{R}$ with respect to $P_{0}$ by $\mathbb{E}[f(X)]$ or by $\mathbb{E}_{X\sim P_{0}}[f(X)]$ if we wish to emphasise the sampling distribution. To learn about $P_{0}$ from $x_{1:n}$, we follow the tenants of classical Bayesian analysis, and posit a model class $\mathcal{M}_{\Theta}:=\{P_{\theta}:\theta\in\Theta\}$ and a prior $\Pi\in\mathcal{P}(\Theta)$ over elements of $\mathcal{M}_{\Theta}$, where $\mathcal{P}(\Theta)$ denotes the space of probability measures on $\Theta$.
Instead of attempting to learn the best possible parameter, as in classical (or even generalised) Bayesian inference, we focus on learning the best possible predictive distribution. To rigorously formulate this optimality, we use scoring rules.
Definition 1 (Scoring Rule).
Let $S:\mathcal{P}(\mathcal{X})\times\mathcal{X}\to\mathbb{R}\cup\{\infty\}$ be an extended real-valued function. Then $S$ is a proper scoring rule, relative to a class $\mathcal{P}$, if:
  • (i)
    the expected score $\mathcal{S}(P,P^{\prime}):=\mathbb{E}_{X\sim P^{\prime}}\left\{S(P,X)\right\}$, is finite for some $P,P^{\prime}\in\mathcal{P}$;
  • (ii)
    $\mathcal{S}(P^{\prime},P^{\prime})\leq\mathcal{S}(P,P^{\prime})$ for all $P,P^{\prime}\in\mathcal{P}$, so that $\mathcal{S}$ induces the statistical divergence $\mathcal{D}_{S}(P,P^{\prime}):=\mathcal{S}(P,P^{\prime})-\mathcal{S}(P^{\prime},P^{\prime})\geq 0$ for which $\mathcal{D}_{S}(P,P^{\prime})=0\Longleftrightarrow P=P^{\prime}$.
Scoring rules are foundational statistical objects concerned with measuring accuracy of predictive distributions (see, e.g., Gneiting and Raftery, 2007). Given a target distribution $P_{0}$, a candidate predictive distribution $P$ is judged to be more accurate than an alternative candidate $P^{\prime}$ by an expected score $\mathcal{S}$ if and only if $\mathcal{S}(P,P_{0})\leq\mathcal{S}(P^{\prime},P_{0})$. The logarithmic scoring rule, given by $S(P,x)=-\log P(x)$, is the most common example and is related to the Kullback-Leibler divergence via $\mathrm{KL}(P^{\prime}\|P)=\mathcal{D}_{S}(P,P^{\prime})$. Other popular choices include the spherical score, cumulative ranked probability score (CRPS), the energy score, and squared maximum mean discrepancy (MMD), which for a kernel function $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{+}$ and the Dirac measure at $x$ denoted by $\updelta_{x}$ is given by
                                                                                                    $\mathrm{MMD}^{2}(P,\updelta_{x})$ ${}:={}$ $\mathbb{E}_{X\sim P,X^{\prime}\sim P}[k(X,X^{\prime})]-2\mathbb{E}_{X\sim P}[k(X,x)].$                                                                                                      (1)
Throughout the paper, we restrict our analysis to any score that satisfies the following minimal convexity assumption, which is satisfied for all scoring rules mentioned above.
Assumption 1.
The scoring rule $S(\cdot,x)$ is convex in its first argument for all $x\in\mathcal{X}$, and its corresponding divergence $\mathcal{D}_{S}$ is jointly convex in both arguments.
1.1 Key Ideas
When $P_{\theta}$ has density $p_{\theta}$, the Bayes posterior $\Pi_{n}$ and its associated posterior predictive distribution are
                                                                                                    $\Pi_{n}(A):=\frac{\int_{A}p_{\theta}(x_{1:n})\mathsf{d}\Pi(\theta)}{\int_{\Theta}p_{\theta}(x_{1:n})\mathsf{d}\Pi(\theta)},\quad\quad P_{\Pi_{n}}:=\int_{\Theta}P_{\theta}\mathsf{d}\Pi_{n}(\theta).$                                                                                                      (2)
In this set-up, the predictive distribution $P_{\Pi_{n}}$ is derived as a consequence of the Bayes posterior $\Pi_{n}$—which itself is primarily geared towards quantifying the uncertainty about which parameter $\theta$ best describes the observed data $x_{1:n}$. In this paper, we reverse that order: the posterior predictive $P_{Q}:=\int_{\Theta}P_{\theta}\mathsf{d}Q(\theta)$ is the object of interest, and we obtain a predictively oriented (PrO) posterior as the measure $Q_{n}$ which quantifies uncertainty about $\theta$ in a manner that is optimal for prediction.
The key motivation for this novel methodology is that while $P_{\Pi_{n}}$ is predictively optimal if $P_{0}\in\mathcal{M}_{\Theta}$ (see Aitchison, 1975), this optimality breaks down once $P_{0}\notin\mathcal{M}_{\Theta}$. Indeed, the Bayes posterior predictive can yield extremely poor predictive performance under even seemingly innocuous model misspecification (see, e.g., Grünwald and van Ommen, 2017). Though this is well-understood, robust Bayesian methods usually still treat prediction as a second-order concern, and are more concerned with parameter uncertainty. One family emblematic of this kind of research are Gibbs posteriors. (e.g. Jiang and Tanner, 2008; Bissiri et al., 2016; Zhang, 2006; Alquier, 2024; Syring and Martin, 2019; Martin and Syring, 2022; Knoblauch et al., 2018; Schmon et al., 2020; Chérief-Abdellatif and Alquier, 2020; Matsubara et al., 2022, 2023; Altamirano et al., 2023a, b; McLatchie et al., 2025; Duran-Martin et al., 2024; Laplante et al., 2025; Frazier et al., 2025). Many of these posteriors are based on an scoring rule evaluated on the sample, and given by, for $A\subseteq\Theta$,
                                                                                                    $Q_{n}^{\dagger}(A)$ ${}:={}$ $\frac{\int_{A}\exp\left\{-\frac{\lambda_{n}}{n}\sum_{i=1}^{n}S(P_{\theta},x_{i})\right\}\mathsf{d}\Pi(\theta)}{\int_{\Theta}\exp\left\{-\frac{\lambda_{n}}{n}\sum_{i=1}^{n}S(P_{\theta},x_{i})\right\}\mathsf{d}\Pi(\theta)},\quad\quad P_{Q_{n}^{\dagger}}:=\int_{\Theta}P_{\theta}\mathsf{d}Q_{n}^{\dagger}(\theta),$                                                                                                      (3)
where the tuning parameter $\lambda_{n}$ is often called the learning rate. Using Gibbs posteriors for parameter inference is often called generalised Bayesian inference because they recover the Bayes posterior for the special case where $S$ is the logarithmic scoring rule and $\lambda_{n}=n$. To contrast them with PrO posteriors, it will prove instructive to rewrite Gibbs posteriors as the outcome of the following variational optimisation problem:
                                                                                                    $Q_{n}^{\dagger}$ ${}={}$ $\operatornamewithlimits{argmin}_{Q\in\mathcal{P}(\Theta)}\left\{\frac{\lambda_{n}}{n}\sum_{i=1}^{n}\int S(P_{\theta},x_{i})\mathsf{d}Q(\theta)+\mathrm{KL}(Q\|\Pi)\right\}.$                                                                                                      (4)
The above representation is well-established (see e.g., Theorem 1 of Knoblauch et al., 2022), and makes clear that all Gibbs posteriors—including the Bayes posterior—seek to minimise the averaged scoring rule, rather than the score of the averaged model. This in turn means that the inferences they produce are targeted towards finding the parameter, rather than the predictive, which best describes the data. Crucially, these two goals do not overlap outside of well-specified models, where $P_{0}\in\mathcal{M}_{\Theta}$ (see Aitchison, 1975). The result is that in empirical applications, Bayes posteriors and Gibbs posteriors alike often deliver poorly performing posterior predictives.
What then should we do if we wish to predict from a parametric model, but also suspect that $P_{0}\notin\mathcal{M}_{\Theta}$? Now, we argue that the quantity of interest should be the predictive $P_{Q}:=\int P_{\theta}\mathsf{d}Q(\theta)$, and should thus be scored directly. Doing so gives rise to the PrO posterior given by
${Q}_{n}=\operatornamewithlimits{argmin}_{Q\in\mathcal{P}(\Theta)}\left\{\frac{\lambda_{n}}{n}\sum_{i=1}^{n}\widehat{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),x_{i}\right)+\mathrm{KL}(Q\|\Pi)\right\},$
where we write $\widehat{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),x_{i}\right)$ instead of ${S}\left(\int P_{\theta}\mathsf{d}Q(\theta),x_{i}\right)$ to highlight that throughout our theoretical and methodological development, PrO posteriors can be computed based on either exact or approximate predictive scoring rules.
1.2 Initial Comparison: Gibbs Posteriors vs PrO Posteriors
As we show later, PrO posteriors have several important properties, but the most fundamentally useful property is that they are adaptive: in contrast to Bayes and Gibbs posteriors, PrO posteriors adapt to model misspecification. We illustrate what this means by comparing the predictive distributions associated to $Q_{n}$ and $Q_{n}^{\dagger}$ under $\mathcal{M}_{\Theta}=\{P_{\theta}:\,\mathsf{N}(\theta,\sigma^{2}=1)\}$ for a range of data distributions. This model class is particularly educational here as the predictive is obtained from the posterior via smoothing with a Gaussian kernel with unit scaling. For predictive scores based on squared MMD, Figure 1 plots the results for the well-specified setting where $P_{0}\in\mathcal{M}_{\Theta}$ and three misspecified settings where $P_{0}\notin\mathcal{M}_{\Theta}$. While the PrO posterior behaves similarly to the MMD-based Gibbs posterior if the model is well-specified, it adapts to misspecification, and continues to produce reasonable predictive distributions across each setting. In contrast, Gibbs posteriors are blind to the mismatch between model and data.
As our example uses Gaussian likelihoods, one may misinterpret the phenomenon displayed as a cherry-picked pathology. In parametric models, however, Bayes posteriors and Gibbs posteriors are usually asymptotically Gaussian, and will concentrate around a single element of $\mathcal{M}_{\Theta}$. As we formally show later on, this property is not shared by PrO posteriors. On the contrary, depending on the severity of misspecification, they can be multi-modal and need not concentrate around a single element of $\mathcal{M}_{\Theta}$. Instead, they stabilise towards the predictively optimal average of elements in $\mathcal{M}_{\Theta}$ for $P_{0}$. As a result—and unlike Bayes posteriors and Gibbs posteriors—PrO posteriors produce predictively meaningful parameter uncertainty even under severe forms of misspecification.
Figure 1: Comparison of PrO posteriors and Gibbs posteriors in a well-specified and three misspecified regimes. Grey bars indicate a histogram of the observed data.
1.3 PrO Posteriors: A Third Modelling Culture?
In his prescient contribution, Breiman (2001) identified two divergent sub-cultures within statistical science. On the one hand, he spoke of a data modelling culture interested in hand-tailoring interpretable models for a given research problem and thus strongly associated to classical statistical concerns. On the other hand, he identified an increasing interest in machine learning and the associated algorithmic modelling culture focused on predictive accuracy and black box density estimation techniques. Does their reliance on a parametric base model $\mathcal{M}_{\Theta}$ make PrO posteriors part of the data modelling culture? Or does their focus on predictive performance make them into an offshoot of machine learning and the associated algorithmic modelling culture?
We believe that PrO posteriors constitute a meaningful and novel attempt to incorporate the most useful aspects of both modelling cultures: they rely on interpretable parameters born out of the data modelling culture, but gear inference towards predictive accuracy. When both interpretability and predictive performance are relevant, practitioners are often forced into a fraught choice: should they prefer an interpretable but predictively suboptimal model; or should they resort to a black box model that will improve predictive performance at the cost of interpretability and meaningful guarantees? PrO posteriors do not force us into this problematic binary choice: they rely on interpretable parameters born out of the data modelling culture, but gear inference towards predictive accuracy, and as a result perform well on both tasks. In Figure 2, we illustrate this by studying body measurements for different penguin species in the Palmer Archipelago (Horst et al., 2020). The PrO posterior not only identifies the data set’s fundamental structure by correctly discovering the three different species in it, but also produces a predictive distribution of excellent quality.
Figure 2: Palmer penguins example. We plot both the kernel density estimate (KDE) and the PrO posterior predictive overlaid with the data (left, middle), as well as the PrO posterior overlaid with the average measurements for each of the three penguin species in the data set (right). For inference with the PrO posterior, we use a bivariate Gaussian model. The results show that the PrO posterior not only correctly identifies the three penguin species in the data (right), but also provides a predictive distribution which can fit the observed data as well as a kernel density estimator. For further details, see Appendix D.2.
As we shall see in the remainder, this behaviour is not an accident: PrO posteriors provide an impactful third way of learning from data that combines predictive accuracy with interpretable uncertainty estimates. This behaviour is desirable for a wide array of scientific problems, including in digital twins, simulation models, causal modelling with observational data in public health and medicine, and forecasting applications in energy markets and finance. In operationalising a compromise between parameter inference and predictive accuracy, PrO posteriors also yield an altogether new kind of uncertainty quantification different from that of classical approaches. The next section elaborates on this, and explains why we can think of PrO posteriors as providing a novel type of epistemological uncertainty beyond what is currently attainable in classical Bayesian inference.
1.4 The Nature of PrO Posterior Uncertainty
To start, it is worth contrasting PrO posterior uncertainty expressed with a set of ideas and methods known under the umbrella of predictive Bayes (see e.g. Berti et al., 2021; Fortini and Petrone, 2020, 2023; Fong et al., 2023; Fortini and Petrone, 2025). Despite similar sounding names, the roles played by prediction in the construction of PrO posteriors and predictive Bayes approaches are fundamentally different: the PrO posterior is geared towards learning the predictive from seen data, predictive Bayes approaches instead posits a sequence of predictive distributions to model the unseen data. Put differently, while PrO posteriors make the predictive distribution the object of posterior inference, predictive Bayes methods specify a predictive distribution instead of a prior and likelihood.
In fact, specifying priors and likelihoods is just a convenient way of specifying a sequence of predictive distributions on unseen data: formalising this in a recent read paper at the Royal Statistical Society, Fong et al. (2023) used Doob’s Theorem to argue that the ultimate source of Bayesian uncertainty is unseen data. Taking this to its logical extreme, if we observe infinitely many data points, no uncertainty about the data remains and we can be perfectly confident in the conclusions we draw from data. This is in fact true more generally, as similar reasoning can be applied to various generalisations of Bayesian uncertainty, including the uncertainty expressed by Gibbs posteriors. One of the key implications is that both Bayesian and generalised Bayesian inference schemes result in a form of epistemological uncertainty that vanishes as the number of observations increase.
But should epistemological uncertainty indeed vanish as more data are observed? Though it may be tempting to answer this question in the affirmative, doing so would overlook an important practical limitation: for interpretability and computational reasons, statistical methods express epistemological uncertainty relative to a parameter $\theta\in\Theta$. Crucially, this implies that the only setting in which larger sample sizes should lead to a decrease in epistemological uncertainty is the rare situation where a singular value of $\theta\in\Theta$ can summarise all properties of $P_{0}$ that we are interested in. In other words, the kind of epistemological uncertainty conferred by (generalised) Bayesian methods must always be viewed relative only to the model class, not with respect to the choice of model class itself, rendering inference and particularly uncertainty quantification unreliable whenever $P_{0}\notin\mathcal{M}_{\Theta}$.
The epistemological uncertainty produced by PrO posteriors is more expansive: it delivers both parameter uncertainty and uncertainty about the inherent ability of $\mathcal{M}_{\Theta}$ to adequately explain the observed data. Indeed, when the model is misspecified, PrO posteriors generally do not collapse to a single element of $\mathcal{M}_{\Theta}$ as more data are observed. When $P_{0}\notin\mathcal{M}_{\Theta}$, they instead stabilise towards an irreducible form of uncertainty that originates in the fact that no single value of $\theta\in\Theta$ is sufficient to summarise all properties of $P_{0}$ we take an interest in. This is epistemological uncertainty in its truest form: because the model is misspecified, we should be uncertain about its ability to represent the observed data, even if we had observed the entire population $P_{0}$.
But what if the model is well-specified, so that $P_{0}\in\mathcal{M}_{\Theta}$? In this case, the epistemological uncertainty produced by Bayes and Gibbs posteriors is sufficient: there exists some $\theta\in\Theta$ that allows us to completely describe the observed data—and as more data accumulates, we should become increasingly certain about the ability of this value to correctly explain the data, and hence $P_{0}$. As it turns out, PrO posteriors square the circle: in the special case where the model is well-specified, they do behave like Bayes and Gibbs posteriors, and concentrate towards the parameter value that recovers $P_{0}$.
1.5 Contributions
We make three main contributions to demonstrate the advantages of PrO posteriors:
  1. C1.
    In Section 2, we establish a new mathematical representation that encompasses all PrO posteriors, as well as previous attempts at specifying similar objects.
  2. C2.
    In Sections 3 and 7, we summarise our substantive theoretical contributions. We formally establish that PrO posteriors predictively dominate Bayes posteriors and Gibbs posteriors alike, and prove that they are adaptive to model misspecification.
  3. C3.
    To make our theoretical contributions actionable, we also provide an asymptotically exact sampling strategy for PrO posteriors based on particle systems derived from Wasserstein gradient flows, and demonstrate its performance on several examples.
Our findings in C1 clarify and unify previous approaches, and craft a common lens through which to view a collection of somewhat disparate contributions over the years that used different justifications for objects that resemble PrO posteriors (e.g. Sheth and Khardon, 2020; Masegosa, 2020; Wei et al., 2021; Jankowiak et al., 2020b, a; Morningstar et al., 2022; Lai and Yao, 2024; Shen et al., 2025). Beyond that, C1 powers much of the theory we develop in C2, which in turn provides a rigorous justification for heuristic strategies in earlier literature which targeted variational approximations of $Q_{n}$ for the special case of the log score (see Sheth and Khardon, 2020; Wei et al., 2021; Jankowiak et al., 2020b, a). One important part of this new theory is a significantly strengthened and generalised version of the PAC-Bayes type bounds in related approaches, such as in Masegosa (2020) and Morningstar et al. (2022) and Morningstar et al. (2022) in Lemma 1. Unlike these previous attempts, our result is not restricted to any one particular score or approximation, but instead applies universally under mild regularity conditions. Despite the generality of our approach, we manage to dramatically improve previous bounds for the special case of the logarithmic scoring rule. In fact, our bounds are sufficiently sharp to allow for a characterisation of the qualitative behaviours of PrO posteriors (see Theorems 12). Lastly, C3 provides a generic approach for drawing samples from any PrO posterior—despite the fact that such objects generally have no known analytic form. This is very significant: aside from the special case of MMD-based PrO posteriors in Shen et al. (2025), all previous work had been focused on variational approximations (see e.g. Sheth and Khardon, 2020; Jankowiak et al., 2020b, a; Lai and Yao, 2024), and it was not clear whether a PrO posterior constructed from any other score could even be sampled from at all. As we show, this can indeed be achieved by suitably building on the Wasserstein gradient flow and the ideas of Wild et al. (2023), Wild et al. (2025), and Shen et al. (2025).
2 Background and Theoretical Groundwork
To set the stage for our theoretical developments, we introduce three key ideas that guide our methodological and theoretical innovations. In Section 2.1, we study representations and approximations of scoring rules whose structure leads to a tractable objective, from which we derive a novel PAC-Bayes type bound in Section 2.2, which serves as the foundation of our investigation into the predictive performance of PrO posteriors.
2.1 Exact and Approximate Representations of Predictive Scores
To simplify language and to highlight that PrO posteriors apply scoring rules to the predictive distribution $P_{Q}=\int P_{\theta}\mathsf{d}Q(\theta)$, the remainder of this paper will use the shorthand predictive scoring rule to refer to a scoring rule which is applied to the predictive distribution $P_{Q}$. One of the challenges of working with predictive scoring rules is that they do not admit simple representations. However, Assumption 1 and Jensen’s inequality imply that
                                                                                                    $S\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)$ ${}={}$ $\int S\left(P_{\theta},x\right)\mathsf{d}Q(\theta)-\Delta(Q,x),$                                                                                                      (5)
for all $x\in\mathcal{X}$ and some non-negative remainder term $\Delta:\mathcal{P}(\Theta)\times\mathcal{X}\to\mathbb{R}_{+}$. This representation has important practical ramifications: depending on the form of $\Delta(Q,x)$, a predictive score is either intractable, or can be tractably approximated. In the remainder, we call a predictive scoring rule tractable when it can be written as the expectation over a product measure for any $x\in\mathcal{X}$; so that for some $k\in\mathbb{N}_{+}$ and a loss $L:\Theta^{k}\times\mathcal{X}\rightarrow\mathbb{R}$, one obtains
$S\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)=\int L(\theta_{1:k},x)\mathsf{d}Q^{k}(\theta_{1:k}),\quad Q^{k}(\theta_{1:k})=\bigotimes_{i=1}^{k}Q(\theta_{i}).$ (6)
If one can sample from $Q$, this allows approximation to arbitrary precision via Monte Carlo methods. For example, kernel scores, such as MMD and CRPS, are tractable with $k=2$ (see Lemma 3 in Appendix B).
While tractable predictive scoring rules facilitate computation, several predictive scores of interest are intractable—most notably the logarithmic predictive scoring rule (see Lemma 4 in Appendix B). From the decomposition in (2.1), its associated Jensen’s gap depends on an intermediate value between $p_{\theta}(x)$ and $\int p_{\theta}(x)\mathsf{d}Q(\theta)$ denoted $g_{x}(p_{\theta},Q)$, and is
                                                                                                    $\Delta_{\log}(Q,x)$ ${}={}$ $\int\frac{1}{2g_{x}(p_{\theta_{1}},Q)^{2}}\left\{p_{\theta_{1}}(x)-\int p_{\theta_{2}}(x)\mathsf{d}Q(\theta_{2})\right\}^{2}\mathsf{d}Q(\theta_{1}),$                                                                                                      (7)
which is clearly intractable due to the dependence of $g_{x}(p_{\theta_{1}},Q)$ on $\int p_{\theta}(x)\mathsf{d}Q(\theta)$. To approximate $\Delta_{\log}$ in a tractable manner, Masegosa (2020) assumed that $\sup_{\theta\in\Theta,x\in\mathcal{X}}p_{\theta}(x)\leq u$ for some $u>0$, so that $\Delta_{\log}(Q,x)\geq\int\frac{1}{2\sup_{\theta^{\prime}\in\Theta}p_{\theta^{\prime}}(x)}\left\{p_{\theta_{1}}(x)-p_{\theta_{2}}(x)\right\}^{2}\mathsf{d}Q^{2}(\theta_{1:2})$, which leads to a tractable approximation, $\widehat{S}_{\operatorname{DI}}$, that upper-bounds the predictive logarithmic score via
                                                                                                    $S\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)\leq\widehat{S}_{\operatorname{DI}}\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)$ ${}:={}$ $\int L_{\operatorname{DI}}(\theta_{1:2},x)\mathsf{d}Q^{2}(\theta_{1:2}),$                                                                                                      (8)
                                                                                                    $L_{\operatorname{DI}}(\theta_{1:2},x)$ ${}:={}$ $-\log p_{\theta_{1}}(x)-\frac{\left\{p_{\theta_{1}}(x)-p_{\theta_{2}}(x)\right\}^{2}}{2\sup_{\theta\in\Theta}p_{\theta}(x)}.$                                                                                                      (9)
Since the term $\Delta_{\log}$ can be viewed as a variance or diversity term, Masegosa (2020) referred to this approximation strategy as diversity inducing. We will pay homage to this work by referring to all tractable predictive scoring rules constructed from the representation in (2.1)—whether exact or approximate—as diversity inducing (DI).
While the approximations suggested by Masegosa (2020) in the log score setting allowed for the derivation of a second-order PAC-Bayes bound, they are not practically appealing: $\sup_{x\in\mathcal{X},\theta\in\Theta}p_{\theta}(x)\leq u$ means that the bound only applies to models whose density or mass function is bounded, which rules out even simple models such as Gaussians. To address this deficiency, Morningstar et al. (2022) proposed a second type of tractable approximation based on multi sample (MS) bounds in variational inference. While originally proposed exclusively for the logarithmic predictive scoring rule, the approach clearly generalises to any convex score. In particular, for any $k\in\mathbb{N}$, we can define the tractable upper bound approximation $\widehat{S}_{\operatorname{MS}}$ as
                                                                                                    $S\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)\leq\widehat{S}_{\operatorname{MS}}^{k}\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)$ ${}:={}$ $\int L_{\operatorname{MS}}^{k}(\theta_{1:k},x)\mathsf{d}Q^{k}(\theta_{1:k}),$                                                                                                      (10)
                                                                                                    $L_{\operatorname{MS}}^{k}(\theta_{1:k},x)$ ${}:={}$ $S\left(\frac{1}{k}\sum_{j=1}^{k}P_{\theta_{j}},x\right).$                                                                                                      (11)
For larger choices of $k$, this bound tightens, and the approximation error vanishes as $k\to\infty$. For the special case of the logarithmic score, Morningstar et al. (2022) attempted to translate this into a PAC-Bayesian generalisation bound for a PrO posterior. However, the resulting bound was vacuous, and failed to discern whether the resulting posteriors conferred any advantages over Bayes posteriors. We rectify this omission, and deliver such a result: under mild regularity conditions and for arbitrary scoring rules, PrO posteriors based on $\widehat{S}^{k}_{\operatorname{MS}}\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)$ strictly improve predictive performance relative to Bayes posteriors and Gibbs posteriors.
To end this section, we will summarise and generalise our observations on previous approaches through a mild assumption that restricts how the exact or approximate predictive scores giving rise to PrO posteriors will be computed. Importantly, this assumption will ensure that we will not have to differentiate explicitly between the case of exact or approximate predictive scores. Instead, we only assume that the (exact or approximate) predictive score can be computed as an average of form
                                                                                                    $\widehat{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)$ ${}:={}$ $\int L(\theta_{1:k},x)\mathsf{d}Q^{k}(\theta_{1:k}),$                                                                                                      (12)
where $L$ is subject to further requirements that generalise both the DI-based approach of Masegosa (2020) and the MS-based approach of Morningstar et al. (2022).
Assumption 2.
$\widehat{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)=\int L(\theta_{1:k},x)\mathsf{d}Q^{k}(\theta_{1:k})$ is the approximate or exact predictive scoring rule, where $L:\Theta^{k}\times\mathcal{X}\to\mathbb{R}$ satisfies conditions $\operatorname{(DI)}$ or $\operatorname{(MS)}$ below.
  • $\operatorname{(DI)}$
    $L(\theta_{1:k},x)=S(P_{\theta_{1}},x)-\delta(\theta_{1:k},x)$, for some $\delta:\Theta^{k}\times\mathcal{X}\to\mathbb{R}_{+}$ and $\Delta(Q,x)\geq\int\delta(\theta_{1:k},x)\mathsf{d}Q(\theta)$, for all $x\in\mathcal{X}$ and $\theta_{1:k}\in\Theta^{k}$. Further, for any $\varphi\in\Theta$, if $\theta_{1:k}=(\varphi,\varphi,\dots\varphi)$, then $\delta(\theta_{1:k},x)=0$, for all $x\in\mathcal{X}$.
  • $\operatorname{(MS)}$
    $L(\theta_{1:k},x)=S(\frac{1}{k}\sum_{j=1}^{k}P_{\theta_{j}},x)$.
2.2 A Key Lemma for PrO Posteriors
Assumption 2 allows us to formally define the PrO posterior irrespective of the sub-case considered, which can now be conveniently represented in the form:
$Q_{n}=\operatornamewithlimits{argmin}_{Q\in\mathcal{P}(\Theta)}\left\{\lambda_{n}\int L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\mathrm{KL}(Q\|\Pi)\right\},$ (13)
where $\lambda_{n}>0$ is a positive scalar,
                                                                                                    $L_{n}(\theta_{1:k})$ ${}={}$ $\frac{1}{n}\sum_{i=1}^{n}L(\theta_{1:k},x_{i})$                                                                                                      (14)
denotes the in-sample loss corresponding to the (exact or approximate) score. This representation is the basis for a key Lemma that we build on throughout the remainder, and which is inspired by arguments prevalent for the study of Gibbs posteriors.
Crucially, the theoretical analysis of $Q_{n}$ will deviate sharply from standard approaches due to the appearance of product measures, the generally non-standard loss $L_{n}(\theta_{1:k})$, and the focus on the predictive $P_{Q}:=\int P_{\theta}\;\mathsf{d}Q(\theta)$. Despite these differences, we impose Hoeffding-type moment conditions on the function $L_{n}(\theta_{1:k})$ that are very similar to those arising in the study of Gibbs posteriors.
Assumption 3.
Define the expectation $\theta_{1:k}\mapsto\mathcal{L}(\theta_{1:k}):=\mathbb{E}[L(\theta_{1:k},X)]$. For any $\lambda\in\mathbb{R}$ and some $C_{k}>0$, $\mathop{\text{\large$\int$}}\nolimits\mathbb{E}\left[e^{\lambda\{\mathcal{L}(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\right]\mathsf{d}\Pi^{k}(\theta_{1:k})\leq e^{\lambda^{2}C_{k}^{2}/n}.$
Usually, the above condition is unproblematic: it automatically holds for scoring rules based on the squared MMD with uniformly bounded kernels. For the log score, the situation is more delicate. The log score must always be approximated, so Assumption 3 must be checked for a given approximation strategy. While this is a well-known drawback of imposing Hoeffding-type assumptions in the context of logarithmic scores (see e.g. Germain et al., 2016), it is generally hard to avoid them for this kind of analysis. Indeed, Assumption 3 was also crucial for the developments of both Masegosa (2020) and Morningstar et al. (2022).
With our three main assumptions in place, we can now state our first generic guarantee for PrO posteriors. Recalling that $P_{Q}=\int P_{\theta}\mathsf{d}Q(\theta)$, and $\mathcal{S}(P,P_{0})=\mathbb{E}[S(P,X)]$ with $\mathcal{D}_{S}(P_{Q},P_{0})=\mathcal{S}(P_{Q},P_{0})-\mathcal{S}(P_{0},P_{0})$, we obtain a PAC-Bayes type inequality.
Lemma 1.
Under Assumptions 1-3, the following two statements hold:
                                                                                                    $\mathbb{E}\left[\mathcal{S}(P_{Q_{n}},P_{0})\right]$ ${}\leq{}$ $\mathbb{E}\inf_{Q\in\mathcal{P}(\Theta)}\left\{\int L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\frac{\mathrm{KL}({Q^{k}\|\Pi^{k}})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}{k}C_{k}^{2}}{n};$                                                                                                      (15)
if the scoring rule is exact so that $\widehat{S}=S$, i.e., (6) holds, then
                                                                                                    $\mathbb{E}\left[\mathcal{D}_{S}(P_{Q_{n}},P_{0})\right]\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(P_{Q},P_{0}\right)+\frac{\mathrm{KL}({Q^{k}\|\Pi^{k}})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}{k}C_{k}^{2}}{n}.$                                                                                                      (16)
Lemma 1 delivers a meaningful preliminary bound on the predictive performance of PrO posteriors. Further, this result applies to PrO posterior based on any predictive score—whether approximate or exact—so long as it satisfies Assumptions 1-3. In the specific case of logarithmic score, it also recovers an expectation version of the bounds in both Masegosa (2020) and Morningstar et al. (2022). The implications are immediate. Lemma 1 provides the first known theoretical guarantee for the MMD-based PrO posterior proposed in Shen et al. (2025), and lends theoretical rigour to justify variational approaches seeking to approximate $Q_{n}$ (e.g. Sheth and Khardon, 2020; Wei et al., 2021; Jankowiak et al., 2020b, a; Lai and Yao, 2024).
3 Properties of PrO Posteriors
While Lemma 1 superficially resembles a PAC-Bayes bound used in the study of Gibbs posteriors, this in no way implies that PrO posteriors behave similarly to Gibbs posteriors. In fact, we will show that unlike Gibbs posteriors, PrO posteriors adapt to the level of discord between an assumed model and the observed data.
3.1 Model Misspecification Settings
Before investigating performance and adaptivity of PrO posteriors in more detail, we posit a taxonomy of model misspecification regimes under which we will compare the behaviour of the Gibbs and PrO posteriors.
Definition 2.
We say that $\mathcal{M}_{\Theta}$ is well-specified if $P_{0}\in\mathcal{M}_{\Theta}$, and misspecified if $P_{0}\notin\mathcal{M}_{\Theta}$. Further, we call a misspecified $\mathcal{M}_{\Theta}$ trivially misspecified with respect to a scoring rule $\mathcal{S}$ if there exists $\theta^{\prime}\in\Theta$ so that $\inf_{Q\in\mathcal{P}(\Theta)}\mathcal{S}(P_{Q},P_{0})\geq\mathcal{S}(P_{\theta^{\prime}},P_{0})$, and non-trivially misspecified if $\mathcal{S}(P_{Q},P_{0})<\inf_{\theta\in\Theta}\mathcal{S}(P_{\theta},P_{0})$ for some $Q\in\mathcal{P}(\Theta)$. Further, we say that a misspecified $\mathcal{M}_{\Theta}$ allows convex recovery if there exists $Q\in\mathcal{P}(\Theta)$ such that $P_{Q}=P_{0}$.
To aid understanding, Figure 3 provides an illustration of these different forms of model misspecification. What we call non-trivial misspecification is a generalisation of what Grünwald and van Ommen (2017) termed bad misspecification. It rules out forms of misspecification where there exists a singular model parametrisation that cannot be outperformed by any convex combination of models. Convex recovery is a special case of this, and implies that the extended model class $\{P_{Q}:Q\in\mathcal{P}(\Theta)\}$ obtained via convex combinations of $\mathcal{M}_{\Theta}$ contains $P_{0}$ and is therefore well-specified.
$\mathcal{M}_{\Theta}$$P_{\theta^{\star}}=\arg\min_{\theta}\mathcal{S}(P_{\theta},P_{0})$$P_{0}$
(a) Trivial misspecification.
$\mathcal{M}_{\Theta}$$P_{\theta_{2}}$$P_{\theta_{1}}$$P_{Q^{\star}}$$P_{\theta^{\star}}$$P_{0}$
(b) Non-trivial misspecification.
[…]
(c) Convex recoverability.
Figure 3: Illustration of key ideas introduced in Definition 2. Panel (a) depicts trivial model misspecification: relative to the chosen score $\mathcal{S}$, a singular element of $\mathcal{M}_{\Theta}$ provides the best fit for $P_{0}$, and there is no benefit to averaging. In contrast, Panel (b) illustrates the case of non-trivial model misspecification: constructing ${Q^{\star}}=\operatornamewithlimits{argmin}_{Q}\mathcal{S}(P_{Q},P_{0})$ as a convex combination of $P_{\theta_{1}}$ and $P_{\theta_{2}}$ yields a better predictive for $P_{0}$ than any one point in $\mathcal{M}_{\Theta}$. Panel (c) provides pictorial intuition for convex recoverability: it is the special case where $P_{Q^{\star}}$ and $P_{0}$ coincide.
3.2 Key Results
Equipped with the different forms of misspecification, we now compare the impact of different misspecification regimes on PrO posteriors $Q_{n}$, and compare them to Gibbs posteriors $Q_{n}^{\dagger}$. Whenever $Q_{n}$ and $Q_{n}^{\dagger}$ are compared directly, we assume that both are constructed from the same scoring rule. While Gibbs posteriors are always assumed to be constructed based on exact scoring rules, PrO posteriors may be based on approximations which adhere to $\operatorname{(DI)}$ or $\operatorname{(MS)}$ of Assumption 2.
We first present results applicable to exact and tractable predictive scores, and then obtain similar results for the setting of approximate predictive scores. Doing so reveals important differences in the approximation quality of approaches based on $\operatorname{(DI)}$ versus $\operatorname{(MS)}$, and allows us to obtain corollaries for PrO posteriors based on squared MMD as well as the logarithmic scoring rule. A full discussion of all underlying regularity conditions and a careful dissemination of the associated technical challenges is deferred to Section 7.
3.2.1 Tractable Predictive Scoring Rules
Before turning to the case of approximate predictive scores, it is informative to investigate the best case scenario: when the exact predictive scoring rule is tractable and (6) holds with equality, we obtain the strongest results.
Theorem 1.
Suppose that the predictive scoring rule is tractable so that (6) holds with equality. If Assumptions 13 and additional regularity conditions hold, then the following are satisfied for $n$ sufficiently large.
  • $\operatorname{(WS)}$
    If the model is well-specified, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]-\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]\leq\dfrac{\log(n)}{n^{1/2}}$;
  • $\operatorname{(NT)}$
    If the model is non-trivially misspecified, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]<\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]$.
    With further assumptions, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq\inf_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{\mathcal{S}}(P_{Q},P_{0})+\dfrac{\log(n)}{n^{1/2}}$;
  • $\operatorname{(CR)}$
    If the model allows convex recovery, then $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq\dfrac{\log(n)}{n^{1/2}}$.
[…]
The above result assumes that the exact predictive scoring rule is tractable, so that $\widehat{S}=S$ and (6) holds. Lemma 3 in Appendix B shows that this is satisfied for a number of scoring rules, for which Theorem 1 applies immediately.
Lemma 2.
For all kernel scores such as the squared MMD, the energy score, and the CRPS, the predictive scoring rule is tractable so that (6) holds exactly.
Interestingly, this result constitutes the only guarantees for the approach recently introduced in Shen et al. (2025), and shows that what was motivated as little more than a heuristic does in fact deliver strong and rigorous performance guarantees.
3.2.2 Approximate Scores
While the special case of exact predictive scores is very useful, certain scoring rules do not take the form in (6), and need to be approximated. To analyse what happens in this setting, one requires a refined version of Theorem 1 which introduces the difference between the exact and approximate score as $\operatorname{Gap}(k)$. While the below result denotes this difference generically, its magnitude and behaviour depend on the exact approximation chosen.
Theorem 2.
If Assumptions 13 and additional regularity conditions hold, then for the same $C_{k}$ as in Lemma 1, the following are satisfied for $n$ sufficiently large: for $n$ sufficiently large but finite, it holds for the same $C_{k}$ as in Lemma 1 that
  • $\operatorname{(WS)}$
    If the model is well-specified, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]-\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]\leq C_{k}\log(n)\sqrt{\frac{k}{n}}$;
  • $\operatorname{(NT)}$
    If the model is non-trivially misspecified, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]<\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]$.
    With further assumptions, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq\inf_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{\mathcal{S}}(P_{Q},P_{0})+\operatorname{Gap}(k)+C_{k}\log(n)\sqrt{\frac{k}{n}}$;
  • $\operatorname{(CR)}$
    If the model allows convex recovery, then $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq\operatorname{Gap}(k)+C_{k}\log(n)\sqrt{\frac{k}{n}}$.
Intriguingly, the above shows that so long as predictive scoring rules are approximated in line with Assumption 2, our key conclusions regarding the predictive performance of PrO posteriors are the same as in Theorem 1: the PrO posterior performs comparably well to the Gibbs posterior in the well-specified setting, and strictly dominates it in the non-trivially misspecified setting. In fact, the $\operatorname{Gap}(k)$ term is only relevnt in situations where we bound the expected predictive performance of the PrO posterior directly, and without comparison to the Gibbs posterior. To tackle these bounds, we need to delineate along the axes of $\operatorname{(DI)}$ and $\operatorname{(MS)}$ defined in Assumption 2. The case of DI-based approximations to predictive scores is straightforward: since $k$ is fixed and cannot be varied, the gap term reverts to a fixed constant. In other words, relative to exact predictive scores, PrO posteriors based on inexact DI-based approximations may lead to an irreducible error of size $\operatorname{Gap}(k)$. In contrast, for MS-based PrO posteriors, $\operatorname{Gap}(k)\to 0$ as $k\to\infty$. To ensure that the bounds become non-vacuous however, the increase in $k$ needs to be carefully balanced against $n$ to prevent $C_{k}\log(n)\sqrt{\frac{k}{n}}$ from diverging. The rate at which $k$ can be allowed to increase will vary depending on the score chosen. For example, for the special case of the logarithmic scores, we obtain the following result.
Corollary 1.
Under the conditions of Theorem 1, if we approximate the logarithmic predictive scoring rule as in $\operatorname{(MS)}$ of Assumption 2 and choose $k\propto\log(n)/n^{1/3}$, then there exists a constant $C>0$ so that $\operatorname{Gap}(k)+C_{k}\log(n)\sqrt{\frac{k}{n}}\leq C\cdot\log(n)/n^{1/3}$.
As expected, the result reveals that relative to computing PrO posteriors based on exact predictive scoring rules, MS-based approximations incur an additional cost and slow the rate of convergence to $\log(n)/n^{1/3}$. Despite the fact that the convergence slows, this is far from a negative finding. Indeed, to the best of our knowledge, Corollary 1 is the first known guarantee of its kind for PrO posterior type objects based on logarithmic scoring rules, and constitutes a strict improvement over previous bounds derived in Masegosa (2020) and Morningstar et al. (2022).
3.2.3 Posterior Concentration
The results in $\operatorname{(NT)}$ and $\operatorname{(CR)}$ of both Theorems 1 and 2 show that the PrO posterior predictive concentrates towards to an element close to $P_{0}$ in terms of $\mathcal{D}_{\mathcal{S}}$. For the special case of well-specified models, this suggests that the PrO posterior should also behave similarly to Gibbs posteriors in the parameter space $\Theta$, and concentrate towards a point mass at $\theta^{\star}=\operatornamewithlimits{argmin}_{\theta\in\Theta}\mathcal{D}_{\mathcal{S}}(P_{\theta},P_{0})$. Below, we formally prove this.
Corollary 2.
If Assumptions 13 and additional regularity conditions hold, if the model is well-specified, and if the predictive scoring rule satisfies $\operatorname{(MS)}$ of Assumption 2 or is exact, then $Q_{n}$ concentrates to $\theta^{\star}$ at rate $(\log(n)\sqrt{k/n})^{1/\alpha}$ for some $\alpha>1$.
The fact that the above result has to exclude PrO posteriors using DI-based approximate predictive scores is a direct consequence of the difficulty of controlling the rate at which its corresponding gap term $\operatorname{Gap}(k)$ vanishes as $n$ increases. However, the result still applies in most settings of interest, including for the exact predictive scores associated to the logarithmic score or the squared MMD. For both cases, one can show that the result holds with $\alpha=2$, which implies that the PrO posteriors based on these scores contract at the same rate. While the exact predictive logarithmic score is not computationally tractable, Corollary 2 also extends its MS-based approximation. This case is somewhat more involved however, and discussion is deferred to Sections 7.1.4 and 7.2.
4 Computation of PrO Posteriors
The tools for the computation of PrO posteriors are limited. In fact, most previous work targeting related objects did not even consider the possibility of exact computation, and instead focused on approximating $Q_{n}$ with a parameterised variational family (see e.g. Jankowiak et al., 2020a, b; Masegosa, 2020; Morningstar et al., 2022; Lai and Yao, 2024). This approach has several shortcomings: on a practical level, most variational approaches have to be fine-tuned to a given scoring rule, and depend on biased gradient estimates that suffer from high variance. More importantly however, as a PrO posterior is only defined through an optimisation problem and does not have an analytic form, it is almost impossible to ascertain if a given variational family is appropriate for capturing its most important characteristics. The only exception to this is the work of Shen et al. (2025), which built on the ideas in Wild et al. (2023) and Del Moral (2013) to derive an asymptotically exact sampler for the special case of PrO posteriors that used the squared MMD as the scoring rule. In the remainder, we will explain how this idea generalises to any PrO posterior whose scoring rule satisfies Assumption 2 so that $\widehat{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)=\int L(\theta_{1:k},x)\mathsf{d}Q^{k}(\theta_{1:k})$.
4.1 A Sampling Method for PrO Posteriors
Assuming for simplicity that $\Pi$, $Q$ and $Q^{k}$ have corresponding Lebesgue densities $\pi$, $q$ and $q^{k}$, we may rewrite the objective minimised in (13) as
                                                                                                    $\mathscr{L}(q)$ ${}={}$ $\lambda_{n}\int L_{n}(\theta_{1:k})q^{k}(\theta_{1:k})\mathsf{d}\theta_{1:k}-\int\log\pi(\theta)q(\theta)\mathsf{d}\theta+\int\log q(\theta)q(\theta)\mathsf{d}\theta,$                                                                                                      (17)
which we can think of as a polynomial function in $q$ with a strictly convex regulariser in the form of the entropy of $q$. For functionals like this, the Wasserstein gradient flow (WGF) can be used to construct a gradient descent scheme on the space of measures with finite moments. This in turn will allow us to compute it PrO posteriors with finite second moment by optimising the above functional. Implementing this optimisation procedure yields an interacting particle system derived from a McKean-Vlasov process. The resulting algorithm has been extensively discussed and analysed under the name of mean field Langevin dynamics in the monograph of Del Moral (2013) and was previously suggested in a similar context by Wild et al. (2023) and Shen et al. (2025).
To construct the particle system driving this algorithm, we first require the Wasserstein gradient of $\mathscr{L}(q)$. Denoting $v^{(\mapsto j)}[\vartheta]=(v_{1},\dots v_{j-1},\vartheta,v_{j},\dots v_{k-1})$ as the vector obtained by inserting $\vartheta\in\Theta$ at location $j$ of the vector $v$ with length k1k-1, we define $\widetilde{L}_{n}(\vartheta,\theta_{1:(k-1)},j)=L_{n}(\theta_{1:(k-1)}^{(\mapsto j)}[\vartheta])$ and obtain the first variation of $\mathscr{L}$ with respect to $q$ at $\vartheta$ as
                                                                                                    $\frac{\delta\mathscr{L}}{\delta q}[\vartheta]$ ${}={}$ $\lambda_{n}\int\sum_{j=1}^{k}\widetilde{L}_{n}(\vartheta,\theta_{1:(k-1)},j)\;q^{k-1}(\theta_{1:(k-1)})\mathsf{d}\theta_{1:(k-1)}-\log\pi(\vartheta)+\log q(\vartheta),$                                                                                                      (18)
from which one obtains the Wasserstein gradient after differentiating with respect to $\vartheta$. To simplify notation, we introduce the symmetrised loss function $L_{n}^{\operatorname{sym}}(\vartheta,\theta_{1:(k-1)})=\frac{1}{k}\sum_{j=1}^{k}\widetilde{L}_{n}(\vartheta,\theta_{1:(k-1)},j)$, denote its gradient with respect to $\vartheta$ by $\nabla_{1}L_{n}^{\operatorname{sym}}(\vartheta,\theta_{1:(k-1)})$, and the gradients of $\log\pi(\vartheta)$ and $\log q(\vartheta)$ by $\nabla\log\pi(\vartheta)$ and $\nabla\log q(\vartheta)$. With this, one obtains the Wasserstein gradient of $\mathscr{L}$ with respect to $q$ evaluated at $\vartheta$ as
                                                                                                    $\nabla_{\operatorname{W}}\mathscr{L}(q)[\vartheta]$ ${}={}$ $\lambda_{n}k\int\nabla_{1}L_{n}^{\operatorname{sym}}(\vartheta,\theta_{1:(k-1)})\;q^{k}(\theta_{1:(k-1)})\mathsf{d}\theta_{1:(k-1)}-\nabla\log\pi(\vartheta)+\nabla\log q(\vartheta).$                                                                                                      (19)
The WGF $\{q_{t}\}_{t>0}$ defined by this gradient is simply the continuous-time process that evolves in the direction of $\nabla_{\operatorname{W}}\mathscr{L}(q)$: the direction of steepest descent relative to the convex functional $\mathscr{L}(q)$. As $\mathscr{L}(q)$ is convex, and the minimiser of $\mathscr{L}(q)$ is precisely the PrO posterior, this means that we should think of the WGF as a continuously-valued version of gradient descent, and intuitively expect $q_{t}$ to converge to the PrO posterior as $t\to\infty$.
While we could in principle use the WGF to construct an optimisation scheme directly on the space of measures with finite second moments via a partial differential equation (PDE), this would generally lead to an intractable algorithm, as PDE solvers scale poorly with $\operatorname{dim}(\Theta)$ and $k$. With this in mind, we will instead resort to its alternative mean-field representation, which gives rise to the stochastic differential equation (SDE)
                                                                                                    $\mathsf{d}\vartheta_{t}$ ${}={}$ $\left\{\lambda_{n}k\hskip-2.84544pt\int\hskip-2.84544pt\nabla_{1}L_{n}^{\operatorname{sym}}(\vartheta_{t},\theta_{1:(k-1)})q_{t}^{k-1}(\theta_{1:(k-1)})\mathsf{d}\theta_{1:(k-1)}-\nabla\log\pi(\vartheta_{t})\right\}\mathsf{d}t+\sqrt{2}\mathsf{d}B_{t};\;\;\vartheta_{t}\sim q_{t},$                                                                                                      (20)
where $B_{t}$ denotes a Brownian motion. This term appears because the WGF of entropy corresponds exactly to the heat equation (Jordan et al., 1998), which in turn governs the evolution of Brownian motion. Unfortunately, the above SDE still depends on an intractable integral with respect to $q_{t}^{k-1}$, which makes exact forward simulation impossible. To remedy this, we use the approximation $q_{t}(\theta)\approx\frac{1}{p}\sum_{j=1}^{p}\updelta_{\vartheta_{t}^{(j)}}$, where $\updelta_{x}$ denotes the Dirac measure at $x$, and $\vartheta_{t}^{(j)}$ for $j=1,2,\dots p$ are a set of $p$ interacting particles. To state the necessary evolution equations, we take $\sigma(j,p,k)$ as the set of all $p-1\choose p-k$} possible permutations of the set $\{1,\dots j-1,j+1,\dots p\}$ with size k1k-1, and write the corresponding set of particles for any set of indices $\mathcal{I}\in\sigma(j,p,k)$ with $\mathcal{I}=\{i_{1},i_{2},\dots i_{k-1}\}$ as $\vartheta_{t}^{[\mathcal{I}]}=(\vartheta_{t}^{(i_{1})},\vartheta_{t}^{(i_{2})},\dots\vartheta_{t}^{(i_{k-1})})$. Based on this, we evolve the particles as
                                                                                                    $\mathsf{d}\vartheta_{t}^{(j)}$ ${}={}$ $\left\{\frac{\lambda_{n}k}{{p-1\choose p-k}}\sum_{\mathcal{I}\in\sigma(j,p,k)}\nabla_{1}L_{n}^{\operatorname{sym}}(\vartheta_{t}^{(j)},\vartheta_{t}^{[\mathcal{I}]})-\nabla\log\pi(\vartheta_{t}^{(j)})\right\}\mathsf{d}t+\sqrt{2}\mathsf{d}B_{t}^{(j)},$                                                                                                      (21)
where $\{B_{t}^{(j)}:j=1,2,\dots p\}$ denotes a set of $p$ independent Brownian motions. To obtain a high-fidelity approximation of the PrO posterior from this particle system in practice, we first have to discretise it into time steps $t_{1},t_{2},\dots$, and then evolve it for sufficiently many particles and a sufficiently long time. Afterwards, we average over all time steps past an initial burn-in period of length $\tau$ and approximate the PrO posterior as
                                                                                                    $Q_{n}(\theta)\approx\frac{1}{|\{t_{i}:t_{i}>\tau\}|}\sum_{t_{i}>\tau}\widehat{Q}[t_{i}](\theta)\quad\text{ for }\quad\widehat{Q}[t_{i}](\theta)=\frac{1}{p}\sum_{j=1}^{p}\updelta_{\vartheta_{t_{i}}^{(j)}}(\theta).$                                                                                                      (22)
Here, the notation highlights that this type of approximation should be thought of as an average of $|\{t_{i}:t_{i}>\tau\}|$ atomic measures $\widehat{Q}[t]$, each of which will recover $Q_{n}$ exactly in the limit as both $p\to\infty$ and $t\to\infty$. (Establishing this formally can be achieved via minor modifications of the arguments in Lemma 6 and Appendix E of Wild et al., 2023)
4.2 Examples: MMD and logarithmic score
Our above development recovers the algorithm of Shen et al. (2025) for the special case of the squared MMD. For this case, $k=2$, and the objective takes a particularly simple form. In particular, we have that $\widetilde{L}_{n}(\vartheta,\theta)=\frac{1}{n}\sum_{i=1}^{n}\mathsf{L}_{\operatorname{MMD}}(\vartheta,\theta,x_{i})$, where $\mathsf{L}_{\operatorname{MMD}}(\vartheta,\theta,x)=\langle\mu(P_{\vartheta})-\mu(\updelta_{x}),\mu(P_{\theta})-\mu(\updelta_{x})\rangle_{\mathcal{H}}$ is a symmetric function so that $\mathsf{L}_{\operatorname{MMD}}(\vartheta,\theta,x)=\mathsf{L}_{\operatorname{MMD}}(\theta,\vartheta,x)$. This symmetry simplifies the evolution equation and leads to
                                                                                                    $\mathsf{d}\vartheta_{t}^{(j)}=\left\{\frac{2\lambda_{n}}{p-1}\sum_{\begin{subarray}{c}\ell=1\\ \ell\neq j\end{subarray}}^{p}\nabla_{1}\widetilde{L}_{n}(\vartheta_{t}^{(j)},\vartheta_{t}^{(\ell)})-\nabla\log\pi(\vartheta_{t}^{(j)})\right\}\mathsf{d}t+\sqrt{2}\mathsf{d}B_{t}^{(j)}.$                                                                                                      (23)
Similarly, choosing MS-based approximations to the log score of order $k$, we find that $\widetilde{L}_{n}(\vartheta,\theta_{1:{(k-1)}})=\frac{1}{n}\sum_{i=1}^{n}\mathsf{L}_{\log}(\vartheta,\theta_{1:(k-1)},x_{i})$ for $\mathsf{L}_{\log}(\vartheta,\theta_{1:(k-1)},x)=\log\left\{\frac{1}{k}\sum_{j=1}^{k-1}P_{\theta_{j}}(x)+\frac{1}{k}P_{\vartheta}\right\}$, which is again symmetric in $(\vartheta,\theta_{1:(k-1)})$. Denoting by $\xi(j,p,k)$ the set of $p-1\choose k-1$} subsets (rather than permutations) of the set $\{1,\dots j-1,j+1,\dots p\}$ with size k1k-1, this symmetry simplifies the evolution equation as
$\mathsf{d}\vartheta_{t}^{(j)}=-\left\{\frac{\lambda_{n}k}{{p-1\choose k-1}}\sum_{\mathcal{I}\in\xi(j,p,k)}\nabla_{1}L_{n}^{\operatorname{sym}}(\vartheta_{t}^{(j)},\vartheta_{t}^{[\mathcal{I}]})+\nabla\log\pi(\vartheta_{t}^{(j)})\right\}\mathsf{d}t\\ +\sqrt{2}\mathsf{d}B_{t}^{(j)}.$
4.3 Practicalities and Implementation
Figure 4: Computation for the example of Section 1.2. The particle trajectories targeting the MMD-based PrO-posterior (in blue) compared to the path of a single MALA chain targeting the MMD-based Gibbs posterior (in red). The vertical dashed line indicates the burn-in period $\tau$. In the axes on the right are the empirical posterior densities from these two trajectories.
To provide intuition about how the sampling algorithm described in this section differs from more conventional Bayesian computation procedures, Figure 4 compares it to the Metropolis adjusted Langevin algorithm (MALA; Roberts and Tweedie, 1996; Roberts and Rosenthal, 1998) on the normal location example in Section 1.2. While the $p=32$ particles which target the PrO posterior are initialised randomly from the prior, they soon repel each other, and lead to a multi-modal distribution. In contrast, the single particle evolved according to MALA bounces around a singular mode.
While the interacting particle approach has the distinct advantage of being able to sample from any PrO posterior, it introduces several hyper-parameters that may require careful tuning in practice. The most important of these are the step sizes for the discretisation of the underlying SDE, the number of particles $p$, as well as the exact choice of $\lambda_{n}$ and $k$. Here, the choice of step size and the number of particles $p$ constitute a direct trade-off between approximation quality and computational efficiency: choosing the step size too small or the number of particles too large will in principle increase accuracy, but do so at the cost of exploding computational cost and substantively slowed convergence. While little is known about choosing these hyper-parameters in our setting, recent work by Chazal et al. (2025) may finally provide a principled way for this. In contrast, we have clear ideas how $k$ and $\lambda_{n}$ should be set from the theory developed in Section 3. While $k$ is generally fixed for DI-based approximations, we can freely choose it for MS-based approximations: by way of example, for the logarithmic score, one can show that $k\asymp\log(n)/n^{1/3}$ is the predictively optimal rate (see Corollay 1). Likewise, the learning rate $\lambda_{n}$ should be set to maximise the predictive performance. For the exact scoring rules as covered in Theorem 1 as well as approximate DI-based approaches, this is achieved by setting $\lambda_{n}\asymp n^{1/2}/\log(n)$. For MS-based approximations, the optimal choice depends on the optimal rate of growth for $k$. In the special case of the logarithmic score covered in Corollary 1, the optimal rate is the same as that for $k$, and given by $\lambda_{n}\asymp n^{1/3}/\log(n)$.
5 Illustrations
We end our development on PrO posteriors by illustrating some of their key behaviours. First, we explore the practical ramifications of convex recovery and the choice of different scoring rule approximations in Section 5.1 in the context of classification. Next, we show that similar conclusions hold for the continuous case by constructing in Section 5.2, and provide a more thorough investigation into how PrO posteriors achieve better predictive performance than Gibbs posteriors. To conclude, we highlight the potential future use of PrO posteriors for diagnosing model misspecification in a real world data set in Section 5.3.
5.1 Binary classification
For a first assessment of the predictive benefits conferred by PrO posteriors, we consider a simple synthetic classification example in which the model allows for convex recovery as set out in Definition 2. Labels $y_{i}$ are assigned deterministically as a function of the two-dimensional covariate vector $(x_{i}^{(0)},x_{i}^{(1)})\in[-2,2]^{2}$. Throughout, $(x_{i}^{(0)},x_{i}^{(1)})$ are sampled uniformly. When the covariates fall into the top-left quadrant of $2,2]^{2}$ (so that $x_{i}^{(0)}<0,x_{i}^{(1)}>0$), $y_{i}=0$. In contrast, the bottom-left quadrant (for which $x_{i}^{(0)}>0,x_{i}^{(i)}<0$) yields $y_{i}=1$, while labels in the remaining two quadrants are drawn from a Bernoulli distribution with parameter $0.5$.
We fit these data with a logistic classifier, and show the results in Figure 5. The plot compares the predictive performance of the Bayes posterior with that of a PrO posterior based on an approximation to the predictive log score. [^1] Here, we slightly abuse terminology: the Bayes posterior in this section is in fact tempered by $\lambda_{n}=n^{1/2}$. While this scaling term is required for our theory to apply, the experimental evidence in this section and its interpretation would be the same if the Bayes posterior were not tempered and $\lambda_{n}=1$. Since this approximation can be obtained either by following the DI-based approach of Masegosa (2020) that adheres to part $\operatorname{(DI)}$ of Assumption 2 or the MS-based approach which satisfies part $\operatorname{(MS)}$ and was originally put forward by Morningstar et al. (2022), we plot both of them. From the pictorial evidence, it is clear that relative to either approximation scheme, the Gibbs posterior produces a subpar decision boundary, and leads to overconfident and incorrect predictions—particularly in the top-right and bottom-left quadrants. In contrast, the PrO posteriors identify the correct decision boundaries, and recover a predictive distribution $P_{Q_{n}}$ for which $\mathrm{KL}(P_{0}\|P_{Q_{n}})$ is close to zero. This mirrors the predictions of Theorems 1 and 2, and is a consequence of the fact that the PrO posterior predictive allows for convex recovery of the exact decision boundary. Notably, there is little practical difference between the PrO posteriors in this setting: the predictive performance is similar whether we use the DI-based approximation or an MS-based approximation for $k=2$. This reinforces one of the key points of Theorem 2: even if the predictive score has to be approximated with an approximation error that cannot be recovered, the resulting PrO posterior will still predictively dominate the Bayes posterior.
Figure 5: Binary classification example. The first row shows the posterior predictive distributions of the Bayes posterior and the PrO posterior using two different approximations of the predictive logarithmic scoring rule; and is overlaid with the raw data. The second row describes the modes of the posterior predictive distributions, which gives rise to their decision boundaries.
5.2 Linear Regression
The previous section compared Gibbs and PrO posteriors for discrete-valued data. The following example extends this analysis, and compares performance in the context of a continuously-valued response when the model is non-trivially misspecified and allows for convex recovery. To do so, we generate outcomes from a mixture of linear regressions with the two-dimensional covariate $x_{i}\in\mathbb{R}^{2}$ distributed as standard normal, and with Gaussian errors $\varepsilon_{i}\sim\mathsf{N}(0,\sigma^{2})$. In particular, we generate them as
$y_{i}=\begin{cases}x_{i}^{\top}(0,2)+\varepsilon_{i}&\text{with probability }0.5\\ x_{i}^{\top}(0,-2)+\varepsilon_{i}&\text{otherwise}.\end{cases}$
We then fit these data using an MMD-based Gibbs posterior and a PrO posterior using the exact predictive score. To this end, we follow the formulation of MMD in the context of regression formulated by Alquier and Gerber (2024). Additional details on this and data generation protocols are deferred to Appendices D.1 and D.6.
We summarise our findings in the three panels of Figure 6. The left panel compares the posterior densities on the parameter space. It shows that the Gibbs posterior concentrates onto some convex combination of the underlying regression coefficients given by $(\theta_{1},\theta_{2})=(0,-c\cdot 2+(1-c)\cdot 2)$ for some $c\in(0,1)$. In contrast, the PrO posterior seems to concentrate onto the actual coefficients $(\theta_{1},\theta_{2})\in\{(0,-2),(0,2)\}$ which generated the data. This difference translates directly into predictive performance: the middle panel illustrates how the overconfidence of the Gibbs posterior results in a predictive distribution that is under-dispersed. The PrO posterior does not share this issue, and provides excellent predictive fit. This difference in predictive performance is quantified in the right panel, which demonstrates that the Gibbs posterior predictive performs roughly as well as the PrO posterior for data drawn from the one mode it focuses on, and substantively worse for all data from the mode which it neglects. Further details are provided in Appendix D.6.
Figure 6: Synthetic linear regression data. We show, from left to right, the joint posterior distribution of the PrO posterior and of the Gibbs posterior overlaid with the true regression coefficients (in black crosses); draws from their respective posterior predictive distributions are overlaid with the empirical distribution of the test data (in grey); and the difference between the PrO and Gibbs log point-wise predictive density for each test point. If the difference is positive, the PrO posterior predictive at that point is more accurate.
5.3 Putting Data
We end the section by considering a real world data set which can help us to compare the behaviour of Bayes and PrO posteriors in the case of non-trivial misspecification. Gelman and Nolan (2002) present data on professional golf putting success as a function of the distance from which the shot is taken. We show the proportion of successful putts in Figure 7 and fit a logistic regression to these data. The top row of this plot shows that the Bayes posterior predictive is overconfident, which is due to a highly concentrated parameter posterior. In contrast, the PrO posterior predictive is much better calibrated: it adapts to the present level of model misspecification, and does not concentrate around a single point, but instead onto two modes; each representing a distinct class of putting ability. In related but distinct contexts, this notable difference in uncertainties has led previous authors to speculate that one may be able to use PrO posterior-like methods as a diagnostic tool for model misspecification (see Lai and Yao, 2024; Chazal et al., 2025). While previous work argued for this on the basis of intuition and heuristics, our findings in Section 3 provide a way forward that would allow us a rigorous way for using PrO posteriors in the detection of model misspecification.
Figure 7: Golf putting data. Points represent the proportion of successful putts made by professional golfers as a function of the distance from the hole. The first column shows the predictive mean and $80\%$ credible intervals of the PrO posterior and of the Gibbs posterior. The second and third columns show the marginals of the corresponding parameter posteriors.
6 Discussion
In this paper, we have advocated for a new statistical paradigm that combines the best aspects of both parameter inference and density estimation. This led to PrO posteriors, which can operate at a similar predictive accuracy as density estimation techniques while simultaneously delivering practically meaningful uncertainty quantification for interpretable model parameters. We argued that PrO posteriors bridge the two main statistical traditions identified by Breiman (2001), and asserted that they lead to a version of epistemological uncertainty that is more expansive, adapts to model misspecification, and can avoid posterior collapse whenever doing so is benefits prediction. Further, we proved that PrO posteriors predictively dominate both Bayes and Gibbs posteriors. After introducing sampling algorithms for PrO posteriors, we illustrated our findings on several numerical examples.
PrO posteriors are a rigorous exploration of themes that have been present in several research communities, and build directly on the pioneering work of Masegosa (2020), Morningstar et al. (2022), and Shen et al. (2025). In a more applied setting, related ideas have also previously been studied for variational methods (Futami et al., 2021; Futami et al., 2022; Wei et al., 2021; Lai and Yao, 2024), in the context of Gaussian processes (Jankowiak et al., 2020a, b), and for mixture models (Wang and Liu, 2019). Within statistical learning, an important area of research with significant parallels to PrO posteriors are PAC-Bayesian bounds which are concerned with introducing diversity into an ensemble. The most important example is weighted majority voting, whereby a set of classifiers are assigned weights, and the final prediction is the label supported by the largest total weight across the ensemble (Lacasse et al., 2006; Laviolette et al., 2011; Germain et al., 2015; Laviolette et al., 2017; Lorenzen et al., 2019; Masegosa et al., 2020; Viallard et al., 2021; Wu et al., 2021; Wu and Seldin, 2022; Vignault et al., 2023; Theisen et al., 2024). Beyond majority voting, related PAC-Bayesian theories were also discussed in Audibert (2004), Ortega et al. (2022), Hauptvogel and Igel (2024) and Pitas et al. (2025).
Throughout, robustness to model misspecification was an important motivation in our development of PrO posteriors. Robustness is also one of the key drivers behind the recent emergence of generalised Bayesian methods, and particularly Gibbs posteriors: for example, Altamirano et al. (2023a, b); Matsubara et al. (2022, 2023); Laplante et al. (2025); Duran-Martin et al. (2024) all propose Gibbs posteriors which produce posteriors that are Bayes-like, but robust to outliers. However, these methods guard against misspecification only if our main interest remains parameter inference and our misspecification can be described through the famous Huber contamination model (Huber, 1964). If the misspecification falls far outside of this class of model perturbations, Gibbs posteriors constructed with this setting in mind generally perform poorly. Though the underlying meta-model is more intricate than the Huber contamination model, a similar conclusion holds for the coarsened Bayes posteriors of Miller and Dunson (2019). In contrast, PrO posteriors do not explicitly rely on a meta-model: while the results in Theorems 1 and 2 demonstrate strong performance guarantees whenever the data-generating process is a mixture over elements in $\mathcal{M}_{\Theta}$, they also show that PrO posteriors continue to dominate the predictive outputs of Gibbs and Bayes posteriors well beyond this setting.
We presented our findings as a post-Bayesian idea: grounded in the type of epistemological uncertainty produced by Bayesian methods, but not reliant on Bayes’ rule. It is interesting to note that one could also interpret our methodology as a nonparametric frequentist estimation technique: for the case of the logarithmic scoring rule, one could even interpret the PrO posterior objective as an KL-regularised version of a nonparametric maximum likelihood estimation (e.g. Laird, 1978; Lindsay, 1995). Extending this logic, PrO posteriors for general scoring rules could be thought of as frequentist KL-regularised nonparametric minimum distance estimators for the model class $\{P_{Q}:Q\in\mathcal{P}(\Theta)\}$ parameterised by a measure $Q\in\mathcal{P}(\Theta)$. However, as our parameter of interest is actually $\theta\in\Theta$, and since PrO posteriors directly incorporate our prior beliefs, $\Pi$, we argue that they should foremost be viewed as belief distributions over the parameters of interest, and are therefore more closely related to Bayesian and generalised Bayesian methods.
Despite a thorough exploration of PrO posteriors in the current paper, much remains to be done. Perhaps the most pressing issue is computation: to improve the sampling scheme we proposed, more insight is needed on hyper-parameter tuning and convergence diagnostics. We are optimistic that the recent paper of Chazal et al. (2025) may remedy this: on top of providing a new suite of diagnostics and convergence guarantees, it also proposes new algorithms applicable to PrO posteriors. Beyond computation, there are several promising methodological avenues: for instance, we should be able to derive statistical tests that use PrO posteriors to diagnose model misspecification. Unlike previous tests, they would not operate on posterior summaries, but directly on the difference in uncertainty between PrO posteriors and Gibbs posteriors. Further, the current theory is limited to observations which are independently drawn from a common data-generating mechanism. As there is considerable interest in forecasting and prediction for time series, graphical models, and spatial settings, it would be of great practical interest to extend our findings to dependent data. While this would be a technically challenging undertaking, our experimental evidence suggests that PrO posteriors should be expected to perform well in these settings. Other questions that are currently unaddressed include the choice of scoring rule: in finite samples, are there settings where one scoring rule should be preferred over another? What are the differences between PrO posteriors and their posterior predictives as we vary the underlying scoring rules? Likewise, we have not discussed prior choice in the current paper: though it seems intuitive to think of the prior as being conceptually similar to the generalised Bayesian case, we have not attempted to formalise this analogy. Doing so would have the potential of shedding more light on the role of prior regularisation in PrO posteriors, and could raise several key questions. For example, can we quantify how exactly the prior impacts the PrO posterior? Similarly, is it possible to adapt the arguments given by Catoni (2007) to PrO posteriors, and to derive priors which are in some sense predictively optimal?
In conclusion then, PrO posteriors open up many new possibilities, and allow us to reason about parameter uncertainty and prediction-centric statistical methods in a new light. They mark a clear departure from existing paradigms focused on either parameter inference or density estimation, and introduce a novel form of epistemological uncertainty unaccounted for by classical statistical methods. While the introduction of PrO posteriors results in numerous challenges that must be resolved before they are ready for wide-spread adoption, we believe that the scale of these challenges is matched by their transformative potential for modern statistical science.
Acknowledgements
This manuscript has benefited tremendously from interactions with many brilliant colleagues. Most notably, we feel deep gratitude towards Prof. Chris Oates for many conversations and insightful remarks. Further, we would like to thank Dr. Louis Sharrock and Dr. Sam Power for their help with the computational aspects of this paper. We are also grateful for correspondence with Prof. Pierre Alquier that pointed to the connection with PAC-Bayesian majority voting. Finally, we want to thank Dr. Evgenii Chzhen for insightful discussions on sub-Gaussianity. DT Frazier acknowledges funding from the Australian Research Council (DE200101070, DP200101414).
Appendix
Appendix A Notations
In this section we re-describe the main notations introduced in the main text to ensure the appendix is as self-contained as possible. We consider observed data ${x}_{1:n}=(x_{1},\dots,x_{n})$ for $x_{i}\in\mathcal{X}$ and $i=1,\dots,n$, where each $x_{i}$ is generated from an unknown distribution $P_{0}$. We denote expectations of functions $g:\mathcal{X}\to\mathbb{R}$ with respect to $X\sim P_{0}$ by $\mathbb{E}[g(X)]$ or by $\mathbb{E}_{X\sim P_{0}}[g(X)]$ if we wish to emphasise the distribution over which integration occurs. The model class is given by $\mathcal{M}_{\Theta}:=\{P_{\theta}:\theta\in\Theta\}$ and a prior measure $\Pi\in\mathcal{P}(\Theta)$ over elements of $\mathcal{M}_{\Theta}$. Likewise, let $\mathcal{P}(\Theta)$ denote the set of distributions over $\Theta$, and recall $P_{Q}=\int P_{\theta}\mathsf{d}Q(\theta)$ for $Q\in\mathcal{P}(\Theta)$. The function $S(P,x)$ denotes a scoring rule for the distribution $P$, while $\mathcal{S}(P,P)=\mathbb{E}_{X\sim P}[S(P,X)]$ denotes the expected score, and $\mathcal{D}_{S}(P,P_{0})=\mathcal{S}(P,P_{0})-\mathcal{S}(P_{0},P_{0})$ is the divergence associated to $P$, evaluated at the true distribution $P_{0}$. Further, define $S_{n}(P_{\theta})=\frac{1}{n}\sum_{i=1}^{n}S(P_{\theta},x_{i})$, and $L_{n}(\theta_{1:k})=\frac{1}{n}\sum_{i=1}^{n}L(\theta_{1:k},x_{i})$, where $L(\theta_{1:k},x)$ is as defined in Assumption 2, and which differs by the choice of approximation.
Appendix B Lemmas
This section contains several useful lemmas that are used throughout the remainder of the text. First, we demonstrate our claim that PrO posterior based on any kernel scoring rule are tractable with $k=2$. This implies that for scores based on squared maximum mean discrepancy (MMD2) and Cumulative Ranked Probability Score (CRPS), the function $\Delta(Q,x)$ in (2.1) is constant in $x$.
To prove this, take $x,x^{\prime}\in\mathcal{X}$, let $\kappa(x,x^{\prime})$ denote a positive-definite kernel function with feature map $\phi:\mathcal{X}\to\mathcal{H}$, such that $\kappa(x,x^{\prime})=\langle\phi(x),\phi(x^{\prime})\rangle_{\mathcal{H}}$. For fixed $\theta\in\Theta$, define the mean embedding $\mu_{\theta}\in\mathcal{H}$ by $\mu_{\theta}:=\int\phi(z)\,\mathsf{d}P_{\theta}(z)$ and the inner product $\langle\mu_{\theta},\mu_{\theta^{\prime}}\rangle_{\mathcal{H}}=\int\!\!\int\kappa(z,z^{\prime})\,\mathsf{d}P_{\theta}(z)\mathsf{d}P_{\theta^{\prime}}(z^{\prime}).$ Following Gneiting et al. (2007), kernel scoring rules can be expressed as $S(P,x)=\mathbb{E}_{X,X^{\prime}\sim P}[\kappa(X,X^{\prime})]-2\mathbb{E}_{X\sim P}[\kappa(X,x)]+\kappa(x,x).$
Lemma 3.
For any kernel scoring rule, and for all $x\in\mathcal{X}$: (i) The score $P_{\theta}\mapsto S(P_{\theta},x)$ is convex; (ii) $S(P_{Q},x)=\int L(\theta_{1:2},x)\mathsf{d}Q^{2}(\theta_{1:2})$ with $L(\theta_{1:2},x)=S(P_{\theta_{1}},x)-\frac{1}{2}\|\mu_{\theta_{1}}-\mu_{\theta_{2}}\|^{2}_{\mathcal{H}}.$
Proof.
To see that (i) is satisfied, use the feature map representation to write the score as
$S(P_{\theta},x)=\|\mu_{\theta}\|_{\mathcal{H}}^{2}-2\langle\mu_{\theta},\phi(x)\rangle_{\mathcal{H}}+\|\phi(x)\|_{\mathcal{H}}^{2}.$
The map $P_{\theta}\mapsto\mu_{\theta}$ is linear, so that $P_{\theta}\mapsto\|\mu_{\theta}\|_{\mathcal{H}}^{2}$ is quadratic. Thus, $P_{\theta}\mapsto S(P_{\theta},x)$ is the composition of a linear map and a convex function, and is convex. To deduce the second result, write
[…]
Expand the first integral, and use the mean embedding definition for $\mu_{\theta}$, to write
$S\left(P_{Q},x\right)=\int\int\int\int\kappa(z,z^{\prime})\,\mathsf{d}P_{\theta}(z)\mathsf{d}P_{\theta^{\prime}}(z^{\prime})\mathsf{d}Q(\theta)\mathsf{d}Q(\theta^{\prime})$ $=\langle\mathbb{E}_{\theta\sim Q}(\mu_{\theta}),\mathbb{E}_{\theta^{\prime}\sim Q}(\mu_{\theta^{\prime}})\rangle$
$=\|\mathbb{E}_{\theta\sim Q}(\mu_{\theta})\|^{2}_{\mathcal{H}}.$
Similarly, write
$\int S\left(P_{\theta},x\right)\mathsf{d}Q(\theta)$ $=\int\int\int\kappa(z,z^{\prime})\,\mathsf{d}P_{\theta}(z)\mathsf{d}P_{\theta}(z^{\prime})\mathsf{d}Q(\theta)=\mathbb{E}_{\theta\sim Q}[\|\mu_{\theta}\|^{2}_{\mathcal{H}}].$
Hence, for $\Delta(Q,x)$ as in (2.1),
$-\Delta(Q,x)=\|\mathbb{E}_{\theta\sim Q}[\mu_{\theta}]\|^{2}_{\mathcal{H}}-\mathbb{E}_{\theta\sim Q}[\|\mu_{\theta}\|^{2}_{\mathcal{H}}]$ $=-\mathbb{E}_{\theta\sim Q}\left[\|\mu_{\theta}-\mathbb{E}_{\theta^{\prime}\sim Q}[\mu_{\theta^{\prime}}]\|^{2}_{\mathcal{H}}\right]$
$=-\frac{1}{2}\int\|\mu_{\theta}-\mu_{\theta^{\prime}}\|^{2}_{\mathcal{H}}\mathsf{d}Q(\theta)\mathsf{d}Q(\theta^{\prime}).$
Lemma 4.
The log score satisfies Assumption 1. Further, for all $x\in\mathcal{X}$,
                                                                                                    $-\log\int p_{\theta}(x)\mathsf{d}Q(\theta)$ ${}={}$ $-\int\log p_{\theta}(x)\mathsf{d}Q(\theta)-\Delta_{\log}(Q,x)$                                                                                                      (29)
where for $g(p_{\theta}(x),p_{Q}(x))$ some intermediate value between $p_{\theta}(x)$ and $\int p_{\theta}(x)\mathsf{d}Q(\theta)$, we have
                                                                                                    $\Delta_{\log}(Q,x)$ ${}={}$ $\int\frac{1}{2g(p_{\theta}(x),p_{Q}(x))^{2}}\left\{p_{\theta}(x)-\int p_{\theta^{\prime}}(x)\mathsf{d}Q(\theta^{\prime})\right\}^{2}\mathsf{d}Q(\theta).$                                                                                                      (30)
Proof of Lemma 4.
For $\mu(x)=\int p_{\theta}(x)\mathsf{d}Q(\theta)$, the function $\log\mu(x)$ is convex, since $\log\mu(x)$ is concave. Thus, Assumption 1 is satisfied for the log score. Further, it is well-known that the KL-divergence is jointly convex in both arguments.
To prove the second part of the result, let $y(x)=p_{\theta}(x)$, and consider a second-order Taylor expansion of $\log y(x)$, around $\mu(x)$: for $g(y(x),\mu(x))$ is a real value between $y(x)$ and $\mu(x)$,
$\log y(x)=$ $\log\mu(x)+\frac{1}{\mu(x)}(y(x)-\mu(x))-\frac{1}{2g(y(x),\mu(x))^{2}}(y(x)-\mu(x))^{2}$
$=$ $\log\mathbb{E}_{\theta^{\prime}\sim Q(\theta^{\prime})}[p_{\theta^{\prime}}(x)]-\frac{1}{2g\{p_{\theta}(x),\int p_{\theta}(x)\mathsf{d}Q(\theta)\}^{2}}\left\{p_{\theta}(x)-\mathbb{E}_{\theta^{\prime}\sim Q(\theta^{\prime})}[p_{\theta^{\prime}}(x)]\right\}^{2}.$
Therefore,
$-\log\int p_{\theta}(x)\mathsf{d}Q(\theta)=$ $-\int\log p_{\theta}(x)\mathsf{d}Q(\theta)-\int\frac{\left\{p_{\theta}(x)-\int p_{\theta^{\prime}}(x)\mathsf{d}Q(\theta^{\prime})\right\}^{2}}{2g\{p_{\theta}(x),p_{Q}(x)\}^{2}}\mathsf{d}Q(\theta)$
$\Delta_{\log}(Q,x)=$ $\int\frac{\left\{p_{\theta}(x)-\int p_{\theta^{\prime}}(x)\mathsf{d}Q(\theta^{\prime})\right\}^{2}}{2g\{p_{\theta}(x),p_{Q}(x)\}^{2}}\mathsf{d}Q(\theta).$
Lemma 5.
Under Assumptions 1-4, for any $\lambda_{n}\geq d_{S}/r_{0}$, we have
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}Q_{n}^{\dagger}(\theta)\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\frac{d_{S}}{\lambda_{n}}\,\log\left(\dfrac{e\cdot c_{S}\lambda_{n}}{d_{S}}\right)+\frac{\lambda_{n}C^{2}}{n}\,.$
Furthermore, for the specific choice $\lambda_{n}=n^{1/2}\,\sqrt{\frac{d_{S}}{C^{2}}\,\log\left(\frac{ec}{d_{S}}\right)}$, for any $n\,\geq\frac{d_{S}C^{2}}{r_{0}^{2}\,\log\left(\frac{e\cdot c_{S}}{d_{S}}\right)}$, we have
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}^{\dagger}},P_{0}\right)\right]\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+C\left(\log n/n\right)^{1/2}.$
Proof.
The proof follows similar arguments to those used to prove Lemma 1 but where we take $L(\theta,\theta^{\prime})=R(\theta)=\mathcal{S}(P_{\theta},P_{0})-\mathcal{S}(P_{0},P_{0})$ and $L_{n}(\theta,\theta^{\prime})=R_{n}(\theta)=S_{n}(P_{\theta})-S_{n}(P_{0}).$ If Assumption 3 is satisfied for $R(\theta)$ and $R_{n}(\theta)$, then
$1\geq\mathbb{E}\int e^{\lambda_{n}\left\{R(\theta)-R_{n}(\theta)\right\}-\frac{\lambda_{n}^{2}C^{2}}{n}}\mathsf{d}\Pi(\theta).$
Apply the Donsker-Varadhan lemma with $h(\theta)=R(\theta)-R_{n}(\theta)$ to obtain
$1\geq\mathbb{E}e^{\sup_{Q\in\mathcal{P}(\Theta)}\lambda_{n}\int\left[R(\theta)-R_{n}(\theta)\right]\mathsf{d}Q(\theta)-\mathrm{KL}(Q\|\Pi)-\frac{\lambda_{n}^{2}C^{2}}{n}}.$
Apply Jensen’s inequity and take logs to obtain
$0\geq\mathbb{E}\sup_{Q\in\mathcal{P}(\Theta)}\left\{\lambda_{n}\int\left[R(\theta)-R_{n}(\theta)\right]\mathsf{d}Q(\theta)-\mathrm{KL}(Q\|\Pi)-\frac{\lambda_{n}^{2}C^{2}}{n}\right\}.$
Since the above is true for any $Q\in\mathcal{P}(\Theta)$, it is also true for $Q=Q^{\dagger}_{n}$. Applying this choice and re-arranging terms yields
$\mathbb{E}\int R(\theta)\mathsf{d}Q^{\dagger}_{n}(\theta)\leq\mathbb{E}\left\{\int R_{n}(\theta)\mathsf{d}Q^{\dagger}_{n}(\theta)+\frac{\mathrm{KL}({Q^{\dagger}_{n}\|\Pi})}{\lambda_{n}}+\frac{\lambda_{n}C^{2}}{n}\right\}.$
Use the definitions of $R(\theta)$ and $R_{n}(\theta)$, and Assumption 1, to obtain the inequality
$\mathbb{E}\left[\mathcal{S}(P_{Q_{n}^{\dagger}},P_{0})-\mathcal{S}(P_{0},P_{0})\right]$
$\leq$ $\mathbb{E}\left[\int\{\mathcal{S}(P_{\theta},P_{0})-\mathcal{S}(P_{0},P_{0})\}\mathsf{d}Q^{\dagger}_{n}(\theta)\right]$ (31)
$\leq$ $\mathbb{E}\left\{\int\left\{S_{n}(P_{\theta})-S_{n}(P_{0})\right\}\mathsf{d}Q^{\dagger}_{n}(\theta)+\frac{\mathrm{KL}({Q^{\dagger}_{n}\|\Pi})}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}$
$\leq$ $\mathbb{E}\inf_{Q\in\mathcal{P}(\Theta)}\left\{\int\left\{S_{n}(P_{\theta})-S_{n}(P_{0})\right\}\mathsf{d}Q(\theta)+\frac{\mathrm{KL}({Q\|\Pi})}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}$
$=$ $\mathbb{E}\inf_{Q\in\mathcal{P}(\Theta)}\left\{\int\left\{S_{n}(P_{\theta})-S_{n}(P_{0})\right\}\mathsf{d}Q(\theta)+\frac{\mathrm{KL}({Q\|\Pi})}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}$ (32)
Define the distributions $\Pi_{\mathcal{B}_{r}}$ indexed by a radius $r>0$ with corresponding density:
$\pi_{\mathcal{B}_{r}}(\theta):=\frac{1(\theta\in\mathcal{B}_{r})\pi(\theta)}{\Pi(\mathcal{B}_{r})}\quad\textnormal{with}\quad\mathcal{B}_{r}:=\left\{\theta\in\Theta:\mathcal{D}_{S}(P_{\theta},P_{0})\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+r\right\}.$
We can now further upper bound (31) by restricting the infimum on the right-hand side of (32) to distributions belonging to the set $\{\Pi_{\mathcal{B}_{r}}:r>0\}\subset\mathcal{P}(\Theta)$, which yields
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}Q_{n}^{\dagger}(\theta)$ $\leq\mathbb{E}\inf_{r>0}\left\{\int\left\{S_{n}(P_{\theta})-S_{n}(P_{0})\right\}\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\mathrm{KL}({\Pi_{\mathcal{B}_{r}}\|\Pi})}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}$
$\leq\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\mathrm{KL}({\Pi_{\mathcal{B}_{r}}\|\Pi})}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}$
$=\inf_{r>0}\left\{\mathop{\text{\large$\int$}}\nolimits\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\log\left\{1/\Pi(\mathcal{B}_{r})\right\}}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}$
$\leq\inf_{0<r\leq r_{0}}\left\{\mathop{\text{\large$\int$}}\nolimits\left\{\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+r\right\}\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{d_{S}\,\log\left(c_{S}/r\right)}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}$
$=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\inf_{0<r\leq r_{0}}\left\{r+\frac{d_{S}\,\log\left(c_{S}/r\right)}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}\,.$
As soon as $\lambda_{n}\geq\frac{d_{S}}{r_{0}}$, taking $r=\frac{d_{S}}{\lambda_{n}}\leq r_{0}$ then leads to
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}Q_{n}^{\dagger}(\theta)$ $\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\frac{d_{S}}{\lambda_{n}}+\frac{d_{S}}{\lambda_{n}}\,\log\left(\dfrac{c_{S}\lambda_{n}}{d_{S}}\right)+\frac{\lambda_{n}C^{2}}{n}$
$=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\frac{d_{S}}{\lambda_{n}}\,\log\left(\dfrac{e\cdot c_{S}\lambda_{n}}{d_{S}}\right)+\frac{\lambda_{n}C^{2}}{n}\,$,
which establishes the first result.
To obtain the second, substitute $\lambda_{n}=n^{1/2}\,\sqrt{\frac{d_{S}}{C^{2}}\,\log\left(\frac{e\cdot c_{S}}{d_{S}}\right)}$ in the bound, for $n\,\geq\frac{d_{S}C^{2}}{r_{0}^{2}\,\log\left(\frac{e\cdot c_{S}}{d_{S}}\right)}$, to obtain
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}Q_{n}^{\dagger}(\theta)$
$\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\frac{C\sqrt{d_{S}}}{n^{1/2}}\frac{\log\left(n^{1/2}\,\dfrac{e\cdot c_{S}}{C\sqrt{d_{S}}}\sqrt{\log\left(\frac{e\cdot c_{S}}{d_{S}}\right)}\right)}{\sqrt{\log\left(\frac{e\cdot c_{S}}{d_{S}}\right)}}+\frac{C\sqrt{d_{S}}}{n^{1/2}}\sqrt{\log\left(\frac{e\cdot c_{S}}{d_{S}}\right)}\,$,
and by convexity, Assumption 1 implies that $\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]\leq\mathbb{E}\mathop{\text{\large$\int$}}\nolimits\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}Q_{n}^{\dagger}(\theta),$ which ends the proof. ∎
Appendix C Proofs of Main results
Proof of Lemma 1.
Let us consider any fixed integer $k>1$. Recall $Q^{k}=\bigotimes_{j=1}^{k}Q$. Apply the Donsker-Varadhan Lemma with $h(\theta_{1:k})=k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}$, to see that, $P_{0}$-almost surely, $\forall Q\in\mathcal{P}(\Theta)$, we have
$\mathop{\text{\large$\int$}}\nolimits k\lambda_{n}\left\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\right\}\mathsf{d}Q^{k}(\theta_{1:k})\leq\text{KL}(Q^{k}\|\Pi^{k})+{\log}\mathop{\text{\large$\int$}}\nolimits e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k}).$
Rearranging terms, we have $P_{0}$-almost surely, $\forall Q\in\mathcal{P}(\Theta)$,
$\mathop{\text{\large$\int$}}\nolimits L(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})\leq\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}+\frac{{\log}\mathop{\text{\large$\int$}}\nolimits e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k})}{k\lambda_{n}},$
and $P_{0}$-almost surely, for $Q=Q_{n}$ in particular:
$\mathop{\text{\large$\int$}}\nolimits L(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})\leq\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})+\frac{\text{KL}(Q_{n}^{k}\|\Pi^{k})}{k\lambda_{n}}+\frac{{\log}\mathop{\text{\large$\int$}}\nolimits e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k})}{k\lambda_{n}}.$
Taking expectation with respect to $P_{0}$ on both sides and applying Jensen’s inequality to the log then leads to:
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})$ $\leq\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})+\frac{\mathbb{E}\text{KL}(Q_{n}^{k}\|\Pi^{k})}{k\lambda_{n}}+\frac{\mathbb{E}\log\mathop{\text{\large$\int$}}\nolimits e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k})}{k\lambda_{n}}$
$\leq\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})+\frac{\mathbb{E}\text{KL}(Q_{n}^{k}\|\Pi^{k})}{k\lambda_{n}}+\frac{\log\mathbb{E}\mathop{\text{\large$\int$}}\nolimits e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k})}{k\lambda_{n}}.$
The last term on the right-hand-side is controlled using Tonelli’s theorem and Assumption 3, under which:
$\log\mathbb{E}\mathop{\text{\large$\int$}}\nolimits e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k})=\log\mathop{\text{\large$\int$}}\nolimits\mathbb{E}e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k})\leq\frac{k^{2}\lambda_{n}^{2}C_{k}^{2}}{n},$
which then gives
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})$ $\leq\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})+\frac{\mathbb{E}\text{KL}(Q_{n}^{k}\|\Pi^{k})}{k\lambda_{n}}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$=\mathbb{E}\left[\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})+\frac{\text{KL}(Q_{n}^{k}\|\Pi^{k})}{k\lambda_{n}}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}.$
Recall that $Q_{n}$ is the minimiser of the quantity inside the expectation on the right-hand-side to finally write:
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})\leq\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}.$
To obtain the first result of the theorem, we just need to lower bound the expected integral on the left-hand-side. From the convexity in Assumption 1, for the DI-based approximation (Assumption 2 (DI)), we have
$\mathcal{S}\left(\int P_{\theta}\mathsf{d}Q_{n}(\theta),P_{0}\right)$ $=\int\mathcal{S}(P_{\theta},P_{0})\mathsf{d}Q_{n}(\theta)-\mathbb{E}\{\Delta(Q_{n},X)\}$
$\leq\int\mathcal{S}(P_{\theta},P_{0})\mathsf{d}Q_{n}(\theta)-\int\delta(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})\equiv\int L(\theta_{1:k})\mathsf{d}Q_{n}(\theta_{1:k}),$
where $\delta(\theta_{1:k})=\mathbb{E}[\delta(\theta_{1:k},X)]$; alternatively, by Assumption 1, under the MS-class of approximations (Assumption 2 (MS)), we have
$\mathcal{S}\left(\int P_{\theta}\mathsf{d}Q_{n}(\theta),P_{0}\right)$ $\leq\int\mathcal{S}\left(\frac{1}{k}\sum_{j=1}^{k}P_{\theta_{j}},P_{0}\right)\mathsf{d}Q_{n}^{k}(\theta_{1:k})\equiv\int L(\theta_{1:k})\mathsf{d}Q_{n}(\theta_{1:k}).$
Finally, and recalling $P_{Q_{n}}=\int P_{\theta}\mathsf{d}Q_{n}(\theta)$, this gives us
$\mathbb{E}\left[\mathcal{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq\mathbb{E}\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$ (33)
as desired.
To obtain the second result, notice that when no approximation is used, we have for any $Q$ and any $x$ that
$S\left(P_{Q},x\right)=\widehat{S}\left(P_{Q},x\right)\equiv\int L(\theta_{1:k},x)\mathsf{d}Q^{k}(\theta_{1:k}),$
which implies that for any $Q$, we have
$S_{n}(P_{Q})\equiv\frac{1}{n}\sum_{i=1}^{n}S\left(P_{Q},x_{i}\right)=\frac{1}{n}\sum_{i=1}^{n}\int L(\theta_{1:k},x_{i})\mathsf{d}Q^{k}(\theta_{1:k})\equiv\int L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k}).$
Using the above, we can rewrite equation (33) as
$\mathbb{E}\left\{\mathcal{S}\left(P_{Q_{n}},P_{0}\right)\right\}$ $\leq\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$=\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{S_{n}(P_{Q})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{S}(P_{Q},P_{0})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}.$
Recalling that $\mathcal{S}(P_{0},P_{0})=\mathbb{E}S_{n}(P_{0})$, we can subtract $\mathcal{S}(P_{0},P_{0})$ from both sides and write
$\mathbb{E}\{\mathcal{D}_{S}(P_{Q_{n}},P_{0})\}$ $=\mathbb{E}\left\{\mathcal{S}\left(P_{Q_{n}},P_{0}\right)-\mathcal{S}(P_{0},P_{0})\right\}$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{S}(P_{Q},P_{0})-\mathcal{S}(P_{Q},P_{0})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$=\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}(P_{Q},P_{0})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}.$
Proof of Theorem 3.
Recalling the definition of $\mathcal{B}_{r}$ in Assumption 4, define the candidate density
$\pi_{\mathcal{B}_{r}}(\theta):=\frac{1(\theta\in\mathcal{B}_{r})\pi(\theta)}{\Pi(\mathcal{B}_{r})}.$
Now, the RHS of the inequality in Lemma 1 is true for any $Q\in\mathcal{P}(\Theta)$, and remains true for $Q=\Pi_{\mathcal{B}_{n}}$. Under this choice for $Q$, and for $\delta_{n}(Q^{k})=\frac{1}{n}\sum_{i=1}^{n}\int\delta(\theta_{1:k},x_{i})\mathsf{d}Q^{k}(\theta_{1:k})$, equation (33) in Lemma 1 can be restated as
$\mathbb{E}\left[\mathcal{D}_{S}(P_{Q_{n}},P_{0})\right]$ $\leq\mathbb{E}\left[\inf_{r>0}\left\{\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}\Pi_{\mathcal{B}_{r}}^{k}(\theta_{1:k})-\mathcal{S}(P_{0},P_{0})+\frac{\text{KL}(\Pi_{\mathcal{B}_{r}}^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$=\mathbb{E}\left[\inf_{r>0}\left\{\mathop{\text{\large$\int$}}\nolimits S_{n}(\theta)\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta_{k})-\mathcal{S}(P_{0},P_{0})-\delta_{n}(\Pi_{\mathcal{B}_{r}}^{k})+\frac{\text{KL}(\Pi_{\mathcal{B}_{r}}^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\inf_{r>0}\left[\mathbb{E}\left\{\mathop{\text{\large$\int$}}\nolimits S_{n}(\theta)\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)-\mathcal{S}(P_{0},P_{0})-\delta_{n}(\Pi_{\mathcal{B}_{r}}^{k})+\frac{\text{KL}(\Pi_{\mathcal{B}_{r}}^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)-\mathbb{E}\delta_{n}(\Pi^{k}_{\mathcal{B}_{r}})+\frac{\mathrm{KL}(\Pi_{\mathcal{B}_{r}}^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{k\lambda_{n}C_{k}^{2}}{n}$
$\leq\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\mathrm{KL}(\Pi_{\mathcal{B}_{r}}^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{k\lambda_{n}C_{k}^{2}}{n},$
where the last line follows since $\delta_{n}(\Pi^{k}_{\mathcal{B}_{r}})\geq 0$ for all $r>0$.
Note that $\text{KL}(Q^{k}\|\Pi^{k})=k\cdot\text{KL}(Q\|\Pi)$, and that, for $0\leq r\leq r_{0}$, over $\mathcal{B}_{r}$, $\mathcal{D}_{S}(P_{\theta},P_{0})\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+r$, so that we obtain
$\mathbb{E}\left[\mathcal{D}_{S}(P_{Q_{n}},P_{0})\right]$ $\leq\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\mathrm{KL}(\Pi_{\mathcal{B}_{r}}^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{k\lambda_{n}C_{k}^{2}}{n},$
$=\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\mathrm{KL}(\Pi_{\mathcal{B}_{r}}\|\Pi)}{\lambda_{n}}\right\}+\frac{k\lambda_{n}C_{k}^{2}}{n}$
$=\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\log\left\{1/\Pi(\mathcal{B}_{r})\right\}}{\lambda_{n}}\right\}+\frac{k\lambda_{n}C_{k}^{2}}{n},$
$\leq\inf_{r_{0}\geq r>0}\left\{\int\left\{\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+r\right\}\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{d_{S}\log(c_{S}/r)}{\lambda_{n}}\right\}+\frac{k\lambda_{n}C_{k}^{2}}{n}$
$=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\inf_{r_{0}\geq r>0}\left\{r+\frac{d_{S}\log(c_{S}/r)}{\lambda_{n}}\right\}+\frac{k\lambda_{n}C_{k}^{2}}{n}$
The remainder of the proof follows the same as in the proof of Lemma 5. ∎
Proof of Theorem 4.
Assumption 5 ensures the existence of some $Q^{\star}\in\arg\inf_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{S}\left(P_{Q},P_{0}\right)$ such that
${\mathsf{JenGap}}:=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\mathcal{D}_{S}\left(P_{Q^{\star}},P_{0}\right)-{\mathsf{AppGap}}(Q^{\star},k)>0,$
with ${\mathsf{AppGap}}(Q^{\star},k)\geq 0$ under Assumptions 1-2. According to Assumption 6, there also exists some $\widetilde{Q}^{\star}$ such that $\textnormal{KL}(\widetilde{Q}^{\star}\|\Pi)<+\infty$ and
$\left|\int L(\theta_{1:k})\mathsf{d}Q^{\star}(\theta)-\int L(\theta_{1:k})\mathsf{d}\widetilde{Q}^{\star}(\theta)\right|\leq\frac{\mathsf{JenGap}}{4}\,.$
Denoting
${\mathsf{AppGap}}_{n}(Q,k):=\frac{1}{n}\left\{\widehat{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),x_{i}\right)-S\left(\int P_{\theta}\mathsf{d}Q(\theta),x_{i}\right)\right\}\,$,
we can re-arrange equation (33) in the proof of Lemma 1 (valid due to Assumptions 1-3) as
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$
$\leq\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),x_{i}\right)-S\left(P_{0},x_{i}\right)\right)+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$=\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{\frac{1}{n}\sum_{i=1}^{n}\left(S\left(\int P_{\theta}\mathsf{d}Q(\theta),x_{i}\right)-S\left(P_{0},x_{i}\right)\right)+{\mathsf{AppGap}}_{n}(Q,k)+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\left(S\left(\int P_{\theta}\mathsf{d}Q(\theta),x_{i}\right)-S\left(P_{0},x_{i}\right)\right)+{\mathsf{AppGap}}_{n}(Q,k)+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right]\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$=\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+{\mathsf{AppGap}}(Q,k)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}.$ (34)
Hence, since $\mathrm{KL}(Q^{k}\|\Pi^{k})=k\cdot\mathrm{KL}(Q\|\Pi)$, we can obtain:
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+{\mathsf{AppGap}}(Q,k)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$=\inf_{Q\ll\Pi}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+{\mathsf{AppGap}}(Q,k)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}\widetilde{Q}^{\star}(\theta),P_{0}\right)+{\mathsf{AppGap}}(\widetilde{Q}^{\star},k)+\frac{\textnormal{KL}(\widetilde{Q}^{\star}\|\Pi)}{\lambda_{n}}+\frac{k\lambda_{n}C_{k}^{2}}{n}$
$=\underbrace{\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}\widetilde{Q}^{\star}(\theta),P_{0}\right)+{\mathsf{AppGap}}(\widetilde{Q}^{\star},k)}_{\equiv\int L(\theta_{1:k})\mathsf{d}\widetilde{Q}^{\star}(\theta_{1:k})-S(P_{0},P_{0})}-\underbrace{\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q^{\star}(\theta),P_{0}\right)-{\mathsf{AppGap}}(Q^{\star},k)}_{\equiv\int L(\theta_{1:k})\mathsf{d}{Q}^{\star}(\theta_{1:k})-S(P_{0},P_{0})}$
$\quad+\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q^{\star}(\theta),P_{0}\right)+{\mathsf{AppGap}}({Q}^{\star},k)+\frac{\textnormal{KL}(\widetilde{Q}^{\star}\|\Pi)}{\lambda_{n}}+\frac{k\lambda_{n}C_{k}^{2}}{n}$
$\leq\frac{\mathsf{JenGap}}{4}+\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q^{\star}(\theta),P_{0}\right)+{\mathsf{AppGap}}({Q}^{\star},k)+\frac{\textnormal{KL}(\widetilde{Q}^{\star}\|\Pi)}{\lambda_{n}}+\frac{k\lambda_{n}C_{k}^{2}}{n}$
$=\frac{\mathsf{JenGap}}{4}+\mathcal{D}_{S}\left(P_{Q^{\star}},P_{0}\right)+\left\{\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\mathcal{D}_{S}\left(P_{Q^{\star}},P_{0}\right)-{\mathsf{JenGap}}\right\}+\frac{\textnormal{KL}(\widetilde{Q}^{\star}\|\Pi)}{\lambda_{n}}+\frac{k\lambda_{n}C_{k}^{2}}{n}$
$=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\frac{\textnormal{KL}(\widetilde{Q}^{\star}\|\Pi)}{\lambda_{n}}+\frac{k\lambda_{n}C_{k}^{2}}{n}-\frac{3}{4}\mathsf{JenGap}.$
Hence using Assumption 4, for $n$ large enough so that
$\frac{\textnormal{KL}(\widetilde{Q}^{\star}\|\Pi)}{\lambda_{n}}+\frac{k\lambda_{n}C_{k}^{2}}{n}\leq\frac{\mathsf{JenGap}}{4}\,$,
we have
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\frac{\mathsf{JenGap}}{2}\,.$
Furthermore, from Lemma 5,
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}^{\dagger}(\theta),P_{0}\right)\right]\leq C\frac{\log(n)}{n^{1/2}}+\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})$
so that for $n$ large enough
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}^{\dagger}(\theta),P_{0}\right)\right]\geq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\frac{\mathsf{JenGap}}{4}\,$,
so that for $n$ large enough but finite,
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$ $\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\frac{\mathsf{JenGap}}{2}$
$\leq\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}^{\dagger}(\theta),P_{0}\right)\right]-\frac{\mathsf{JenGap}}{4}$
$<\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}^{\dagger}(\theta),P_{0}\right)\right].$
To obtain the second result, return to equation (34), and use the fact that $Q^{\star}\in\operatornamewithlimits{arginf}_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{S}(P_{Q},P_{0})$ satisfies $\mathrm{KL}(Q^{\star}\|\Pi)<\infty$, to obtain
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+{\mathsf{AppGap}}(Q,k)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\mathcal{D}_{S}(P_{Q^{\star}},P_{0})+{\mathsf{AppGap}}(Q^{\star},k)+\frac{\text{KL}(Q^{\star}\|\Pi)}{\lambda_{n}}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\frac{C}{\lambda_{n}}+\frac{\lambda_{n}kC_{k}^{2}}{n}+{\mathsf{AppGap}}(Q^{\star},k).$
Choosing $\lambda_{n}\propto n^{1/2}/\log(n)$ then yields the stated result. ∎
Proof of Theorem 5.
We break the result up into cases depending on whether (i) or (ii) in Assumption 7 is satisfied.
Assumption 7(i). From equation (34) in the proof of Theorem 4, we have that
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]$ $\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+{\mathsf{AppGap}}(Q,k)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\inf_{Q\ll\Pi}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+{\mathsf{AppGap}}(Q,k)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$=\mathcal{D}_{S}(P_{Q^{\star}},P_{0})+{\mathsf{AppGap}}(Q^{\star},k)+\frac{\text{KL}(Q^{\star}\|\Pi)}{\lambda_{n}}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq{\mathsf{AppGap}}(Q^{\star},k)+C\frac{\log(n)}{n^{1/2}},$
where the last inequality follows from the fact that $\mathrm{KL}(Q^{\star}\|\Pi)<\infty$ and choosing $\lambda_{n}=(n/\log n)^{1/2}$. Since in this case $\mathcal{D}_{S}(P_{Q^{\star}},P_{0})=0$,
${\mathsf{JenGap}}=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-{\mathsf{AppGap}}(Q^{\star},k)>0,$
so that under Assumption 5, there exist an $n$ large enough so that
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]$ $\leq{\mathsf{AppGap}}(Q^{\star},k)-\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+C\frac{\log(n)}{n^{1/2}}+\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})$
$=-{\mathsf{JenGap}}+C\frac{\log(n)}{n^{1/2}}+\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})$
$<-\frac{\mathsf{JenGap}}{4}+\mathcal{D}_{S}(P_{\theta^{\star}},P_{0}).$
Further, for $n$ large enough, from Lemma 5, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}^{\dagger}},P_{0}\right)\right]\geq-\frac{\mathsf{JenGap}}{4}+\mathcal{D}_{S}(P_{\theta^{\star}},P_{0}).$ Putting these inequalities together, we see that, for $n$ large enough,
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]$ $<-\frac{\mathsf{JenGap}}{4}+\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})\leq\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}^{\dagger}},P_{0}\right)\right].$
Assumption 7(ii). Recall $\mathcal{B}_{r,j}=\{\theta:\mathcal{D}_{S}(P_{\theta},P_{\theta_{j}^{\star}})\leq r\}$, denote $\mathsf{d}Q_{r}^{\star}(\theta)=\sum_{j=1}^{K}\omega_{j}\mathsf{d}\pi^{\star}_{r,j}(\theta)$ where $\omega\in\Delta^{K-1}$, and
$\mathsf{d}\pi^{\star}_{r,j}(\theta)=\begin{cases}\frac{\pi(\theta)}{\Pi(\mathcal{B}_{r,j})}&\text{ if }\theta\in\mathcal{B}_{r,j}\\ 0&\text{ else }\end{cases}.$
Using $Q^{\star}_{r}$ within equation (34) in the proof of Theorem 4, yields
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+{\mathsf{AppGap}}(Q,k)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\inf_{r>0}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q^{\star}_{r}(\theta),P_{0}\right)+{\mathsf{AppGap}}(Q^{\star}_{r},k)+\frac{\textnormal{KL}(Q^{\star}_{r}\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}C_{k}^{2}}{n}$
$=\inf_{r>0}\left\{\mathcal{D}_{S}\left(\sum_{j=1}^{K}\omega_{j}\int P_{\theta}\,\mathsf{d}\pi^{\star}_{r,j}(\theta),\sum_{j=1}^{K}\omega_{j}P_{\theta_{j}^{\star}}\right)+\frac{\textnormal{KL}\left(\sum_{j=1}^{K}\omega_{j}\Pi^{\star}_{r,j}\|\Pi\right)}{\lambda_{n}}+{\mathsf{AppGap}}(Q^{\star}_{r},k)\right\}+\frac{\lambda_{n}C_{k}^{2}}{n}$
$\leq\inf_{r>0}\left\{\sum_{j=1}^{K}\omega_{j}\,\int\,\mathcal{D}_{S}\left(P_{\theta},P_{\theta_{j}^{\star}}\right)\mathsf{d}\pi^{\star}_{r,j}(\theta)+\frac{\sum_{j=1}^{K}\omega_{j}\textnormal{KL}\left(\Pi^{\star}_{r,j}\|\Pi\right)}{\lambda_{n}}+{\mathsf{AppGap}}(Q^{\star}_{r},k)\right\}+\frac{\lambda_{n}C_{k}^{2}}{n}$
$\leq\inf_{r>0}\left\{\sum_{j=1}^{K}\omega_{j}\,\int\,\mathcal{D}_{S}\left(P_{\theta},P_{\theta_{j}^{\star}}\right)\mathsf{d}\pi^{\star}_{r,j}(\theta)+\frac{\sum_{j=1}^{K}\omega_{j}\log\{1/\Pi(\mathcal{B}_{r,j})\}}{\lambda_{n}}+{\mathsf{AppGap}}(Q^{\star}_{r},k)\right\}+\frac{\lambda_{n}C_{k}^{2}}{n}$
$\leq\inf_{r>0}\left\{\sum_{j=1}^{K}\omega_{j}\,\int\,\mathcal{D}_{S}\left(P_{\theta},P_{\theta_{j}^{\star}}\right)\mathsf{d}\pi^{\star}_{r,j}(\theta)+\frac{d_{s}\log(c_{s}/r)}{\lambda_{n}}+{\mathsf{AppGap}}(Q^{\star}_{r},k)\right\}+\frac{\lambda_{n}C_{k}^{2}}{n}$
$\leq\inf_{r>0}\left\{r+\frac{d_{s}\log(c_{s}/r)}{\lambda_{n}}\right\}+{\mathsf{AppGap}}(Q^{\star}_{r},k)+\frac{\lambda_{n}C_{k}^{2}}{n}.$
Optimising $r$ using similar arguments to the end of Lemma 5 then yields the stated result. Furthermore, when equation (6) is satisfied with equality, ${\mathsf{AppGap}}(Q^{\star}_{r},k)=0$ for any $r>0$. ∎
Proof of Corollary 3.
The stated claim follows from the results in Theorem 3 and Lemma 5. In particular, from the proof of Theorem 3, we have,
$\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{0})]\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\inf_{r>0}\left\{r+\frac{d_{s}\log(c_{s}/r)}{\lambda_{n}}\right\}+\frac{\lambda_{n}C_{k}^{2}}{n},$
and optimizing the choice of $\lambda_{n}$ as in the proof of Lemma 5 then yields
$0\leq\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq$ $\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+C\frac{\log(n)}{n^{1/2}}$
$\leq$ $C\frac{\log(n)}{n^{1/2}}+\mathbb{E}[\mathcal{D}_{S}(P_{Q^{\dagger}_{n}},P_{0})]+\mathbb{E}\left\{\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\mathcal{D}_{S}(P_{Q^{\dagger}_{n}},P_{0})\right\}$
$\leq$ $C\frac{\log(n)}{n^{1/2}}+\mathbb{E}[\mathcal{D}_{S}(P_{Q^{\dagger}_{n}},P_{0})],$
where the third line follows from Lemma 5. ∎
Proof of Corollary 4 .
Since $\mathsf{d}_{2}(\cdot,\cdot)$ satisfies a weak triangle inequality, for some $C>0$:
                                                                                                    $\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}},P_{Q_{n}^{\dagger}})]$ ${}\leq{}$ $C\mathbb{E}\{\mathsf{d}_{2}(P_{Q_{n}},P_{0})+\mathsf{d}_{2}(P_{Q_{n}^{\dagger}},P_{0})\}.$                                                                                                      (35)
Recall that, by Theorem 5, $\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{0})]\leq C\log(n)/n^{1/2}$, while from Lemma 5, we also have $\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]\leq C\log(n)/n^{1/2}$. Hence, from Assumption 8(iii), and concavity of the map $x\mapsto x^{1/\alpha}$ for $\alpha\geq 1$,
$\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}},P_{0})]\leq$ $\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{0})^{1/\alpha}]\leq\left\{\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{0})]\right\}^{1/\alpha}\leq C\{\log(n)/n^{1/2}\}^{1/\alpha},$
$\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}^{\dagger}},P_{0})]\leq$ $\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})^{1/\alpha}]\leq\left\{\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]\right\}^{1/\alpha}\leq\{\log(n)/n^{1/2}\}^{1/\alpha}.$
where the first bound comes from Theorem 5, and the second from Lemma 5. Applying the above bounds into (C), we obtain
                                                                                                    $\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}},P_{Q_{n}^{\dagger}})]$ ${}\leq{}$ $C\{\log(n)/n^{1/2}\}^{1/\alpha}.$                                                                                                      (36)
From equation (C), and Assumption 8(i) - identifiability and continuity of $\theta\mapsto P_{\theta}$ - conclude that $P_{Q_{n}}\rightarrow P_{0}$ and $P_{Q_{n}^{\dagger}}\rightarrow P_{0}$. Hence, by Assumption 8(ii) there exists $N_{1}$ large enough so that, for $n\geq N_{1}$,
$\mathbb{E}[\mathsf{d}_{1}(Q_{n},Q_{n}^{\dagger})]\leq C\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}},P_{Q_{n}^{\dagger}})].$
Since $Q_{n}^{\dagger}\rightarrow\updelta_{\theta^{\star}}$, and $P_{0}=\int P_{\theta}\mathsf{d}\updelta_{\theta^{\star}}$, there exists an $N_{2}$ large enough so that for all $n\geq N_{2}$,
$\mathbb{E}[\mathsf{d}_{1}(Q_{n}^{\dagger},\updelta_{\theta^{\star}})]\leq C\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}^{\dagger}},P_{0})].$
For $n\geq\max\{N_{1},N_{2}\}$, apply the triangle inequality and equation (C) to obtain
$\mathbb{E}[\mathsf{d}_{1}(Q_{n},\updelta_{\theta^{\star}})]\leq\mathbb{E}[\mathsf{d}_{1}(Q_{n},Q_{n}^{\dagger})]+\mathbb{E}[\mathsf{d}_{1}(Q_{n}^{\dagger},\updelta_{\theta^{\star}})]$ $\leq C\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}},P_{Q_{n}^{\dagger}})]+C\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}^{\dagger}},P_{0})]$
$\leq C\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}},P_{0})]+2C\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}^{\dagger}},P_{0})]$
$\leq C(\log n/n^{1/2})^{1/\alpha}.$
Proof of Theorem 6.
The proof exploits the following upper bound from Lemma 1:
$\mathbb{E}$ $\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\widehat{S}_{k}\left(P_{Q},P_{0}\right)-\mathcal{S}(P_{0},P_{0})+\frac{\textnormal{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\widehat{S}_{k}\left(P_{Q},P_{0}\right)-\mathcal{S}\left(P_{Q},P_{0}\right)+\mathcal{S}\left(P_{Q},P_{0}\right)+\frac{\textnormal{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\widehat{S}_{k}\left(P_{Q^{\star}},P_{0}\right)-\mathcal{S}\left(P_{Q^{\star}},P_{0}\right)+\frac{\textnormal{KL}(Q^{\star}\|\Pi)}{\lambda_{n}}+\frac{\lambda_{n}kC_{k}^{2}}{n}+\mathrm{KL}\left(P_{0}\big\|P_{Q^{\star}}\right).$
In the following, let us abuse notation and set $Q^{\star}(\theta_{1:k})=\bigotimes_{j=1}^{K}Q^{\star}(\theta)$. Consider
$\widehat{S}_{k}\left(P_{Q^{\star}},P_{0}\right)-S\left(P_{Q^{\star}},P_{0}\right)$
$\overset{(0)}{=}\mathbb{E}_{x\sim P_{0}}\left[\int-\log\left(\frac{1}{k}\sum_{j=1}^{k}p_{\theta_{j}}(x)\right)\mathsf{d}Q^{\star}(\theta_{1:k})\right]-\mathbb{E}_{x\sim P_{0}}\left[-\log\left(\int p_{\theta}(x)\,\mathsf{d}Q^{\star}(\theta)\right)\right]$
$\overset{(1)}{=}\mathbb{E}_{x\sim P_{0}}\left[\mathop{\text{\large$\int$}}\nolimits-\log\left(1+\frac{\frac{1}{k}\sum_{j=1}^{k}p_{\theta_{j}}(x)-\int p_{\theta}(x)\mathsf{d}Q^{\star}(\theta)}{\int p_{\theta}(x)\mathsf{d}Q^{\star}(\theta)}\right)\mathsf{d}Q^{\star}(\theta_{1:k})\right]$
$\overset{(2)}{=}\mathbb{E}_{x\sim P_{0}}\left[\mathop{\text{\large$\int$}}\nolimits\left\{\frac{\frac{1}{k}\sum_{j=1}^{k}p_{\theta_{j}}(x)-\int p_{\theta}(x)\mathsf{d}Q^{\star}(\theta)}{\int p_{\theta}(x)\mathsf{d}Q^{\star}(\theta)}-\log\left(1+\frac{\frac{1}{k}\sum_{j=1}^{k}p_{\theta_{j}}(x)-\int p_{\theta}(x)\mathsf{d}Q^{\star}(\theta)}{\int p_{\theta}(x)\mathsf{d}Q^{\star}(\theta)}\right)\right\}\mathsf{d}Q^{\star}(\theta_{1:k})\right]$
$\overset{(3)}{\leq}\mathbb{E}_{x\sim P_{0}}\left[\mathop{\text{\large$\int$}}\nolimits\left|\frac{\frac{1}{k}\sum_{j=1}^{k}p_{\theta_{j}}(x)-\int p_{\theta}(x)\mathsf{d}Q^{\star}(\theta)}{\int p_{\theta}(x)\mathsf{d}Q^{\star}(\theta)}\right|\cdot\left|\log\left(1+\frac{\frac{1}{k}\sum_{j=1}^{k}p_{\theta_{j}}(x)-\int p_{\theta}(x)\mathsf{d}Q^{\star}(\theta)}{\int p_{\theta}(x)\mathsf{d}Q^{\star}(\theta)}\right)\right|\mathsf{d}Q^{\star}(\theta_{1:k})\right]$
$\overset{(4)}{\leq}\mathbb{E}_{x\sim P_{0}}\bigg[\sqrt{\mathop{\text{\large$\int$}}\nolimits\left(\frac{\frac{1}{k}\sum_{j=1}^{k}p_{\theta_{j}}(x)-\int p_{\theta}(x)\mathsf{d}Q^{\star}(\theta)}{\int p_{\theta}(x)\mathsf{d}Q^{\star}(\theta)}\right)^{2}\mathsf{d}Q^{\star}(\theta_{1:k})}$
$\cdot\sqrt{\mathop{\text{\large$\int$}}\nolimits\log\left(1+\frac{\frac{1}{k}\sum_{j=1}^{k}p_{\theta_{j}}(x)-\int p_{\theta}(x)\mathsf{d}Q^{\star}(\theta)}{\int p_{\theta}(x)\mathsf{d}Q^{\star}(\theta)}\right)^{2}\mathsf{d}Q^{\star}(\theta_{1:k})}\bigg]$
$\overset{(5)}{=}\mathbb{E}_{x\sim P_{0}}\left[\sqrt{\mathbb{V}_{\theta_{1:k}\sim Q^{\star}(\theta_{1:k})}\left[\frac{1}{k}\sum_{j=1}^{k}\frac{p_{\theta_{j}}(x)}{\mathbb{E}_{\theta\sim Q^{\star}}\left[p_{\theta}(x)\right]}\right]}\cdot\sqrt{\mathbb{E}_{\theta_{1:k}\sim Q^{\star}(\theta_{1:k})}\left[\log\left(\frac{1}{k}\sum_{j=1}^{k}\frac{p_{\theta_{j}}(x)}{\mathbb{E}_{\theta\sim Q^{\star}}\left[p_{\theta}(x)\right]}\right)^{2}\right]}\right]$
$\overset{(6)}{=}\mathbb{E}_{x\sim P_{0}}\left[\sqrt{\frac{1}{k}\cdot\mathbb{V}_{\theta\sim Q^{\star}}\left[\frac{p_{\theta}(x)}{\mathbb{E}_{\theta\sim Q^{\star}}\left[p_{\theta}(x)\right]}\right]}\cdot\sqrt{\mathbb{E}_{\theta_{1:k}\sim Q^{\star}(\theta_{1:k})}\left[\log\left(\frac{1}{k}\sum_{j=1}^{k}\frac{p_{\theta_{j}}(x)}{\mathbb{E}_{\theta\sim Q^{\star}}\left[p_{\theta}(x)\right]}\right)^{2}\right]}\right]$
$\overset{(7)}{\leq}\mathbb{E}_{x\sim P_{0}}\bigg[\sqrt{\frac{1}{k}\cdot\mathbb{V}_{\theta\sim Q^{\star}}\left[\frac{p_{\theta}(x)}{\mathbb{E}_{\theta\sim Q^{\star}}\left[p_{\theta}(x)\right]}\right]}\times$
$\cdot\frac{M^{\prime}}{\sqrt{k}}\cdot\mathbb{E}_{\theta\sim Q^{\star}}\left[\left(\frac{p_{\theta}(x)}{\mathbb{E}_{\theta\sim Q^{\star}}\left[p_{\theta}(x)\right]}\right)^{4}\right]^{1/4}\cdot\left(\mathbb{E}_{\theta\sim Q^{\star}}\left[\log\left(\frac{p_{\theta}(x)}{\mathbb{E}_{\theta\sim Q^{\star}}\left[p_{\theta}(x)\right]}\right)^{4}\right]^{1/4}+1\right)\bigg]$
$\overset{(8)}{\leq}\frac{M^{\prime}}{k}\cdot\mathbb{E}_{x\sim P_{0}}\left[\mathbb{E}_{\theta\sim Q^{\star}}\left[\left(\frac{p_{\theta}(x)}{\mathbb{E}_{\theta\sim Q^{\star}}\left[p_{\theta}(x)\right]}\right)^{4}\right]^{1/2}\cdot\left(\mathbb{E}_{\theta\sim Q^{\star}}\left[\log\left(\frac{p_{\theta}(x)}{\mathbb{E}_{\theta\sim Q^{\star}}\left[p_{\theta}(x)\right]}\right)^{4}\right]^{1/4}+1\right)\right].$
Lines $(0)$ and $(1)$ are basic algebra; $(2)$ comes from the fact that the first term in the fourth line is centered with respect to $\theta_{1:k}$; $(3)$ is a direct application of Equation (31) Lemma 5 in Cherief-Abdellatif et al. (2025); $(4)$ is a consequence of Cauchy’s inequality; $(5)$ follows from rewriting integrals into expectations and variances; $(6)$ writes the variance of an empirical mean of independent random variables (conditional on $x$) as a single variance divided by the number of random variables; $(7)$ is an application of Equation (ii) Lemma 6 in Cherief-Abdellatif et al. (2025); and finally $(9)$ upper bounds a variance by the square root of a fourth moment.
So we have
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$ $\leq\widehat{S}_{k}\left(\int P_{\theta}\,\mathsf{d}Q^{\star}(\theta),P_{0}\right)-S\left(\int P_{\theta}\,\mathsf{d}Q^{\star}(\theta),P_{0}\right)+\frac{\textnormal{KL}(Q^{\star}\|\Pi)}{\lambda_{n}}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$+\mathrm{KL}\left(P_{0}\big\|\int P_{\theta}\mathsf{d}Q^{\star}(\theta)\right)$
$\leq\frac{M_{2}}{k}+\frac{\textnormal{KL}(Q^{\star}\|\Pi)}{\lambda_{n}}+\frac{\lambda_{n}kC_{k}^{2}}{n}+\mathrm{KL}\left(P_{0}\big\|\int P_{\theta}\mathsf{d}Q^{\star}(\theta)\right),$
with
$M_{2}=M^{\prime}\cdot\mathbb{E}_{x\sim P_{0}}\left[\mathbb{E}_{\theta\sim Q^{\star}}\left[\left(\frac{p_{\theta}(x)}{\mathbb{E}_{\theta\sim Q^{\star}}\left[p_{\theta}(x)\right]}\right)^{4}\right]^{1/2}\cdot\left(\mathbb{E}_{\theta\sim Q^{\star}}\left[\log\left(\frac{p_{\theta}(x)}{\mathbb{E}_{\theta\sim Q^{\star}}\left[p_{\theta}(x)\right]}\right)^{4}\right]^{1/4}+1\right)\right].$
So the choices $\lambda_{n}\propto(n/kC_{k}^{2})^{1/2}$ and $kC_{k}\propto n^{1/2}$ lead to
                                                                                                    $\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]\leq M\sqrt{\frac{kC_{k}^{2}}{n}}+\mathrm{KL}\left(P_{0}\big\|\int P_{\theta}\mathsf{d}Q^{\star}(\theta)\right),$                                                                                                      (37)
for some constant $M>0$, which ends the proof. ∎
Proof of Theorem 7 .
The only change from the proof of Theorem 6 here consists in establishing that Assumption 3 holds with $C_{k}^{2}\propto C^{2}\log k$ under Assumption 10. The rest of the proof directly follows from Equation (C).
We now show Assumption 3 holds with $C_{k}^{2}\propto C^{2}\log k$ under Assumption 10. Our proof relies on three key observations:
  1. 1.
    Everywhere vs. Joint sub-Gaussianity. As discussed in Section 7.2, the everywhere counterpart of Assumption 3
    $\forall(\theta_{1:k})\in\Theta^{k},\quad\mathbb{E}\left[e^{\lambda(\mathcal{L}(\theta_{1:k})-L_{n}(\theta_{1:k}))}\right]\leq e^{\lambda^{2}C_{k}^{2}/n},$
    is stronger than the joint version we use in the paper, so we directly derive the everywhere version in this proof.
  2. 2.
    Sub-Gaussianity and Concentration. Assumption 3 is essentially a sub-Gaussianity condition. The factor $1/n$ in the bound arises because we consider empirical averages:
    $L_{n}(\theta_{1:k})=\frac{1}{n}\sum_{i=1}^{n}L(\theta_{1:k},x_{i}).$
    Therefore, it suffices to show the sub-Gaussian bound for a single variable:
    $\mathbb{E}_{X}\left[e^{\lambda(\mathcal{L}(\theta_{1:k})-L(\theta_{1:k},X))}\right]\leq e^{\lambda^{2}C_{k}^{2}}.$
  3. 3.
    Alternative Definition of Sub-Gaussianity. In the proof, we use the following characterization of sub-Gaussianity:
    $\mathbb{E}\left[\exp\left(\frac{(Z-\mathbb{E}[Z])^{2}}{\mathcal{C}_{k}^{2}}\right)\right]\leq 2,$
    where the minimal $\mathcal{C}_{k}^{2}$ (the sub-Gaussian norm) is proportional to the minimal $C_{k}^{2}$ used in Assumption 3 (the sub-Gaussian parameter), see e.g. Leskelä and Zhukov (2025).
We can now state the result.
Theorem 8.
Assume that Assumption 10 is satisfied, that is, for any single $\theta\in\Theta$, the log-likelihood is sub-Gaussian: for any $\lambda\in\mathbb{R}$,
$\mathbb{E}_{X}\left[\exp\left(\lambda\Big\{\log p_{\theta}(X)-\mathbb{E}_{X}[\log p_{\theta}(X)]\Big\}\right)\right]\leq e^{\lambda^{2}C^{2}}.$
Then, for any $(\theta_{1:k})\in\Theta^{k}$, there exists a constant $\gamma>0$ independent of $(\theta_{1:k})$ such that, for all $\lambda\in\mathbb{R}$,
$\mathbb{E}_{X}\left[\exp\left(\lambda\Bigg\{\log\frac{1}{k}\sum_{j=1}^{k}p_{\theta_{j}}(X)-\mathbb{E}_{X}\Big[\log\frac{1}{k}\sum_{j=1}^{k}p_{\theta_{j}}(X)\Big]\Bigg\}\right)\right]\leq e^{\lambda^{2}\gamma C^{2}\log k}.$
In other words, $C_{k}^{2}\propto C^{2}\log k$ for the log-score.
Proof.
We look for a constant $\mathcal{C}_{k}>0$ such that
$\mathbb{E}_{X}\exp\Bigg(\frac{\Big(\log\big(\frac{1}{k}\sum_{j=1}^{k}p_{\theta_{j}}(X)\big)-\mathbb{E}_{X}\big[\log\big(\frac{1}{k}\sum_{j=1}^{k}p_{\theta_{j}}(X)\big)\big]\Big)^{2}}{\mathcal{C}_{k}^{2}}\Bigg)\leq 2.$ (38)
Once (38) is established, we immediately obtain that there exists $C_{k}^{2}=\gamma\mathcal{C}_{k}^{2}$ such that
$\mathbb{E}_{X}\Big[\exp\big(\lambda\big\{\log(\frac{1}{k}\sum_{j=1}^{k}p_{\theta_{j}}(X))-\mathbb{E}_{X}[\log(\frac{1}{k}\sum_{j=1}^{k}p_{\theta_{j}}(X))]\big\}\big)\Big]\leq e^{\lambda^{2}\gamma C_{k}^{2}}$
for some $\gamma>0$, by the equivalence between sub-Gaussian norms and parameters.
Step 1: Rewrite using the log-sum-exp function. For any $(\theta_{1},\dots,\theta_{k})\in\Theta^{k}$, almost surely:
$\log\frac{1}{k}\sum_{j=1}^{k}p_{\theta_{j}}(X)-\mathbb{E}_{X}\big[\log\frac{1}{k}\sum_{j=1}^{k}p_{\theta_{j}}(X)\big]=\text{LSE}(\{\log p_{\theta_{j}}(X)\}_{1\leq j\leq k})-\mathbb{E}_{X}[\text{LSE}(\{\log p_{\theta_{j}}(X)\}_{1\leq j\leq k})],$
where $\text{LSE}(z_{1:k})=\log\sum_{j=1}^{k}e^{z_{j}}$.
Step 2: Apply Jensen and Lipschitz properties. For any $\mathcal{C}_{k}>0$, we can write
$\mathbb{E}_{X}\exp\Bigg(\frac{\big(\text{LSE}(\{\log p_{\theta_{j}}(X)\})-\mathbb{E}_{X}[\text{LSE}(\{\log p_{\theta_{j}}(X)\})]\big)^{2}}{\mathcal{C}_{k}^{2}}\Bigg)$
$=\mathbb{E}_{X}\exp\Bigg(\frac{\big(\text{LSE}(\{\log p_{\theta_{j}}(X)\})-\mathbb{E}_{Y}[\text{LSE}(\{\log p_{\theta_{j}}(Y)\})]\big)^{2}}{\mathcal{C}_{k}^{2}}\Bigg)$
$\leq\mathbb{E}_{X,Y}\exp\Bigg(\frac{\big(\text{LSE}(\{\log p_{\theta_{j}}(X)\})-\text{LSE}(\{\log p_{\theta_{j}}(Y)\})\big)^{2}}{\mathcal{C}_{k}^{2}}\Bigg)$
$\leq\mathbb{E}_{X,Y}\exp\Bigg(\frac{\big\|\{\log p_{\theta_{j}}(X)\}-\{\log p_{\theta_{j}}(Y)\}\big\|_{\infty}^{2}}{\mathcal{C}_{k}^{2}}\Bigg)$
$=\mathbb{E}_{X,Y}\exp\Bigg(\frac{\sup_{1\leq j\leq k}\big((\log p_{\theta_{j}}(X)-\mathbb{E}_{X}[\log p_{\theta_{j}}(X)])-(\log p_{\theta_{j}}(Y)-\mathbb{E}_{Y}[\log p_{\theta_{j}}(Y)])\big)^{2}}{\mathcal{C}_{k}^{2}}\Bigg).$
Step 3: Apply the inequality $(a-b)^{2}\leq 2a^{2}+2b^{2}$ and Jensen.
$\leq\mathbb{E}_{X,Y}\exp\left(\frac{\sup_{1\leq j\leq k}\left\{2\left(\log p_{\theta_{j}}(X)-\mathbb{E}_{X}[\log p_{\theta_{j}}(X)]\right)^{2}+2\left(\log p_{\theta_{j}}(Y)-\mathbb{E}_{Y}[\log p_{\theta_{j}}(Y)]\right)^{2}\right\}}{\mathcal{C}_{k}^{2}}\right)$
$\leq\mathbb{E}_{X,Y}\exp\left(\frac{2\sup_{1\leq j\leq k}\left(\log p_{\theta_{j}}(X)-\mathbb{E}_{X}[\log p_{\theta_{j}}(X)]\right)^{2}+2\sup_{1\leq j\leq k}\left(\log p_{\theta_{j}}(Y)-\mathbb{E}_{Y}[\log p_{\theta_{j}}(Y)]\right)^{2}}{\mathcal{C}_{k}^{2}}\right)$
$=\mathbb{E}_{X}\left[\exp\left(\frac{2\sup_{1\leq j\leq k}\left(\log p_{\theta_{j}}(X)-\mathbb{E}_{X}[\log p_{\theta_{j}}(X)]\right)^{2}}{\mathcal{C}_{k}^{2}}\right)\right]^{2}$
$\leq\mathbb{E}_{X}\left[\exp\left(\frac{2\sup_{1\leq j\leq k}\left(\log p_{\theta_{j}}(X)-\mathbb{E}_{X}[\log p_{\theta_{j}}(X)]\right)^{2}}{\mathcal{C}_{k}^{2}}\right)^{2}\right]$
$=\mathbb{E}_{X}\left[\exp\left(\frac{4\sup_{1\leq j\leq k}\left(\log p_{\theta_{j}}(X)-\mathbb{E}_{X}[\log p_{\theta_{j}}(X)]\right)^{2}}{\mathcal{C}_{k}^{2}}\right)\right]$
$=\mathbb{E}_{X}\left[\exp\left(\frac{\sup_{1\leq j\leq k}\left(\log p_{\theta_{j}}(X)-\mathbb{E}_{X}[\log p_{\theta_{j}}(X)]\right)^{2}}{(\mathcal{C}_{k}/2)^{2}}\right)\right]\,.$
Step 4: Bound the supremum via standard sub-Gaussian inequalities. By definition, the last term is bounded by $2$ if $(\mathcal{C}_{k}/2)^{2}$ is taken as the squared sub-Gaussian norm of $\sup_{j}(\log p_{\theta_{j}}(X)-\mathbb{E}_{X}[\log p_{\theta_{j}}(X)])$, which is itself upper bounded (up to a universal factor) by the maximum of the squared sub-Gaussian norms of the individual $\log p_{\theta_{j}}(X)$ multiplied by $\log k$ (see, e.g., Lemma 2.2.2 in Van Der Vaart and Wellner (1996) applied to $\psi(x)=e^{x^{2}}-1$).
Step 5: Conclude. Since each $\log p_{\theta_{j}}(X)-\mathbb{E}_{X}[\log p_{\theta_{j}}(X)]$ has sub-Gaussian parameter $C^{2}$, it also has squared sub-Gaussian norm $C^{2}$ up to a factor, and we finally have
$\mathcal{C}_{k}^{2}\leq\gamma^{\prime}\,C^{2}\log k,$
for some universal constant $\gamma^{\prime}>0$. This establishes (38) and thus proves the theorem. ∎
Appendix D Further experimental details
D.1 Regression with the MMD
sec:synthetic-regression-supp Denote $d=\operatorname{dim}(\Theta)$. In Section 5.2 we use the squared MMD loss to fit the linear regressions, $\{\mathsf{N}(\cdot;\,x^{\top}\theta,\sigma^{2}):\,x\in\mathbb{R}^{d},\,\theta\in\mathbb{R}^{d}\}$, where $\sigma^{2}>0$ is treated as fixed and is estimated from data:
$\hat{\sigma}^{2}=\sum_{i=1}^{n}\frac{(y_{i}-\hat{y}_{i})}{n-d-1};\quad\hat{y}=X\hat{\beta};\quad\hat{\beta}=(X^{\top}X)^{-1}X^{\top}y.$
In the case of the Gibbs posterior, we use the efficient estimator proposed by Alquier and Gerber (2024, Equation 6):
$S_{n}(P_{\theta})=\sum_{i=1}^{n}\ell^{\dagger}(\theta;\,X_{i},y_{i})$
where for all $\theta\in\Theta,\,x\in\mathbb{R}^{d},$ and $y\in\mathbb{R}$,
$\ell^{\dagger}(\theta;\,X,y)=\mathbb{E}_{Y,Y^{\prime}\sim P_{\theta}(\cdot\mid x)}\{\kappa(Y,Y^{\prime})-2\kappa(Y,y)\}.$
In practice we estimate both of these expectations with $m$ Monte Carlo draws:
$\ell^{\dagger}(\theta;\,X,y)\approx\frac{1}{m(m-1)}\sum_{i\neq j}\sum_{j=1}^{m}\kappa(Y_{i},Y^{\prime}_{j})-\frac{2}{m}\kappa(Y_{i},y),$
for $Y_{i},Y^{\prime}_{i}\sim P_{\theta}(\cdot\mid x).$
In the case of the PrO-posterior, we instead use the tandem loss
$L_{n}(\theta,\theta^{\prime})=\sum_{i=1}^{n}\ell(\theta,\theta^{\prime};\,X_{i},y_{i})$
where now
$\ell(\theta,\theta^{\prime};\,X,y)=\mathbb{E}_{\begin{subarray}{c}Y\sim P_{\theta}(\cdot\mid x)\\ Y^{\prime}\sim P_{\theta^{\prime}}(\cdot\mid x)\end{subarray}}\{\kappa(Y,Y^{\prime})-\kappa(Y,y)-\kappa(Y^{\prime},y)\}.$
And again, each of these expectations can be approximated by Monte Carlo:
$\ell(\theta,\theta^{\prime};\,X,y)\approx\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{j=1}^{m}\kappa(Y_{i},Y^{\prime}_{j})-\frac{1}{m}\kappa(Y_{i},y)-\frac{1}{m}\kappa(Y_{i}^{\prime},y),$
for $Y_{i}\sim P_{\theta}(\cdot\mid x)$ and $Y^{\prime}_{i}\sim P_{\theta^{\prime}}(\cdot\mid x)$.
In both cases, we use a Gaussian kernel where the length scale is chosen according to the median heuristic (Garreau et al., 2018).
D.2 Palmer penguins example
The data containing $n=342$ measurements of the penguins’ bill lengths and depths in millimetres were imported from the palmerpenguins package in Python, and jointly centred to have zero-mean and scaled to have unit variance. To these data we fit a Gaussian kernel density estimator implemented in sklearn (Pedregosa et al., 2011), where the bandwidth was chosen with cross-validation across a logarithmic grid. We also fit for comparison a PrO posterior over the mean of an the isotropic bivariate Gaussian location model class, $\mathcal{M}_{\Theta}=\{\mathsf{N}(\theta,\sigma^{2}I):\,\theta\in\mathbb{R}^{2}\}$, where $\sigma^{2}$ is fixed, and which we fit with the squared MMD score. This same bandwidth was used for fitting the PrO posterior with the squared MMD as fitting the KDE. We estimated the squared MMD with Monte Carlo using $m=n$ samples (in the notation of Section D.1), $\lambda_{n}=10^{3}$, and $\sigma^{2}=0.2$. The Wasserstein gradient flow was run with $2^{5}$ particles and $\mathsf{d}t=10^{-1}$ for $10,000$ iterations. In this example, and in all following examples, both Gibbs and PrO posteriors are fit with the same value of $\lambda_{n}$.
D.3 Normal location illustrations
We simulate $n=1,000$ data from each of the data-generating processes:
                                                                                                    $y_{i}$ ${}\sim{}$ $\mathsf{N}(0,1^{2})$ (“Well-specified”)
                                                                                                    $y_{i}$ ${}\sim{}$ $\begin{cases}\mathsf{N}(-2,1^{2}),&\quad\text{with probability }0.2\\ \mathsf{N}(2,1^{2}),&\quad\text{with probability }0.8\end{cases}$ (“Mixture”)
[…]
                                                                                                    $y_{i}$ ${}\sim{}$ $\mathsf{Student}-t_{1.5}$ $\quad\text{(``Heavy-tails'')}.$
For each of these, we estimate the MMD with $m=2^{7}$ Monte Carlo samples, run the Wasserstein gradient flow for $p=2^{5}$ particles for all data, and for $4,000$ iterations similarly to MALA. We use $\lambda_{n}=n$. The values of $\mathsf{d}t$ for the WGF and MALA varied between datasets and were chosen by visual inspection of the trajectory plots.
Figure 8: Illustrative examples of Section 1.2. In the left-most column we show the particle trajectories from a Wasserstein gradient flow targeting the PrO-posterior (in blue) compared to the path of a single MALA chain targeting the Gibbs posterior (in red). In the two axes on the right we show the empirical posterior densities from these two trajectories.
As mentioned before, the predictive distribution in the case of the normal location model is nothing more than a smoothed posterior distribution. We show the posterior distributions of the Gibbs and PrO-posteriors and the Wasserstein gradient flow trajectories in Figure 8.
D.4 Golf putting example
We set $\lambda_{n}=n^{1/2}$ and approximate the log score with a DI approximation as described by Masegosa (2020). The WGF was run with different step sizes for both dimensions, $\mathsf{d}t=(8\times 10^{-4},2\times 10^{-5})$, and with $p=10$ particles for $8,000$ iterations. MALA was run with $\mathsf{d}t=3\times 10^{-7}$ and also for $8,000$ iterations. The initial particle values for both methods were initialised around the posterior mean values (2.23,0.26)(2.23,-0.26) identified by Gelman and Nolan (2002).
D.5 Binary classification example
We simulated $n=1,000$ in the manner described in Section 5.1, and fit all models with $\lambda_{n}=n^{1/2}$. The PrO posterior was based on the log score, and fit under two different approximations: the same DI approximation as in Section 5.3, and an MS approximation with $k=2$. Both were run with $p=2^{6}$ particles, $\mathsf{d}t=5\times 10^{-2}$, and for $10,000$ iterations. MALA was run for $4,000$ steps with $\mathsf{d}t=10^{-4}$.
D.6 Synthetic linear regression
In Section 5.2 we set $\xi=2,\,\sigma^{2}=1,$ and $\eta=0.5$ to generate $n=500$ training data. The log point-wise predictive density (lppd) is estimated from $S$ samples from the posterior $\{\theta^{(s)}\}_{s=1}^{S}$ at a point $\{X_{i},y_{i}\}$ through
$\operatorname{lppd}_{i}=\log\left\{\frac{1}{S}\sum_{s=1}^{S}P_{\theta^{(s)}}(y_{i}\mid X_{i})\right\}.$
In the right-most panel of Figure 6 we plot the difference $\Delta_{i}=\operatorname{lppd}_{i}^{(\operatorname{PrO})}-\operatorname{lppd}_{i}^{(\operatorname{Gibbs})}$. And since we want to maximise the log score, if $\Delta_{i}>0$ we say that the PrO posterior induces better predictive performance. The expected log predictive density (elpd) on a test set, $\{X_{i},y_{i}\}_{i=1}^{n_{\mathrm{test}}}$ is computed as
$\operatorname{elpd}=\sum_{i=1}^{n_{\mathrm{test}}}\operatorname{lppd}_{i}.$
The PrO posterior attains an elpd of 1834.68-1834.68 on $n_{\mathsf{test}}=1,000$ test data, while the Gibbs posterior performs much worse with an elpd of 3593.67-3593.67 on the same data.
The MMD was estimated using $m=2^{5}$ Monte Carlo samples, and the Wasserstein gradient flow was run with $\mathsf{d}t=5\times 10^{-3}$ using $p=2^{4}$ particles and $\lambda_{n}=n^{1/2}$ for $4,000$ iterations, while MALA was run with $\mathsf{d}t=10^{-4}$ also for $4,000$ iterations after $96,000$ warmup iterations to ensure convergence.