### v2

# Predictively Oriented Posteriors

Abstract
We advocate for a new statistical principle that combines the most desirable aspects of both parameter inference and density estimation. This leads us to the predictively oriented (PrO) posterior, which expresses uncertainty as a consequence of predictive ability. Doing so leads to inferences which predictively dominate both classical and generalised Bayes posterior predictive distributions: up to logarithmic factors, PrO posteriors converge to the predictively optimal model average. Whereas classical and generalised Bayes posteriors only achieve this rate if the model can recover the data-generating process, PrO posteriors adapt to the level of model misspecification. This means that they concentrate around the true model in the same way as Bayes and Gibbs posteriors if the model can recover the data-generating distribution, but do not concentrate in the presence of non-trivial forms of model misspecification. Instead, they stabilise towards a predictively optimal posterior whose degree of irreducible uncertainty admits an interpretation as the degree of model misspecification—a sharp contrast to how Bayesian uncertainty and its existing extensions behave. Lastly, we show that PrO posteriors can be sampled from by evolving particles based on mean field Langevin dynamics, and verify the practical significance of our theoretical developments on a number of numerical examples.
[…]
Often motivated through its favourable decision-theoretic properties, the primary goal of Bayesian statistical inference is to quantify uncertainty about a parameter of interest given a set of prior beliefs (e.g. Robert et al., 2007). All other inferential tasks are derivatives of this primary directive. In the current era of black-box prediction models and the wide-spread adoption of machine learning techniques across science and society, this principle leads to a fundamental tension with Bayesian statistical methods: in many modern problems, prediction is the primary goal of interest, with parameter inference being only a secondary concern. Because much of Bayesian methodology is centered on parameter inference, its predictive performance quickly falters outside of the narrow conditions that make Bayesian belief updates optimal (see Zellner, 1988; Aitchison, 1975). In other words, the practice of Bayesian statistics is overwhelmingly focused on parameter inference, not prediction.
One of the clearest ways in which this manifests is posterior collapse: as data accumulate, the Bayes posterior generally concentrates around a single parameter value (see e.g van der Vaart, 1998, Section 10.2). This is sensible for parameter inference when
uncertainty in a parameter of interest arises due to missing observations. If a complete population is observed, then the parameter of interest can be assumed to be known precisely.
(Fong et al., 2023)
In the context of prediction however, posterior collapse is generally not desirable: it implies that as the sample size increases, marginalising over the Bayes posterior ceases to function as a genuine combination of individual predictive distributions. Instead of averaging over models, such predictive distributions essentially predict from a single model. Crucially, this behaviour persists even for generalised Bayesian methods that replace a likelihood-based approach with a loss-based belief update (Bissiri et al., 2016; Jiang and Tanner, 2008; Zhang, 2006; Knoblauch et al., 2022; Alquier, 2024; Syring and Martin, 2019; Martin and Syring, 2022). As a result, the posterior predictives obtained from both classical and generalised Bayes approaches quickly resemble those obtained from a point estimator as sample size grows (McLatchie et al., 2025).
To better understand the impact of this phenomenon on prediction, it is useful to distinguish between two regimes. When a statistical model is well-specified so that the data-generating process is recoverable by a singular parameter value, posterior concentration is the desired outcome: as more data accumulates, uncertainty contracts around the true parameter, and the posterior predictive distribution is well-calibrated. In practice however, statistical analysis is almost always subject to some degree of model misspecification. In this regime, no single parameter value is capable of describing the data generating process, so that the posterior predictives based on standard or generalised Bayesian posteriors become poorly calibrated (Syring and Martin, 2019). Importantly, this issue is not caused by obtaining predictive distributions through the marginalisation of a posterior, but by marginalising over a posterior which collapses to a point mass. This in turn provides strong motivation for methods whose parameter uncertainty can be guaranteed not to collapse to a point mass when this
is appropriate and desirable, since […] some uncertainty always remains about the true (idealised) distribution, no matter how much data are observed.
(Miller and Dunson, 2019)
In this paper, we synthesise the above discussion into a new statistical principle for inference that results in a predictively oriented (PrO) posterior. The resulting approach is distinct from existing inference frameworks in two main ways: first, it treats prediction as its main goal without requiring a black box approach, while naturally inducing uncertainty quantification for a parameter of interest as a secondary inferential outcome; second, it is adaptive: the PrO posterior collapses to a point mass only in the appropriate setting where the model is well-specified. When the model is misspecified however, PrO posteriors acknowledge that no single parameter is sufficient for capturing the true data-generating process, and yield an irreducible level of uncertainty, even with increasing sample size.
Conceptually, PrO posteriors represent a useful intermediary between existing statistical paradigms and compromise between the two tenants of data modelling and algorithmic modelling sub-cultures in statistics that were famously identified by Breiman (2001). In practical terms, this means that PrO posteriors reconcile parameter inference and predictive inference in situations where the choice between either is difficult. Intriguingly, they do so through a posterior that expresses epistemic uncertainty in a way that is meaningfully distinct from standard and generalised Bayesian approaches.
We strengthen this intuitive appeal by demonstrating that PrO posteriors possess strong theoretical guarantees and an algorithmically viable approach for computation. To do so, we first set the scene and formally introduce PrO posteriors (Section 1). Next, we investigate the nature of the uncertainty which they confer, show that they are adaptive to model misspecification, and explain their broader place in the statistical sciences (Section 2). From there, we go on to formally establish that PrO posteriors predictively dominate classical and generalised Bayesian posteriors, and characterise their asymptotic behaviour (Sections 3 and 8). We end by introducing a sampling algorithm which can be used to sample from PrO posteriors (Section 6), exploring their empirical performance on a range of examples (Section 5), and providing a discussion of our findings (Section 7).
1 Background
Throughout, we consider data ${x}_{1:n}=(x_{1},\dots,x_{n})$ for $x_{i}\in\mathcal{X}$ and $i=1,\dots,n$. We assume that ${x}_{1:n}$ was generated from an unknown distribution $P_{0}$. With this, we denote expectations of a function $f:\mathcal{X}\to\mathbb{R}$ with respect to $P_{0}$ by $\mathbb{E}[f(X)]$ or by $\mathbb{E}_{X\sim P_{0}}[f(X)]$ if we wish to emphasise the sampling distribution. To learn about $P_{0}$ from $x_{1:n}$, we follow the tenants of classical Bayesian analysis, and posit a model class $\mathcal{M}_{\Theta}:=\{P_{\theta}:\theta\in\Theta\}$ and a prior $\Pi\in\mathcal{P}(\Theta)$ over elements of $\mathcal{M}_{\Theta}$, where $\mathcal{P}(\Theta)$ denotes the space of probability measures on $\Theta$. Instead of attempting to learn the best possible parameter, as in classical (or even generalised) Bayesian inference, we focus on learning the best possible predictive distribution. To rigorously formulate this optimality, we use scoring rules.
Definition 1 (Scoring Rule).
Let $S:\mathcal{P}(\mathcal{X})\times\mathcal{X}\to\mathbb{R}\cup\{\infty\}$ be an extended real-valued function. Then $S$ is a proper scoring rule, relative to a class $\mathcal{P}$, if:
  • (i)
    the expected score $\mathcal{S}(P,P^{\prime}):=\mathbb{E}_{X\sim P^{\prime}}\left\{S(P,X)\right\}$, is finite for some $P,P^{\prime}\in\mathcal{P}$;
  • (ii)
    $\mathcal{S}(P^{\prime},P^{\prime})\leq\mathcal{S}(P,P^{\prime})$ for all $P,P^{\prime}\in\mathcal{P}$, so that $\mathcal{S}$ induces the statistical divergence $\mathcal{D}_{S}(P,P^{\prime}):=\mathcal{S}(P,P^{\prime})-\mathcal{S}(P^{\prime},P^{\prime})\geq 0$ for which $\mathcal{D}_{S}(P,P^{\prime})=0\Longleftrightarrow P=P^{\prime}$.
Scoring rules are foundational statistical objects concerned with measuring accuracy of predictive distributions (see, e.g., Gneiting and Raftery, 2007). Given a target distribution $P_{0}$, a candidate predictive distribution $P$ is judged to be more accurate than an alternative candidate $P^{\prime}$ by an expected score $\mathcal{S}$ if and only if $\mathcal{S}(P,P_{0})\leq\mathcal{S}(P^{\prime},P_{0})$. The logarithmic scoring rule, given by $S(P,x)=-\log\mathsf{d}P(x)$, is the most common example and is related to the Kullback-Leibler divergence via $\mathrm{KL}(P^{\prime}\|P)=\mathcal{D}_{S}(P,P^{\prime})$. Other popular choices include the spherical score, cumulative ranked probability score (CRPS), the energy score, and squared maximum mean discrepancy (MMD), which for a kernel function $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{+}$ and the Dirac measure at $x$ denoted by $\updelta_{x}$ is given by $\mathrm{MMD}^{2}(P,\updelta_{x}):=\mathbb{E}_{X\sim P,X^{\prime}\sim P}[k(X,X^{\prime})]+k(x,x)-2\mathbb{E}_{X\sim P}[k(X,x)].$
1.1 Key Ideas
The Bayes posterior $\Pi_{n}$ and its associated posterior predictive distribution are
                                                                                                    $\Pi_{n}(A):=\frac{\int_{A}\mathsf{d}P_{\theta}(x_{1:n})\mathsf{d}\Pi(\theta)}{\int_{\Theta}\mathsf{d}P_{\theta}(x_{1:n})\mathsf{d}\Pi(\theta)},\quad\quad P_{\Pi_{n}}:=\int_{\Theta}P_{\theta}\mathsf{d}\Pi_{n}(\theta).$                                                                                                      (1)
In this set-up, the predictive distribution $P_{\Pi_{n}}$ is derived as a consequence of the Bayes posterior $\Pi_{n}$—which itself is primarily geared towards quantifying the uncertainty about which parameter $\theta$ best describes the observed data $x_{1:n}$. In this paper, we reverse that order: the posterior predictive $P_{Q}:=\int_{\Theta}P_{\theta}\mathsf{d}Q(\theta)$ is the object of interest, and we obtain a predictively oriented (PrO) posterior as the measure $Q_{n}$ which quantifies uncertainty about $\theta$ in a manner that is optimal for prediction.
The key motivation for this novel methodology is that while $P_{\Pi_{n}}$ is predictively optimal if $P_{0}\in\mathcal{M}_{\Theta}$ (see Aitchison, 1975), this optimality breaks down once $P_{0}\notin\mathcal{M}_{\Theta}$. Indeed, the Bayes posterior predictive can yield extremely poor predictive performance under even seemingly innocuous model misspecification (see Grünwald and van Ommen, 2017). Though this is well-understood, robust Bayesian methods usually still treat prediction as a second-order concern, and are more concerned with parameter uncertainty. One family emblematic of this kind of research are Gibbs posteriors (e.g. Zhang, 2006; Jiang and Tanner, 2008; Jewson et al., 2018; Knoblauch et al., 2018; Giummolè et al., 2019; Syring and Martin, 2019; Schmon et al., 2020; Chérief-Abdellatif and Alquier, 2020; Matsubara et al., 2022; Martin and Syring, 2022; Matsubara et al., 2023; Altamirano et al., 2023a, b; Duran-Martin et al., 2024; McLatchie et al., 2025; Laplante et al., 2025; Frazier et al., 2025). Many of these posteriors are based on an scoring rule evaluated on the sample, and for $A\subseteq\Theta$ are given by
                                                                                                    $Q_{n}^{\dagger}(A)$ ${}:={}$ $\frac{\int_{A}\exp\left\{-\frac{\lambda_{n}}{n}\sum_{i=1}^{n}S(P_{\theta},x_{i})\right\}\mathsf{d}\Pi(\theta)}{\int_{\Theta}\exp\left\{-\frac{\lambda_{n}}{n}\sum_{i=1}^{n}S(P_{\theta},x_{i})\right\}\mathsf{d}\Pi(\theta)},\quad\quad P_{Q_{n}^{\dagger}}:=\int_{\Theta}P_{\theta}\mathsf{d}Q_{n}^{\dagger}(\theta),$                                                                                                      (2)
where the tuning parameter $\lambda_{n}$ is often called the learning rate. Using Gibbs posteriors for parameter inference is often called generalised Bayesian inference because they recover the Bayes posterior for the special case where $S$ is the logarithmic scoring rule and $\lambda_{n}=n$. To contrast them with PrO posteriors, it will prove instructive to rewrite Gibbs posteriors as the outcome of the following variational optimisation problem:
                                                                                                    $Q_{n}^{\dagger}$ ${}={}$ $\operatornamewithlimits{argmin}_{Q\in\mathcal{P}(\Theta)}\left\{\frac{\lambda_{n}}{n}\sum_{i=1}^{n}\int S(P_{\theta},x_{i})\mathsf{d}Q(\theta)+\mathrm{KL}(Q\|\Pi)\right\}.$                                                                                                      (3)
The above representation is well-established (see e.g., Theorem 1 of Knoblauch et al., 2022), and makes clear that all Gibbs posteriors—including the Bayes posterior—seek to minimise the averaged scoring rule, rather than the score of the averaged model. This in turn means that the inferences they produce are targeted towards finding the parameter, rather than the predictive, which best describes the data. Crucially, these two goals do not overlap outside of well-specified models, where $P_{0}\in\mathcal{M}_{\Theta}$ (see Aitchison, 1975). The result is that in empirical applications, Bayes posteriors and Gibbs posteriors alike often deliver poorly performing posterior predictives.
What then should we do if we wish to conduct inference on model parameters, but wish to ensure that the prediction made under the model are accurate, while also suspecting that $P_{0}\notin\mathcal{M}_{\Theta}$? We argue that, while the parameters of interest remain $\theta\in\Theta$, we should ensure that the beliefs are generated in such a way that the resulting posterior predictive $P_{Q}:=\int P_{\theta}\mathsf{d}Q(\theta)$ is accurate. To do so, we must score the predictive $P_{Q}$ directly, which gives rise to what we shall call the predictively oriented (PrO) posterior given by
${Q}_{n}=\operatornamewithlimits{argmin}_{Q\in\mathcal{P}(\Theta)}\left\{\frac{\lambda_{n}}{n}\sum_{i=1}^{n}S\left(P_{Q},x_{i}\right)+\mathrm{KL}(Q\|\Pi)\right\}.$ (4)
The PrO posterior can also be represented explicitly through a fixed point equation as
                                                                                                    $\mathsf{d}Q_{n}(\theta)$ ${}\propto{}$ $\mathsf{d}\Pi(\theta)\cdot\exp\left\{-\frac{\lambda_{n}}{n}\sum_{i=1}^{n}\left.\frac{\delta S(P_{Q},x_{i})}{\delta Q}\right|_{Q=Q_{n}}(\theta)\right\},$                                                                                                      (5)
where $\frac{\delta S(P_{Q},x)}{\delta Q}\big|_{Q=Q_{n}}$ denotes the functional derivative of $S(P_{Q},x)$ with respect to $Q$ evaluated at $Q={Q_{n}}$ (see also Appendix B.1). While generally of limited use for gaining intuition about its behaviour, this representation illustrates that the PrO posterior is generally not a coherent belief update in the sense of Bissiri et al. (2016).
1.2 Related Work
Approaches and ideas conceptually related to PrO posteriors have appeared in the machine learning literature under a variety of names and motivations. One prominent line of work is rooted in PAC-Bayesian analyses that introduce diversity into an ensemble such as weighted majority voting, whereby a set of classifiers are assigned weights, and the final prediction is the label supported by the largest total weight across the ensemble (Lacasse et al., 2006; Laviolette et al., 2011; Germain et al., 2015; Laviolette et al., 2017; Lorenzen et al., 2019; Masegosa et al., 2020; Viallard et al., 2021; Wu et al., 2021; Wu and Seldin, 2022; Vignault et al., 2023; Theisen et al., 2024). Related PAC-Bayesian perspectives extending beyond majority voting have also been explored by Audibert (2004), Ortega et al. (2022), Hauptvogel and Igel (2024) and Pitas et al. (2025).
More directly related is the recent work of Shen et al. (2025), which proposes the special case of (4) with the MMD in the context of misspecified deterministic simulation models as a heuristic approach “for arresting the collapse of predictive uncertainty” to avoid that “posterior predictions become deterministic.” This is very different from what we study in the remainder: Shen et al. (2025) make no attempt at advocating for the general formulation in (4) as a new general-purpose methodology for parameter and predictive inference, and do not provide any rigorous investigation into its behaviour.
Several alternative heuristic variational approaches targeting objectives similar to (4) have also been suggested elsewhere (see Wang and Liu, 2019; Sheth and Khardon, 2020; Wei et al., 2021; Jankowiak et al., 2020b, a; Lai and Yao, 2024). Amongst these, the most relevant for our remaining developments are two papers studying PAC-Bayesian guarantees for the special case of an approximated logarithmic scoring rule termed $\operatorname{PAC}_{T}^{2}$-variational posterior (Masegosa, 2020) and $\operatorname{PAC}^{m}$-Bayes (Morningstar et al., 2022). Despite superficial similarities, our development differs significantly from these precursors in both form and function. Most critically, this vein of prior work does not define or study the PrO posterior. Instead, the focus is on deriving computationally efficient methods for Bayesian deep learning via variational approximation algorithms informed by objectives similar to (4), as well as the derivation of PAC-Bayes bounds to justify these algorithms.
In summary, previous contributions studying objectives and objects with similarities to (4) are limited in scope, motivated heuristically, and do not explore the promise of PrO posteriors as a new and universally applicable inferential methodology. In the remainder, we rectify this: we provide practically meaningful guarantees for PrO posteriors and study the nature of the uncertainty they communicate. Based on our findings, we argue that PrO posteriors provide a new form of uncertainty quantification, and are uniquely suited to address the predictively oriented inferences that scientists are increasingly demanding from the statistical sciences.
2 A Motivating Example
Throughout, robustness to model misspecification is an important motivation in our development of PrO posteriors. Robustness is also one of the key drivers behind many Gibbs posteriors, usually to guard against Huber contamination (Altamirano et al., 2023a, b; Matsubara et al., 2022, 2023; Laplante et al., 2025; Duran-Martin et al., 2024; Marusic et al., 2025) . Whenever the nature of misspecification is more complicated than can be described through an outlier model of the Huber type, these approaches fail to provide robustness.
To contrast this behaviour with that of PrO posteriors, we consider a linear regression model which for $x_{i}\in\mathbb{R}$, $z_{i}\in\mathbb{R}^{2}$ and $i=1,\dots n$ is given by
                                                                                                    $x_{i}$ ${}={}$ $z_{i}^{\top}\theta+\varepsilon_{i}.$                                                                                                      (6)
Here, we make two assumptions: the parameter of interest $\theta\in\mathbb{R}^{2}$ is fixed but unknown, and $\varepsilon_{i}\overset{\text{iid}}{\sim}\mathcal{N}(0,\sigma^{2})$ for a fixed and known noise level $\sigma^{2}$. Relative to this model for the data, we explore three different forms of misspecification:
  1. (T)
    Heavy-tailed observation noise $\varepsilon_{i}\overset{\text{iid}}{\sim}t_{3}(0,\sigma^{2})$;
  2. (CR)
    Random coefficients $\theta_{2}=2\cdot(-1)^{\xi_{i}}$, with $\xi_{i}\overset{\text{iid}}{\sim}\operatorname{Ber}(1/2)$ sampled from a Bernoulli;
  3. (NT)
    The combined effect of $\varepsilon_{i}\overset{\text{iid}}{\sim}t_{3}(0,\sigma^{2})$ and $\theta_{2}=2\cdot(-1)^{\xi_{i}}$ with $\xi_{i}\overset{\text{iid}}{\sim}\operatorname{Ber}(1/2)$.
Figure 1: The Gibbs and PrO posterior distributions under different forms of model misspecification. Coefficients $\theta_{1},\theta_{2}$ are shown as black crosses, with $\theta_{1}=\theta_{2}$ under trivial misspecification.
The first setting constitutes a form of trivial (T) misspecification: the inflated noise leads to overconfident inferences, but does not compromise the meaning of the model parameter $\theta$. In line with this, we would expect both Gibbs and PrO posteriors to perform reasonably well at identifying the general location of this parameter. In contrast, the second setting fundamentally compromises the interpretation of inferences on $\theta$. However, it allows convex recovery (CR): a 2-component mixture over the misspecified linear regression model recovers the actual data-generating process. While Gibbs posteriors will be adversely affected by this setting, PrO posteriors will be able to recover the true data-generating process exactly. The last setting merges both of the previous complications, and constitutes a form of non-trivial (NT) misspecification: not even convex combinations of the original model class will recover the true data-generating process exactly. Still, we expect that the PrO posterior will perform well, since a convex combination of the original model will get us very close to the true data-generating mechanism.
Figure 2: Mean and standard error of Gibbs posterior’s predictive loss relative to PrO posteriors evaluated using negative log likelihood (NLL), squared MMD ($\operatorname{MMD}^{2}$), and Continuous Ranked Probability Score (CRPS) on a test set for different forms of misspecification.
2.1 Numerical Results
Our numerical results confirm these intuitions, and are summarised in Figures 1, where we compare posterior inferences for Gibbs and PrO posteriors based on the MMD in all three settings, with implementation details and further information deferred to Appendix D.4. Unlike the Gibbs posterior, the PrO posterior performs uncertainty quantification for both convex recovery and the non-trivially misspecified setting by accounting for the distortions of the data-generating mechanism relative to the posited simple linear regression. Even in the trivially misspecified setting, the PrO posterior’s parameter uncertainty seems a more appropriate reflection of the true underlying uncertainty. The reason is that the PrO posterior is driven by its implied predictive distribution, and therefore is more sensitive to the inflated noise in the trivially misspecified setting. Figure 2 translates the impact of parameter uncertainty into predictive performance, and shows an even clearer advantage of the resulting PrO posterior predictive distributions.
2.2 A Taxonomy of Model Misspecification
Based on these observations, it will prove useful to rigorously define this so far intuitively motivated taxonomy of model misspecification. As we shall see later, doing so will prove critical for comparing the behaviours of Gibbs and PrO posteriors.
Definition 2.
We say that $\mathcal{M}_{\Theta}$ is well-specified if $P_{0}\in\mathcal{M}_{\Theta}$, and misspecified if $P_{0}\notin\mathcal{M}_{\Theta}$. Further, we call a misspecified $\mathcal{M}_{\Theta}$ trivially misspecified with respect to a scoring rule $\mathcal{S}$ if there exists $\theta^{\prime}\in\Theta$ so that $\inf_{Q\in\mathcal{P}(\Theta)}\mathcal{S}(P_{Q},P_{0})\geq\mathcal{S}(P_{\theta^{\prime}},P_{0})$, and non-trivially misspecified if $\mathcal{S}(P_{Q},P_{0})<\inf_{\theta\in\Theta}\mathcal{S}(P_{\theta},P_{0})$ for some $Q\in\mathcal{P}(\Theta)$. Further, we say that a misspecified $\mathcal{M}_{\Theta}$ allows convex recovery if there exists $Q\in\mathcal{P}(\Theta)$ such that $P_{Q}=P_{0}$.
To aid understanding, Figure 3 provides an illustration of these different forms of model misspecification. What we call non-trivial misspecification is a generalisation of what Grünwald and van Ommen (2017) termed “bad misspecification”. Convex recovery is a special case of this, and implies that the extended model class $\{P_{Q}:Q\in\mathcal{P}(\Theta)\}$ obtained via convex combinations of $\mathcal{M}_{\Theta}$ contains $P_{0}$.
$\mathcal{M}_{\Theta}$$P_{\theta^{\star}}=\arg\min_{\theta}\mathcal{S}(P_{\theta},P_{0})$$P_{0}$
(a) Trivial misspecification.
$\mathcal{M}_{\Theta}$$P_{\theta_{2}}$$P_{\theta_{1}}$$P_{Q^{\star}}$$P_{\theta^{\star}}$$P_{0}$
(b) Non-trivial misspecification.
[…]
(c) Convex recovery.
Figure 3: Illustration of key ideas introduced in Definition 2. Panel (a) depicts trivial model misspecification: relative to the chosen score $\mathcal{S}$, a singular element of $\mathcal{M}_{\Theta}$ provides the best fit for $P_{0}$, and there is no benefit to averaging. In contrast, Panel (b) illustrates the case of non-trivial model misspecification: constructing ${Q^{\star}}:=\operatornamewithlimits{argmin}_{Q}\mathcal{S}(P_{Q},P_{0})$ as a convex combination of $P_{\theta_{1}}$ and $P_{\theta_{2}}$ yields a better predictive for $P_{0}$ than any one point in $\mathcal{M}_{\Theta}$. Panel (c) provides pictorial intuition for convex recoverability: it is the special case where $P_{Q^{\star}}$ and $P_{0}$ coincide.
3 Formal Guarantees
The motivating example in the previous section shows empirically that PrO posteriors adapt to model misspecification: when the model is well-specified, they behave like Gibbs or Bayes posteriors, but deliver fundamentally different behaviours when the model is misspecified. In this section, we clarify this adaptivity and formally show that PrO posteriors are predictively superior to Bayes and Gibbs posteriors. Throughout, we will often use the shorthand predictive scoring rule to refer to a scoring rule $S(P_{Q},x)$ applied directly to score a predictive distribution $P_{Q}=\int P_{\theta}\mathsf{d}Q(\theta)$. Additionally, for two sequences $a_{n},b_{n}$ we write $a_{n}\lesssim b_{n}$ to mean that $a_{n}\leq Cb_{n}$ for all $n$ large enough and some universal constant $C>0$. Further, we say that $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$.
3.1 Regularity conditions
Throughout, we impose a minimal convexity assumption satisfied by virtually all scoring rules of interest—including kernel scoring rules and the logarithmic score.
Assumption 1.
The scoring rule $S(\cdot,x)$ is convex in its first argument for all $x\in\mathcal{X}$, and its corresponding divergence $\mathcal{D}_{S}$ is jointly convex in both arguments.
In addition to convexity of the scoring rule, we require regularity conditions on their expectations. While less restrictive conditions can be obtained, the following regularity condition is easy to state, and suffices for our theoretical analysis.
Assumption 2.
Let $r_{n}$ be a positive sequence such that $r_{n}\rightarrow 0$ as $n\rightarrow\infty$ and $\mathcal{P}_{\Pi}(\Theta):=\{Q\in\mathcal{P}(\Theta):Q\ll\Pi\}$, it holds that $\mathbb{E}\left[\sup_{Q\in\mathcal{P}_{\Pi}(\Theta)}\left|\frac{1}{n}\sum_{i=1}^{n}S(P_{Q},X_{i})-\mathcal{S}(P_{Q},P_{0})\right|\right]\lesssim r_{n}$.
As the speed at which the sequence $r_{n}$ decays depends on the scoring rule, we state it in full generality. For the special case of bounded scoring rules, $r_{n}$ will always be of order $\log(n)/{n}^{1/2}$. Further, even for unbounded scoring rules such as the logarithmic scoring rule, the rate $\log(n)/{n}^{1/2}$ is achievable at the expense of additional restrictions and technicalities using techniques such as those in Ghosal and Van Der Vaart (2001). For further discussion and details, see Section 8.
Though Assumption 2 is satisfied by many scoring rules, more interpretable alternatives exist. In particular, Assumption 1 and Jensen’s inequality imply that
                                                                                                    $S\left(P_{Q},x\right)$ ${}={}$ $\int S\left(P_{\theta},x\right)\mathsf{d}Q(\theta)-\Delta(Q,x),$                                                                                                      (7)
for all $x\in\mathcal{X}$ and some non-negative remainder term $\Delta:\mathcal{P}(\Theta)\times\mathcal{X}\to\mathbb{R}_{+}$. In many cases, such as the log predictive score, this remainder does not simplify further. However, kernel scoring rules (Gneiting et al., 2007) can be written as an expectation over a product measure in $\Theta$ for any $x\in\mathcal{X}$ (see Lemma 2 in Appendix B), so that the predictive score is
                                                                                                    $S\left(P_{Q},x\right)$ ${}={}$ $\int\underbrace{\left\{S(P_{\theta_{1}},x_{i})-\delta(x_{i};\theta_{1},\dots,\theta_{k})\right\}}_{\textstyle=:L(\theta_{1:k},x)}\mathsf{d}Q^{k}(\theta_{1:k});\quad Q^{k}(\theta_{1:k})=\bigotimes_{j=1}^{k}Q(\theta_{j}).$                                                                                                      (8)
for some function $\delta:\mathcal{X}\times\Theta^{k}\to\mathbb{R}_{+}$ with $k\in\mathbb{N}_{+}$. This representation allows us to replace Assumption 2 with more interpretable and testable conditions.
Assumption 2.
The predictive scoring rule satisfies (3.1) and for any $\varphi\in\Theta$, if $\theta_{1:k}=(\varphi,\varphi,\dots\varphi)$, then $\delta(x;\varphi,\dots,\varphi)=0$, for all $x\in\mathcal{X}$. Further, for any $\lambda\in\mathbb{R}$, $\mathcal{L}(\theta_{1:k}):=\mathbb{E}[L(\theta_{1:k},X)]$, and some $C>0$, $\mathop{\text{\large$\int$}}\nolimits\mathbb{E}\left[e^{\lambda\{\mathcal{L}(\theta_{1:k})-\frac{1}{n}\sum_{i=1}^{n}L(\theta_{1:k},X_{i})\}}\right]\mathsf{d}\Pi^{k}(\theta_{1:k})\leq e^{\lambda^{2}C^{2}/n}.$
Here, the first part of Assumption 2 is a mild regularity condition, and automatically satisfied by kernel scoring rules for $k=2$. The second part imposes a Hoeffding-type moment condition that generally has to be verified, and which is very similar to those clasically arising in the study of Gibbs posteriors. Though easier to verify, certain scores such as the logarithmic scoring rule violate Assumption 2, so that the more abstract conditions imposed by Assumption 2 must be maintained instead.
3.2 Key Results
Having established the necessary regularity conditions, we proceed by studying the impact of different misspecification regimes on PrO posteriors $Q_{n}$, and compare them to Gibbs posteriors $Q_{n}^{\dagger}$ constructed using the same scoring rule. To state our results, we define
[…]
and take $\lambda_{n}\asymp{n}^{1/2}$ up to logarithmic factors. This allows us to state our first key result on the predictive performance of PrO posteriors.
Theorem 1.
Assumption 1 is satisfied, and Assumption 2 or 2 holds. Under prior regularity conditions, the following holds for $n$ sufficiently large:
  • $\operatorname{(WS)}$
    If the model is well-specified, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]-\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]\lesssim\nu_{n}$;
  • $\operatorname{(NT)}$
    If the model is non-trivially misspecified, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]<\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]$,
    and $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]-\inf_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{\mathcal{S}}(P_{Q},P_{0})\lesssim\nu_{n};$
  • $\operatorname{(CR)}$
    If the model allows convex recovery, then $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\lesssim\nu_{n}.$
Theorem 1 evaluates the PrO posterior predictive $P_{Q_{n}}$ and its relationship to the Gibbs posterior predictive $P_{Q_{n}^{\dagger}}$ in terms of the divergence $\mathcal{D}_{S}$ induced by the score chosen to construct both posteriors. $\operatorname{(WS)}$ concerns the well-specified setting. Here, Gibbs posteriors typically concentrate towards the correct model parameter at the usual $\log(n)/n^{1/2}$ rate. As the result shows, the difference in predictive performance between PrO posteriors and Gibbs posteriors in this setting is small, and usually of order $\log(n)/n^{1/2}$. For the case of non-trivial misspecification studied in $\operatorname{(NT)}$, the situation changes: now, the PrO posteriors strictly outperforms Bayes and Gibbs posteriors. In particular, this performance gap will be of order $\inf_{\theta\in\Theta}\mathcal{D}_{S}(P_{\theta},P_{0})-\inf_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{S}(P_{Q},P_{0})$, which grows larger as the benefits of averaging over $\mathcal{M}_{\Theta}$ become more pronounced. To confirm this intuition, the result in $\operatorname{(CR)}$ considers the special case of convex recovery, where $\inf_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{S}(P_{Q},P_{0})=0$. As expected, the PrO posterior recovers the true data-generating process at rate $\nu_{n}$. In contrast, the Gibbs posterior predictive $P_{Q_{n}^{\dagger}}$ never recovers the true data generating process in this setting, and instead collapses onto a point mass at $\theta^{\star}=\operatornamewithlimits{argmin}_{\theta\in\Theta}\mathcal{D}_{S}(P_{\theta},P_{0})$. Lastly, the result also shows that the only case in which the choice of scoring rule has a notable impact is the non-trivially misspecified case: when the model satisfies convex recovery or is well-specified, any scoring rule will recover the same predictive distribution for $n$ sufficiently large. This suggests that PrO posteriors are generally less sensitive to the choice of scoring rule than Gibbs posteriors.
The above result validates the use of PrO posteriors whenever predictive performance matters, but sheds little light on the qualitative behaviour of PrO posteriors as expressions of parameter uncertainty. Given that they no longer collapse to a point mass, is it safe to assume that they behave well as the sample size increases? Indeed, in the context of parameter inference, convergence to a limiting distribution as more data are observed is fundamental to ensure that one is learning a well-defined target. Our second result assuages these concerns, and shows that PrO posteriors do indeed converge to a well-defined target given by the predictively optimal model average $Q^{*}:=\operatornamewithlimits{argmin}_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{S}(P_{Q},P_{0})$. Below, we state this convergence relative to a distance $\mathsf{d}$ defined on probability distributions and dependent on additional regularity conditions (see Section 8 for details).
Corollary 1.
Assumption 1 and Assumption 2 or 2 holds. Under convex recovery and additional regularity conditions, there exists $\alpha>0$ so that $\mathbb{E}[\mathsf{d}(Q_{n},Q^{\star})]\leq C(\log n/n^{1/2})^{1/\alpha}.$
While the above result relies on convex recovery, our empirical results strongly suggest that this is an artifact of our proof technique, and that this behaviour should be expected to hold under much weaker conditions.
4 Parameter Uncertainty of the PrO Posterior
In his prescient contribution, Breiman (2001) identified two divergent sub-cultures within statistical science. On the one hand, he spoke of a data modelling culture interested in hand-tailoring interpretable models for a given research problem and thus strongly associated to classical statistical concerns. On the other hand, he identified an increasing interest in machine learning and the associated algorithmic modelling culture focused on predictive accuracy and black box density estimation techniques.
We believe that PrO posteriors constitute a meaningful and novel attempt to incorporate the most useful aspects of both modelling cultures: they rely on interpretable parameters born out of the data modelling culture, but gear inference towards predictive accuracy. When both interpretability and predictive performance are relevant, practitioners are often forced into a fraught choice: should they prefer an interpretable but predictively suboptimal model; or should they resort to a black box model that will improve predictive performance at the cost of interpretability and meaningful guarantees? PrO posteriors circumvent this binary choice. Figure 4, illustrates this by using body measurements for different penguin species in the Palmer Archipelago (Horst et al., 2020). Here, the PrO posterior not only identifies the data set’s fundamental structure by correctly discovering the three different species in it, but also produces a predictive distribution of excellent quality.
Figure 4: Palmer penguins example. We plot both the kernel density estimate (KDE) and the PrO posterior predictive overlaid with the data (left, middle), as well as the PrO posterior overlaid with the average measurements for each of the three penguin species in the data set (right). For inference with the PrO posterior, we use a bivariate Gaussian model. The results show that the PrO posterior not only correctly identifies the three penguin species in the data (right), but also provides a predictive distribution which can fit the observed data as well as a kernel density estimator. For further details, see Appendix D.2.
This behaviour is desirable for a wide array of scientific problems, and yields an altogether new kind of epistemic uncertainty which differs starkly from what is currently attainable in classical Bayesian inference and its generalisations.
4.1 The Origin of PrO Posterior Uncertainty
It is worth contrasting PrO posterior uncertainty with a set of ideas and methods known under the umbrella of predictive Bayes (see e.g. Berti et al., 2021; Fortini and Petrone, 2020, 2023; Fong et al., 2023; Fortini and Petrone, 2025). Despite similar sounding names, the roles played by prediction in the construction of PrO posteriors and predictive Bayes approaches are fundamentally different: the PrO posterior is geared towards learning the predictive from seen data, predictive Bayes approaches instead posits a sequence of predictive distributions to model the unseen data. Put differently, while PrO posteriors make the predictive distribution the object of posterior inference, predictive Bayes methods specify predictive distributions instead of priors and likelihoods.
In fact, priors and likelihoods can be recast as a convenient way of specifying a sequence of predictive distributions on unseen data: formalising this in a recent read paper at the Royal Statistical Society, Fong et al. (2023) used Doob’s Theorem to argue that the source of Bayesian uncertainty is unseen data. Taken to its logical extreme, if we observe the entire population $P_{0}$, no epistemic uncertainty remains, and Bayesian methods become perfectly confident in the conclusions they draw from data. In fact, this same logic applies to various generalisations of Bayesian uncertainty, including Gibbs posteriors.
But should epistemic uncertainty indeed vanish as more data are observed? Though it may be tempting to answer this question in the affirmative, doing so would overlook an important practical limitation: for interpretability and computational feasibility, statistical methods express epistemic uncertainty relative to a parameter $\theta\in\Theta$. Crucially, this implies that the only setting in which larger sample sizes should lead to a collapse of parameter uncertainty is the rare situation where a singular value of $\theta\in\Theta$ can summarise all properties of $P_{0}$ that we are interested in. In other words, the epistemic uncertainty conferred by (generalised) Bayesian methods is constructed relative only to a fixed model class, not with respect to the choice of model class itself, which renders inference and uncertainty quantification unreliable whenever $P_{0}\notin\mathcal{M}_{\Theta}$.
The epistemic uncertainty produced by PrO posteriors is more expansive: it delivers both parameter uncertainty and uncertainty about the inherent ability of $\mathcal{M}_{\Theta}$ to adequately explain the observed data. Indeed, when the model is misspecified, PrO posteriors generally do not collapse to a single element of $\mathcal{M}_{\Theta}$ as more data are observed. When $P_{0}\notin\mathcal{M}_{\Theta}$, they instead stabilise towards an irreducible form of uncertainty that originates in the fact that no single value of $\theta\in\Theta$ is sufficient to summarise all properties of $P_{0}$ we take an interest in. This is epistemic uncertainty in its truest form: because the model is misspecified, we should be uncertain about its ability to represent the observed data, even if we had observed the entire population $P_{0}$.
But what if the model is well-specified, so that $P_{0}\in\mathcal{M}_{\Theta}$? In this case, the epistemic uncertainty produced by Bayes and Gibbs posteriors is sufficient: there exists some $\theta\in\Theta$ that allows us to completely describe the observed data—and as more data accumulates, we should become increasingly certain about the ability of this value to correctly explain the data, and hence $P_{0}$. Remarkably, for this special case where the model is well-specified, PrO posteriors do behave like Bayes and Gibbs posteriors, and likewise concentrate towards the parameter value that recovers $P_{0}$.
4.2 Key Comparisons
To round off the above discussion, we contrast the behaviour of PrO posteriors to that of several other related approaches like standard Bayesian inference, Martingale posteriors (Fong et al., 2023), and mixture models. We do so using the golf putting data of Gelman and Nolan (2002), which compares the probability of a successful put $(x\in\{0,1\})$ by professional golfers to the distance ($z\in\mathbb{R}$) from which the shot was made for $n=5,988$ shots. Throughout, we fit this data using the logistic regression model
                                                                                                    $P^{\mathsf{LRM}}_{\theta}(x\mid z)$ ${}={}$ $\sigma(\theta^{\top}z)^{x}\left\{1-\sigma(\theta^{\top}z)\right\}^{1-x}\text{ for }\;\sigma(\theta^{\top}z)=\left\{1+\exp(\theta_{1}+\theta_{2}z)\right\}^{-1},$                                                                                                      (10)
and chart uncertainty estimates across several methods that derive from this set-up.
Figure 5: Golf putting data. Points represent the proportion of successful putts made by professional golfers as a function of the distance from the hole. The first row shows the predictive mean and one standard deviation intervals of the Bayes, Martingale, and PrO posteriors. The second row shows the marginals of the corresponding parameter posteriors.
4.2.1 Bayes Posterior and Martingale Posterior Uncertainty
First, we compare PrO posterior uncertainty with classical Bayes posteriors and a Martingale posterior (Fong et al., 2023) based on the Bayesian bootstrap and the negative log likelihood loss (see Lyddon et al., 2019; Rubin, 1981), and display the results in Figure 5. Bayes posteriors and Martingale posteriors are built around the idea that a single parameter is sufficient to summarise the data-generating process, and that the only source of uncertainty is unseen data. Accordingly, both produce posteriors that concentrate onto a single parameter value. Their predictive distributions mirror this, and are qualitatively similar to what one would obtain by using the maximum likelihood estimator to construct a plug-in predictive distribution (see McLatchie et al., 2025). This sharply contrasts with the behaviour of the PrO posterior: its predictive distribution indicates that while the success of short puts can be estimated with some fidelity, the success of longer puts is associated with substantive levels of uncertainty. One possible explanation of this phenomenon is that the logistic regression model is too coarse, and that there are in fact two sub-groups of professional golfers: those that are very proficient at longer puts, and those that are not. This hypothesis is backed up by the PrO posterior on $\theta_{2}$, a parameter which models the effect of distance on the probability of a successful put: unlike the Bayes posterior or the Martingale posterior, the marginal PrO posterior of $\theta_{2}$ has two modes which could account for these two sub-groups. To investigate this effect more carefully, we also compute the posterior marginal effect of distance on the probability of a successful put, which for a posterior $Q$ can be calculated as
                                                                                                    $\int\frac{\partial P^{\mathsf{LRM}}_{\theta}(x\mid z)}{\partial z}\mathsf{d}Q(\theta)$ ${}={}$ $\int\theta_{2}\cdot P^{\mathsf{LRM}}_{\theta}(x=1\mid z)P^{\mathsf{LRM}}_{\theta}(x=0\mid z)\mathsf{d}Q(\theta),$                                                                                                      (11)
and we plot these curves as a function of the distance $z$ in Figure 6. Inferences across the Bayes, Martingale, and PrO posteriors agrees on the qualitative nature of the marginal effect of distance: while the marginal effect of distance is negative and decreasing until around 5–7 feet, it increases thereafter and monotonically in the range of 7–20 feet, and rises close to a zero marginal effect. This suggests that at the range of 5–7 feet, every professional golfer will opt to put, and that longer puts are generally more difficult. As the distance to the hole increases, only golfers very proficient at putting over long distances opt to take the shot, so that distance modulates a positive marginal effect on success.
Figure 6: Top row: the mean (solid line) and one standard error intervals (shaded area) of the posterior marginal effects of the Bayes, Martingale, and PrO posteriors as a continuous function of distance in the golf putting data. Bottom row: we investigate the posterior marginal effect under PrO posterior at $2,4,6,8$, and 2020 feet.
While all posteriors agree on the shape of this curve, the accompanying uncertainty differs substantively. Unlike for the Bayes and Martingale posteriors, the marginal effects obtained from the PrO posterior at each distance are rich multimodal distributions, and allow a more fine-grained model criticism. More precisely, they show that while posterior uncertainty of the marginal effect is unimodal for very short and very long distances, it bifurcates into a bimodal distribution for middle distances, which once again reaffirms our previous observations: there are two relevant sub-groups in the data set, and middle distances are the most informative to discriminate between them.
We note that this ease at which PrO posteriors can be used for model criticism has led previous authors to speculate that one may be able to use related methods as a diagnostic tool for model misspecification (see Lai and Yao, 2024; Chazal et al., 2025). While this has previously been argued on the basis of heuristics, our findings in Section 3 provide the statistical tools required for making these ideas more rigorous.
4.2.2 Bayesian Mixture Models
The above example raises an immediate question: given that the PrO posteriors seems to indicate the presence of a group structure in the model, should we think of PrO posteriors as mixture models? In fact, since the PrO posterior objective (4) directly targets the marginalised predictive $\int P_{\theta}\mathsf{d}Q(\theta)$—an object which we might call a mixture—how are the aims of PrO posterior inference different from those of mixture modelling?
To answer this, consider the Bayesian mixture model given by
                                                                                                    $P_{\theta^{(1:M)},\omega^{(1:M)}}(x\mid z)$ ${}={}$ $\sum_{m=1}^{M}\omega^{(m)}\cdot P^{\mathsf{LRM}}_{\theta^{(m)}}(x\mid z),$                                                                                                      (12)
where the vector $\omega^{(1:M)}$ is non-negative and sums to one. One of the key practical challenges in Bayesian mixture modelling consists in choosing priors on $M$ and the mixture weights $\omega^{(1:M)}$. This is a key difference: PrO posteriors make no attempt at modelling mixtures over $P_{\theta}^{\textsf{LRM}}$, and thus require no such priors. Instead, their priors are formulated with respect to the typically more interpretable parameters of the base model $P_{\theta}^{\textsf{LRM}}$. Thus, while the PrO posterior $Q_{n}$ is a predictively oriented mixing measure over the base model $P_{\theta}^{\textsf{LRM}}$, this mixing measure itself requires neither a statistical model nor a prior.
This is a sharp contrast to mixture models: as their name implies, they explicitly model the mixture over $P_{\theta}^{\textsf{LRM}}$. This changes the parameter of interest: relative to the simple logistic model in (4.2), a mixture model with $M$ components inflates the number of parameters by a factor of $M$. Though the resulting mixture $P_{\theta^{(1:M)},\,\omega^{(1:M)}}$ is more flexible than $P_{\theta}^{\textsf{LRM}}$, it is also less parsimonious, and thus more difficult to interpret. In fact, mixture models are usually used as density estimators, with parameters in each component being synthetic artifacts without inherent meaning, so that $P_{\theta^{(1:M)},\,\omega^{(1:M)}}=P_{\theta^{\rho(1:M)},\,\omega^{\rho(1:M)}}$ for any permutation $\rho$.
The above reveals several important practical differences between PrO posteriors and Bayesian mixture models: the latter abide by orthodox Bayesian modelling, which comes at the price of having to model the mixture and changing the parameter of interest. In contrast, PrO posteriors continue to draw inferences with respect to a simpler and more interpretable parameter. This allows for conclusions, summaries, and model criticisms that lack clear counterparts in mixture models. For instance, since
                                                                                                    $\log\left\{\frac{P_{\theta}^{\mathsf{LRM}}(x=1\mid z)}{P_{\theta}^{\mathsf{LRM}}(x=0\mid z)}\right\}$ ${}={}$ $\theta_{1}+\theta_{2}z,$                                                                                                      (13)
the parameter $\theta_{2}$ is directly interpretable as the linear effect of distance on the log-odds of a successful put. The PrO posterior uncertainty which indicates two sub-groups of golfers will therefore map directly onto the log-odds of the inferred model. We illustrate this phenomenon in Figure 7, which shows that for larger put distances, the skill differences between these two groups become more pronounced (as measured in log odds). Mixture models would not admit this visualisation: they make it impossible to express $(\theta^{(1:M)},\omega^{(1:M)})$ in terms of their linear contributions to the log-odds of the underlying model $P_{\theta^{(1:M)},\,\omega^{(1:M)}}$.
Figure 7: The distribution of PrO posterior log-odds for the golf putting data. The black dots represent the posterior mean at those distances, and the solid line represents the posterior mean as a continuous function of distance.
5 Illustrations
We supplement the discussion of the previous section by illustrating the promise of PrO posteriors on three statistical problems. In Section 5.1, we highlight the implications of convex recovery in a binary classification problem. Moving past independently sampled data, the remaining examples illustrate that PrO posteriors outperform Gibbs posteriors even when the theory’s assumptions are violated. In this vein, Section 5.2 focuses on predicting temporally dependent water flows of Stansted Brook in England, while Section 5.3 investigates the impact of using different scoring rules in a graphical model for spatially dependent house prices in Boston.
5.1 Binary classification
To strengthen intuitions about the predictive benefits conferred by PrO posteriors, we construct a synthetic data set in which a generalised linear model with logistic link function allows for convex recovery as set out in Definition 2. Labels $y_{i}$ are assigned deterministically as a function of the covariate vector $(x_{i}^{(0)},x_{i}^{(1)})$ which is sampled uniformly on $2,2]^{2}$. When the covariate vector falls into the top-left quadrant of $2,2]^{2}$ (so that $x_{i}^{(0)}<0,x_{i}^{(1)}>0$), $y_{i}=0$. In contrast, the bottom-left quadrant (for which $x_{i}^{(0)}>0,x_{i}^{(i)}<0$) yields $y_{i}=1$, while labels in the remaining two quadrants are sampled as $y_{i}\sim\operatorname{Ber}(0.5)$.
We generate $n=1,000$ observations from this model and compute both the corresponding Bayes posterior and the PrO posterior based on the logarithmic scoring rule. We compare their posterior predictive distributions in Figure 8, and conclude that Bayesian inference produces a subpar decision boundary which leads to an overconfident misinterpretation about the data-generating process in the top-right and bottom-left quadrants. In contrast, the PrO posterior identifies the correct decision boundary, and recovers a predictive distribution $P_{Q_{n}}$ for which $\mathrm{KL}(P_{0}\|P_{Q_{n}})$ is close to zero. This mirrors the predictions of Theorems 1, and is a consequence of the fact that the PrO posterior predictive allows for convex recovery of the exact decision boundary. Figure 9 further illustrates this by probing the effect of $n$, and confirms that the theory derived is practically relevant even for relatively small data set sizes.
Figure 8: Top row: Bayes and PrO posterior predictive distributions overlaid with the raw data. Bottom row: Decision boundaries induced by individual draws from the posterior predictives. While the Bayes posterior concentrates onto a single decision boundary, the PrO posterior is multimodal.
Figure 9: Constructing Bayes and PrO posteriors using $n$ observations, we plot their negative log likelihood (NLL) on test data, with standard error intervals obtained from 2020 replications.
5.2 River flow data
Throughout, our theoretical developments have used the assumption of independently sampled data. In practice however, prediction is often most interesting for sequentially or spatially dependent data. While the theoretical development provided in the current paper cannot account for this, the following example shows that PrO posteriors continue to function as expected in temporally dependent data (see also Shen et al., 2025).
Here, we focus on predicting the gauged daily flow of Stansted Brook north of London using measurements of the UK’s Environment Agency. We model the flow using an auto-regressive model of order seven with an intercept term, and compare the inferences obtained from a Gibbs posterior and a PrO posterior using the squared MMD. With this set-up, we fit the model using data from March to November 2002, and use it to obtain one-step-ahead predictions over the course of the transition from the exceptionally wet winter of 2002/2003 into the country’s driest February–-October period since 1921 (Marsh, 2004). This abrupt shift from heavy rain in January to the dry and warm conditions from February onwards produced severe though short-lived hydrological stress, and offers an interesting regime for testing the adaptivity of the PrO posterior. For further details, see Appendix D.6.
The results of our experiments are summarised in Figure 10. Here, we show the one-step-ahead predictive means over the period of November 2002 to March 2003 in the left column. We find that the richer parameter uncertainty conferred by the PrO posterior is able to handle the variability in the data more easily than the Gibbs posterior. This is also evident from the right column of the Figure, which charts the cumulative performance of the leave-one-out predictions. In summary, despite violating the assumption of independently sampled data, we can recover the clear benefits of PrO posteriors. This suggests that the independence assumptions imposed in our work are not required to justify using the PrO posterior in practice.
Figure 10: River flow data. In the left column we show the water flow (log scaled) in the winter leading up to the drought of 2003 along with the mean of the Gibbs and PrO predictive distributions respectively. In the right column we show the one-step-ahead cumulative mean predictive metrics.
5.3 Boston housing data
Strengthening the observations made in our previous experiment, we now demonstrate that PrO posteriors continue to function reliably in spatial settings by modelling the median house price across $n=506$ different census tracts in Boston in the 1970s and 1980s. These data were originally collected by Harrison and Rubinfeld (1978), and are summarised graphically in the left-most panel of Figure 11. Here, census tracts $i$ and $j$ are displayed as nodes in an undirected graph with adjacency matrix $\Omega$ whose entry at $(i,j)$ is given by $\omega_{i,j}=0$ unless the tracts $i$ and $j$ share at least one vertex, in which case $\omega_{i,j}=1$.
Median house values are then modelled as the conditional auto-regression
$y_{i}\mid y_{1},\dots y_{i-1},y_{i+1},\dots y_{n}\sim\mathsf{N}\left(\theta\sum_{j\neq i}\tilde{\omega}_{i,j}y_{j},\sigma^{2}\right),$
where $\tilde{\omega}_{i,j}$ denote the rescaled $\omega_{i,j}$ so that $\sum_{j\neq i}\tilde{\omega}_{i,j}=1$ for all $i$ and where $\sigma^{2}>0$ is a fixed constant which we estimate from data (for a more detailed explanation, see Appendix D.3). From this model, we construct two Gibbs and PrO posteriors for $\theta$ using the squared MMD and log score respectively; see Appendix D.8 for further details.
Importantly, the underlying data distribution is bimodal, and has its primary mode at around $\mathdollar 20,000$ and a second smaller mode at $\mathdollar 50,000$ (see Appendix D.8), with the second mode is due to censoring values exceeding $\mathdollar 50,000$ (Gilley and Pace, 1996). In other words, the data set has two sub-groups: tracts with median prices close to $\mathdollar 20,000$, and tracts with median prices are more than twice that value. Our results in Figure 11 show the effect on inference: while the Gibbs posterior fails to capture this hidden structure in the data, the PrO posterior transfers it to the parameter space. Its bi-modality reflects the fact that while there is a small group of census tracts with very high house prices surrounded by cheaper areas, a much larger group will be surrounded by similarly priced tracts. While the former group is accounted for through a smaller mode around $\theta\in[1.2,1.4]$, the latter is captured by the larger mode around $\theta\in[0.8,1.0]$. As the rightmost panel of Figure 11 confirms, this is not only a richer description of parameter uncertainty, but also leads to better predictive performance. Confirming our earlier observation that the choice of scoring rule seems to matter less for PrO posteriors, these results and interpretations are qualitatively the same whether one uses the squared MMD or the logarithmic scoring rule.
Figure 11: Top left: illustration of the Boston housing data set. Nodes correspond to census tracts, and shade represents price level. Census tracts are connected by edges if they are spatially adjacent. Middle columns: Gibbs and PrO parameter posteriors under the MMD score (top row) and the log score (bottom row). Right column: mean and standard error of the leave-one-node-out negative log likelihood (lower is better).
6 Computation via Wasserstein Gradient Flows
The tools for the computation of PrO posteriors are limited. In fact, most previous work targeting related objects did not even consider the possibility of exact computation, and instead focused on approximating $Q_{n}$ with a parameterised variational family (see Jankowiak et al., 2020a, b; Masegosa, 2020; Morningstar et al., 2022; Lai and Yao, 2024). The only pre-existing work directly targeting exact computation is Shen et al. (2025), which built on the ideas in Wild et al. (2023) and the mean-field Langevin algorithm of Del Moral (2013) to derive an asymptotically exact sampler for the special case where the scoring rule is the squared MMD. Here, we explain how this can be generalised to any PrO posterior.
We may rewrite the objective minimised in (4) as
                                                                                                    $\mathscr{L}(Q)$ ${}={}$ $\frac{\lambda_{n}}{n}\sum_{i=1}^{n}S\left(P_{Q},x_{i}\right)-\int\log\mathsf{d}\Pi(\theta)\,\mathsf{d}Q(\theta)+\int\log\mathsf{d}Q(\theta)\,\mathsf{d}Q(\theta).$                                                                                                      (14)
For functionals like this, the Wasserstein gradient flow (WGF) can be used to construct a gradient descent scheme via interacting particle systems derived from a McKean-Vlasov process. If the Wasserstein gradient of $\frac{1}{n}\sum_{i=1}^{n}S\left(P_{Q},x_{i}\right)$ is given by $\mathcal{W}(Q)$, the Wasserstein gradient of $\mathscr{L}(Q)$ with respect to $Q$ evaluated at $\vartheta\in\Theta$ is
                                                                                                    $\nabla_{\operatorname{W}}\mathscr{L}(Q)[\vartheta]$ ${}={}$ $\lambda_{n}\mathcal{W}(Q)[\vartheta]-\nabla_{\vartheta}\log\mathsf{d}\Pi(\vartheta)+\nabla_{\vartheta}\log\mathsf{d}Q(\vartheta).$                                                                                                      (15)
The WGF $\{Q_{t}\}_{t>0}$ defined by this gradient is the continuous-time process which evolves in the direction of steepest descent relative to the convex functional $Q\mapsto\mathscr{L}(Q)$ given by $\nabla_{\operatorname{W}}\mathscr{L}(Q)$. As $\mathscr{L}(Q)$ is convex, and as the minimiser of $\mathscr{L}(Q)$ is the PrO posterior, the WGF acts as a continuously-valued version of gradient descent on probability measures, so that $Q_{t}$ converges to the PrO posterior as $t\to\infty$ under mild conditions.
To implement this gradient descent algorithm numerically, we resort to the WGF’s mean-field representation, which gives rise to the stochastic differential equation (SDE)
                                                                                                    $\mathsf{d}\vartheta_{t}$ ${}={}$ $-\left\{\lambda_{n}\mathcal{W}(Q_{t})[\vartheta_{t}]-\nabla_{\vartheta}\log\mathsf{d}\Pi(\vartheta_{t})\right\}\mathsf{d}t+\sqrt{2}\mathsf{d}B_{t},\;\;\text{ for }\vartheta_{t}\sim Q_{t},$                                                                                                      (16)
and where the Brownian motion $B_{t}$ appears as the WGF of entropy corresponds to the heat equation (Jordan et al., 1998). Unfortunately, the above SDE still depends on the unknown measure $Q_{t}$, which makes exact forward simulation impossible. To remedy this, we use the approximations $Q_{t}^{(j)}(\theta)\approx\frac{1}{p}\sum_{\ell=1,\ell\neq j}^{p}\updelta_{\vartheta_{t}^{(\ell)}}$, where $\updelta_{x}$ denotes the Dirac measure at $x$, and $\vartheta_{t}^{(j)}$ for $j=1,2,\dots p$ are a set of $p$ interacting particles. This finally results in a system of $p$ interacting particles whose evolution is governed by the $p$ equations
                                                                                                    $\mathsf{d}\vartheta_{t}^{(j)}$ ${}={}$ $-\left\{\mathcal{W}(Q_{t}^{(j)})[\vartheta_{t}^{(j)}]-\nabla_{\vartheta}\log\mathsf{d}\Pi(\vartheta_{t}^{(j)})\right\}\mathsf{d}t+\sqrt{2}\mathsf{d}B_{t}^{(j)},\quad j=1,2\dots p,$                                                                                                      (17)
where $\{B_{t}^{(j)}:j=1,2,\dots p\}$ denotes $p$ independent Brownian motions. To obtain a high-fidelity approximation of the PrO posterior from this particle system, one discretises it into time steps $t_{1},t_{2},\dots$, and evolves it for a sufficiently long time. Afterwards, one averages over all time steps past an initial burn-in $\tau$ and approximates the PrO posterior as
                                                                                                    $Q_{n}\approx\frac{1}{|\{t_{i}:t_{i}>\tau\}|}\sum_{t_{i}>\tau}\widehat{Q}[t_{i}]\quad\text{ for }\quad\widehat{Q}[t_{i}]=\frac{1}{p}\sum_{j=1}^{p}\updelta_{\vartheta_{t_{i}}^{(j)}}.$                                                                                                      (18)
Here, the notation highlights that this approximation should be thought of as an average of $|\{t_{i}:t_{i}>\tau\}|$ measures $\widehat{Q}[t]$, each of which would recover $Q_{n}$ exactly if there were no discretisation error and both $p\to\infty$ and $t\to\infty$ (see e.g. Lemma 6 in Appendix E of Wild et al., 2023). We showcase this algorithm for the squared MMD and the logarithmic scoring rule in Appendix E.1, and elaborate on practical considerations in Appendix E.2.
7 Discussion
In this paper, we have advocated for a new statistical paradigm that defines epistemic uncertainty as a function of predictive capability. This led to PrO posteriors, which predictively dominate alternative methods for parameter inference like Gibbs posteriors and deliver interpretable and practically meaningful uncertainty quantification. We argued that PrO posteriors bridge the two main statistical traditions identified by Breiman (2001), and asserted that they lead to a version of epistemic uncertainty that is more expansive, adapts to model misspecification, and can avoid posterior collapse whenever this improves predictive performance. Throughout, we rigorously established these benefits and illustrated their practical importance on several instructive examples.
While we have presented evidence which advocates the adoption of PrO posteriors for inference and uncertainty quantification, much remains to be done. Perhaps the most pressing concern is the improvement of the available strategies for computation: sampling schemes obtained from interacting particle systems require further work before they can become attractive off-the-shelf algorithmic framework. Most notably, naïve implementations are infeasible for high-dimensional problems, generally computationally demanding, and require careful tuning of several hyperparameters. However, since research on improving such samplers has recently intensified (see Suzuki et al., 2023; Fu and Wilson, 2023; Liu et al., 2023; Chazal et al., 2025), we are hopeful that advances in the area will facilitate purpose-built algorithms which overcome these limitations.
Beyond computation, there are several outstanding theoretical questions that our current development has not explored yet. The most obvious of these concerns the choice of scoring rule. While our experiments and theory suggest that this choice is relatively unimportant in the case of convex recovery, we expect there to be important qualitative differences in behaviour for certain forms of non-trivial misspecification. For example, with Huber-contaminated data-generating processes of the form $P_{0}=(1-\varepsilon)P_{Q^{*}}+\varepsilon\cdot c$ for some $Q^{*}\in\mathcal{P}(\Theta)$ and some contamination $c\in\mathcal{P}(\mathcal{X})$, existing research on parametric inference (see e.g. Chérief-Abdellatif and Alquier, 2020, 2022; Dellaporta et al., 2022; Alquier et al., 2023; Alquier and Gerber, 2024) would indicate that PrO posteriors based on outlier-robust scoring rules like the squared maximum mean discrepancy should recover posteriors very similar to $Q^{*}$, while logarithmic scoring rules would behave completely differently. A second foundational question concerns the degree to which data can be allowed to be dependent: while our theory was derived under the assumption of independent data, our experiments seemed to suggest that the predictive benefits of PrO posteriors extend beyond this setting. Lastly, while it is clear that PrO posteriors are incoherent belief updates in the sense of Bissiri et al. (2016), it is worth investigating whether they can be re-cast as quasi-coherent updates on an extended space.
In the bigger picture, PrO posteriors have interesting connections to a number of other approaches grounded in frequentist statistics, and studying their intersection may result in mutually beneficial insights. For instance, the PrO posterior can be seen as a regularised frequentist estimate for the mixing measure in a nonparametric mixture model: most obviously, in the special case of the logarithmic scoring rule and when the prior regularisation term is dropped, (4) recovers nonparametric maximum likelihood exactly. Importantly, dropping the KL-term causes several practical issues that are absent from PrO posteriors, including computational difficulties due to non-identifiability (see e.g. Laird, 1978). Further, the corresponding measure-valued nonparametric maximum likelihood estimator generally cannot be interpreted as a tool for probabilistic uncertainty quantification—not only for conceptual reasons, but also because it is generally not unique and discrete; see Lindsay (1995, Theorem 21 in Chapter 5) and Jordan-Squire (2015).
Similarly, our work has clear parallels with the efforts of statistical learning theory to develop optimal aggregation procedures for predictors. For example, in his Saint-Flour lectures, Nemirovski (2000) introduces convex aggregation as the statistical task of finding the optimal convex combination of predictors. Intriguingly, for some model classes such as bounded regression with a finite number of candidates, the minimax-optimal convex aggregation procedure with respect to the squared risk can be shown to coincide precisely with (4) so long as the KL-regularisation term is removed (see, e.g., Lecué, 2013; Yang and Dunson, 2014). Though we have made no attempt at formalising these suspicions, this also suggests that PrO posteriors may be close to minimax-optimal convex aggregation procedures in certain problem classes.
Acknowledgements
This manuscript has benefited tremendously from interactions with many brilliant colleagues. Most notably, we feel deep gratitude towards Prof. Chris Oates for many conversations and insightful remarks. Further, we would like to thank Dr. Louis Sharrock and Dr. Sam Power for their help with the computational aspects of this paper. We are also grateful for correspondence with Prof. Pierre Alquier that pointed to the connection with PAC-Bayesian majority voting. Additionally, we owe a tremendous debt of gratitude to several colleagues who took time to read an earlier version of this manuscript and gave incredibly insightful comments—often pages and pages of them!—that significantly improved the paper; most notably Prof. David Dunson, Prof. Jeff Miller, Prof. Aki Vehtari, Prof. Pierre Jacob, Dr. Andres Masegosa, Dr. Edwin Fong, Dr. Jack Jewson, and Dr. Takuo Matsubara. DTF acknowledges funding from the Australian Research Council (DE200101070, DP200101414), YM from the Heilbronn Institute for Mathematical Research (EP/V521917/1), and JK was supported through the UK’s Engineering and Physical Sciences Research Council (EPSRC) via EP/W005859/1. The authors acknowledge the use of the UCL Myriad High Performance Computing Facility (Myriad@UCL), and associated support services, in the completion of this work. Data were provided by the UK National River Flow Archive.
Appendix
Appendix A Notations
In this section we re-describe the main notations introduced in the main text to ensure the appendix is as self-contained as possible. We consider observed data ${x}_{1:n}=(x_{1},\dots,x_{n})$ for $x_{i}\in\mathcal{X}$ and $i=1,\dots,n$, where each $x_{i}$ is generated from the true unknown distribution $P_{0}$. We denote expectations of functions $g:\mathcal{X}\to\mathbb{R}$ with respect to $X\sim P_{0}$ by $\mathbb{E}[g(X)]$ or by $\mathbb{E}_{X\sim P_{0}}[g(X)]$ if we wish to emphasise the distribution over which integration occurs. The model class is given by $\mathcal{M}_{\Theta}:=\{P_{\theta}:\theta\in\Theta\}$ and a prior measure $\Pi\in\mathcal{P}(\Theta)$ over elements of $\mathcal{M}_{\Theta}$. Likewise, let $\mathcal{P}(\Theta)$ denote the set of distributions over $\Theta$, and recall $P_{Q}=\int P_{\theta}\mathsf{d}Q(\theta)$ for $Q\in\mathcal{P}(\Theta)$. The function $S(P,x)$ denotes a scoring rule for the distribution $P$, while $\mathcal{S}(P,P)=\mathbb{E}_{X\sim P}[S(P,X)]$ denotes the expected score, and $\mathcal{D}_{S}(P,P_{0})=\mathcal{S}(P,P_{0})-\mathcal{S}(P_{0},P_{0})$ is the divergence associated to $P$, evaluated at the true distribution $P_{0}$. Further, define $S_{n}(P_{\theta})=\frac{1}{n}\sum_{i=1}^{n}S(P_{\theta},x_{i})$, and $L_{n}(\theta_{1:k})=\frac{1}{n}\sum_{i=1}^{n}L(\theta_{1:k},x_{i})$, where $L(\theta_{1:k},x)$ is as defined in Assumption 2.
Appendix B Lemmas
This section contains several useful lemmas that are used throughout the remainder of the text. First, we demonstrate our claim that PrO posterior based on any kernel scoring rule are tractable with $k=2$. This implies that for scores based on squared maximum mean discrepancy (MMD2) and Cumulative Ranked Probability Score (CRPS), the function $\Delta(Q,x)$ in (3.1) is constant in $x$.
To prove this, take $x,x^{\prime}\in\mathcal{X}$, let $\kappa(x,x^{\prime})$ denote a positive-definite kernel function with feature map $\phi:\mathcal{X}\to\mathcal{H}$, such that $\kappa(x,x^{\prime})=\langle\phi(x),\phi(x^{\prime})\rangle_{\mathcal{H}}$. For fixed $\theta\in\Theta$, define the mean embedding $\mu_{\theta}\in\mathcal{H}$ by $\mu_{\theta}:=\int\phi(z)\,\mathsf{d}P_{\theta}(z)$ and the inner product $\langle\mu_{\theta},\mu_{\theta^{\prime}}\rangle_{\mathcal{H}}=\int\!\!\int\kappa(z,z^{\prime})\,\mathsf{d}P_{\theta}(z)\mathsf{d}P_{\theta^{\prime}}(z^{\prime}).$ Following Gneiting et al. (2007), all kernel scoring rules can be expressed as $S(P,x)=\mathbb{E}_{X,X^{\prime}\sim P}[\kappa(X,X^{\prime})]-2\mathbb{E}_{X\sim P}[\kappa(X,x)]+\kappa(x,x).$
Lemma 2.
For any kernel scoring rule, and for all $x\in\mathcal{X}$: (i) the score $P_{\theta}\mapsto S(P_{\theta},x)$ is convex; (ii) $S(P_{Q},x)=\int L(\theta_{1:2},x)\mathsf{d}Q^{2}(\theta_{1:2})$ with $L(\theta_{1:2},x)=S(P_{\theta_{1}},x)-\frac{1}{2}\|\mu_{\theta_{1}}-\mu_{\theta_{2}}\|^{2}_{\mathcal{H}}.$
Proof.
To see that (i) is satisfied, use the feature map representation to write the score as
$S(P_{\theta},x)=\|\mu_{\theta}\|_{\mathcal{H}}^{2}-2\langle\mu_{\theta},\phi(x)\rangle_{\mathcal{H}}+\|\phi(x)\|_{\mathcal{H}}^{2}.$
The map $P_{\theta}\mapsto\mu_{\theta}$ is linear, so that $P_{\theta}\mapsto\|\mu_{\theta}\|_{\mathcal{H}}^{2}$ is quadratic. Thus, $P_{\theta}\mapsto S(P_{\theta},x)$ is the composition of a linear map and a convex function, and so is convex. To deduce the second result, write
[…]
Expand the first integral, and use the mean embedding definition for $\mu_{\theta}$, to write
$S\left(P_{Q},x\right)=\int\int\int\int\kappa(z,z^{\prime})\,\mathsf{d}P_{\theta}(z)\mathsf{d}P_{\theta^{\prime}}(z^{\prime})\mathsf{d}Q(\theta)\mathsf{d}Q(\theta^{\prime})$ $=\langle\mathbb{E}_{\theta\sim Q}(\mu_{\theta}),\mathbb{E}_{\theta^{\prime}\sim Q}(\mu_{\theta^{\prime}})\rangle$
$=\|\mathbb{E}_{\theta\sim Q}(\mu_{\theta})\|^{2}_{\mathcal{H}}.$
Similarly, write
$\int S\left(P_{\theta},x\right)\mathsf{d}Q(\theta)$ $=\int\int\int\kappa(z,z^{\prime})\,\mathsf{d}P_{\theta}(z)\mathsf{d}P_{\theta}(z^{\prime})\mathsf{d}Q(\theta)=\mathbb{E}_{\theta\sim Q}[\|\mu_{\theta}\|^{2}_{\mathcal{H}}].$
Hence, for $\Delta(Q,x)$ as in (3.1),
$-\Delta(Q,x)=\|\mathbb{E}_{\theta\sim Q}[\mu_{\theta}]\|^{2}_{\mathcal{H}}-\mathbb{E}_{\theta\sim Q}[\|\mu_{\theta}\|^{2}_{\mathcal{H}}]$ $=-\mathbb{E}_{\theta\sim Q}\left[\|\mu_{\theta}-\mathbb{E}_{\theta^{\prime}\sim Q}[\mu_{\theta^{\prime}}]\|^{2}_{\mathcal{H}}\right]$
$=-\frac{1}{2}\int\|\mu_{\theta}-\mu_{\theta^{\prime}}\|^{2}_{\mathcal{H}}\mathsf{d}Q(\theta)\mathsf{d}Q(\theta^{\prime}).$
Lemma 3.
Under Assumptions 1-3, for any $\lambda_{n}\geq d_{S}/r_{0}$, we have
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}Q_{n}^{\dagger}(\theta)\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\frac{d_{S}}{\lambda_{n}}\,\log\left(\dfrac{e\cdot c_{S}\lambda_{n}}{d_{S}}\right)+\frac{\lambda_{n}C^{2}}{n}\,.$
Furthermore, for the specific choice $\lambda_{n}=n^{1/2}\,\sqrt{{d_{S}}/{C^{2}}\,\log\left({e\cdot c_{S}}/{d_{S}}\right)}$, for any $n\,\geq\frac{d_{S}C^{2}}{r_{0}^{2}\,\log\left(\frac{e\cdot c_{S}}{d_{S}}\right)}$, we have
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}^{\dagger}},P_{0}\right)\right]\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+C\left(\log n/n\right)^{1/2}.$
Proof.
The proof follows similar arguments to those used to prove Lemma 1 but where we take $L(\theta,\theta^{\prime})=R(\theta)=\mathcal{S}(P_{\theta},P_{0})-\mathcal{S}(P_{0},P_{0})$ and $L_{n}(\theta,\theta^{\prime})=R_{n}(\theta)=S_{n}(P_{\theta})-S_{n}(P_{0}).$ If Assumption 2 is satisfied for $R(\theta)-R_{n}(\theta)$, then
$1\geq\mathbb{E}\int e^{\lambda_{n}\left\{R(\theta)-R_{n}(\theta)\right\}-\frac{\lambda_{n}^{2}C^{2}}{n}}\mathsf{d}\Pi(\theta).$
Apply the Donsker-Varadhan lemma with $h(\theta)=R(\theta)-R_{n}(\theta)$ to obtain
$1\geq\mathbb{E}e^{\sup_{Q\in\mathcal{P}(\Theta)}\lambda_{n}\int\left[R(\theta)-R_{n}(\theta)\right]\mathsf{d}Q(\theta)-\mathrm{KL}(Q\|\Pi)-\frac{\lambda_{n}^{2}C^{2}}{n}}.$
Apply Jensen’s inequity and take logs to obtain
$0\geq\mathbb{E}\sup_{Q\in\mathcal{P}(\Theta)}\left\{\lambda_{n}\int\left[R(\theta)-R_{n}(\theta)\right]\mathsf{d}Q(\theta)-\mathrm{KL}(Q\|\Pi)-\frac{\lambda_{n}^{2}C^{2}}{n}\right\}.$
Since the above is true for any $Q\in\mathcal{P}(\Theta)$, it is also true for $Q=Q^{\dagger}_{n}$. Applying this choice and re-arranging terms yields
$\mathbb{E}\int R(\theta)\mathsf{d}Q^{\dagger}_{n}(\theta)\leq\mathbb{E}\left\{\int R_{n}(\theta)\mathsf{d}Q^{\dagger}_{n}(\theta)+\frac{\mathrm{KL}({Q^{\dagger}_{n}\|\Pi})}{\lambda_{n}}+\frac{\lambda_{n}C^{2}}{n}\right\}.$
Use the definitions of $R(\theta)$ and $R_{n}(\theta)$, and Assumption 1, to obtain the inequality
$\mathbb{E}\left[\mathcal{S}(P_{Q_{n}^{\dagger}},P_{0})-\mathcal{S}(P_{0},P_{0})\right]$
$\leq$ $\mathbb{E}\left[\int\{\mathcal{S}(P_{\theta},P_{0})-\mathcal{S}(P_{0},P_{0})\}\mathsf{d}Q^{\dagger}_{n}(\theta)\right]$ (20)
$\leq$ $\mathbb{E}\left\{\int\left\{S_{n}(P_{\theta})-S_{n}(P_{0})\right\}\mathsf{d}Q^{\dagger}_{n}(\theta)+\frac{\mathrm{KL}({Q^{\dagger}_{n}\|\Pi})}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}$
$\leq$ $\mathbb{E}\inf_{Q\in\mathcal{P}(\Theta)}\left\{\int\left\{S_{n}(P_{\theta})-S_{n}(P_{0})\right\}\mathsf{d}Q(\theta)+\frac{\mathrm{KL}({Q\|\Pi})}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}$
$=$ $\mathbb{E}\inf_{Q\in\mathcal{P}(\Theta)}\left\{\int\left\{S_{n}(P_{\theta})-S_{n}(P_{0})\right\}\mathsf{d}Q(\theta)+\frac{\mathrm{KL}({Q\|\Pi})}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}.$ (21)
Define the distributions $\Pi_{\mathcal{B}_{r}}$ indexed by a radius $r>0$ with corresponding density:
$\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta):=\frac{1(\theta\in\mathcal{B}_{r})\mathsf{d}\Pi(\theta)}{\Pi(\mathcal{B}_{r})}\quad\textnormal{with}\quad\mathcal{B}_{r}:=\left\{\theta\in\Theta:\mathcal{D}_{S}(P_{\theta},P_{0})\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+r\right\}.$
We can now further upper bound (20) by restricting the infimum on the right-hand side of (21) to distributions belonging to the set $\{\Pi_{\mathcal{B}_{r}}:r>0\}\subset\mathcal{P}(\Theta)$, which yields
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}Q_{n}^{\dagger}(\theta)$ $\leq\mathbb{E}\inf_{r>0}\left\{\int\left\{S_{n}(P_{\theta})-S_{n}(P_{0})\right\}\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\mathrm{KL}({\Pi_{\mathcal{B}_{r}}\|\Pi})}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}$
$\leq\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\mathrm{KL}({\Pi_{\mathcal{B}_{r}}\|\Pi})}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}$
$=\inf_{r>0}\left\{\mathop{\text{\large$\int$}}\nolimits\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\log\left\{1/\Pi(\mathcal{B}_{r})\right\}}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}$
$\leq\inf_{0<r\leq r_{0}}\left\{\mathop{\text{\large$\int$}}\nolimits\left\{\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+r\right\}\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{d_{S}\,\log\left(c_{S}/r\right)}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}$
$=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\inf_{0<r\leq r_{0}}\left\{r+\frac{d_{S}\,\log\left(c_{S}/r\right)}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}\,.$
As soon as $\lambda_{n}\geq\frac{d_{S}}{r_{0}}$, taking $r=\frac{d_{S}}{\lambda_{n}}\leq r_{0}$ then leads to
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}Q_{n}^{\dagger}(\theta)$ $\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\frac{d_{S}}{\lambda_{n}}+\frac{d_{S}}{\lambda_{n}}\,\log\left(\dfrac{c_{S}\lambda_{n}}{d_{S}}\right)+\frac{\lambda_{n}C^{2}}{n}$
$=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\frac{d_{S}}{\lambda_{n}}\,\log\left(\dfrac{e\cdot c_{S}\lambda_{n}}{d_{S}}\right)+\frac{\lambda_{n}C^{2}}{n}\,$,
which establishes the first result.
To obtain the second, substitute $\lambda_{n}=n^{1/2}\,\sqrt{\frac{d_{S}}{C^{2}}\,\log\left(\frac{e\cdot c_{S}}{d_{S}}\right)}$ in the bound, for $n\,\geq\frac{d_{S}C^{2}}{r_{0}^{2}\,\log\left(\frac{e\cdot c_{S}}{d_{S}}\right)}$, to obtain
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}Q_{n}^{\dagger}(\theta)$
$\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\frac{C\sqrt{d_{S}}}{n^{1/2}}\frac{\log\left(n^{1/2}\,\dfrac{e\cdot c_{S}}{C\sqrt{d_{S}}}\sqrt{\log\left(\frac{e\cdot c_{S}}{d_{S}}\right)}\right)}{\sqrt{\log\left(\frac{e\cdot c_{S}}{d_{S}}\right)}}+\frac{C\sqrt{d_{S}}}{n^{1/2}}\sqrt{\log\left(\frac{e\cdot c_{S}}{d_{S}}\right)}\,$,
and by convexity, Assumption 1 implies that $\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]\leq\mathbb{E}\mathop{\text{\large$\int$}}\nolimits\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}Q_{n}^{\dagger}(\theta),$ which ends the proof. ∎
B.1 Existence and uniqueness of the minimiser
It is interesting to investigate the sufficient conditions under which the objective of (6) admits a unique minimiser. We posit two assumptions in order to do so: one on the score and the other on the prior.
Lemma 4.
Suppose that Assumption 1 holds and further that the score is bounded from below so that $\inf_{Q\in\mathcal{P}(\Theta)}\mathscr{L}(Q)=M>$-$\infty$. Assume further that the prior $\Pi$ admits density with respect to the Lebesgue measure $\mathsf{d}\Pi(\theta)\propto\exp\{-U(\theta)\}$ for $U:\Theta\to\mathbb{R}$ such that:
  1. (i)
    there exist constants $c>0$ and $c^{\prime}\in\mathbb{R}$ such that for all $\theta\in\Theta$,
    $\nabla_{\theta}U(\theta)\cdot\theta\geq c\lVert\theta\rVert^{2}+c^{\prime},$
    and,
  2. (ii)
    $\nabla_{\theta}U$ is Lipschitz continuous.
Then there exists a unique measure $Q_{n}\in\mathcal{P}(\Theta)$ such that $Q_{n}=\operatornamewithlimits{arginf}_{Q\in\mathcal{P}(\Theta)}\mathscr{L}$, and further this $Q_{n}$ is characterised by the implicit equation
$\mathsf{d}Q_{n}(\theta)\propto\mathsf{d}\Pi(\theta)\cdot\exp\left\{\left.-\frac{\lambda_{n}}{n}\sum_{i=1}^{n}\frac{\delta S(P_{Q},x_{i})}{\delta Q}\right|_{Q=Q_{n}}(\theta)\right\}.$
Proof.
We have by Assumption 1 that the score $S(\cdot,\,x)$ is convex is its first argument. In combination with the fact that the mapping $Q\mapsto P_{Q}$ is linear, we have the the mapping $Q\mapsto S(P_{Q},\,x)$ is convex. The existence and uniqueness of a minimiser is then provided under the same conditions in Proposition 2.5 of Hu et al. (2020), and we point the reader to their Section 4 for a proof.
To characterise the form of this minimiser, we take the functional derivative of (6) with respect to $Q$, set it equal to zero, and solve for $Q$. We begin by noting that the functional derivative of the KL regulariser evaluated at $\theta$ is given by
$1+\log\left\{\frac{\mathsf{d}Q}{\mathsf{d}\Pi}(\theta)\right\},$ (22)
while that of the score term is
$\frac{\lambda_{n}}{n}\sum_{i=1}^{n}\frac{\delta S(P_{Q},x_{i})}{\delta Q}(\theta).$ (23)
By the linearity of functional derivatives, setting the sum of (22) and (23) equal to zero and removing constant terms, we find that $Q_{n}$ must satisfy
$0=\frac{\lambda_{n}}{n}\sum_{i=1}^{n}\left.\frac{\delta S(P_{Q},x_{i})}{\delta Q}\right|_{Q=Q_{n}}(\theta)+\log\left\{\frac{\mathsf{d}Q_{n}}{\mathsf{d}\Pi}(\theta)\right\}$
whence rearranging yields the stated result. ∎
The prior coercivity conditions of i and ii are satisfied, for instance, by choosing the prior to be a multivariate standard Gaussian prior, common to many applications (Hu et al., 2020). And the condition that the infimum be attained at non-infinite $M$ is immediate, for example, for the squared MMD under bounded kernels (see also Shen et al., 2025, Theorem 1). In the case of the log score, we refer the reader to Liu et al. (2025, Proposition 1) who establish the necessary conditions.
Appendix C Proofs of Main results
Proof of Lemma 1.
The result depends on slightly different arguments depending on whether Assumption 2 or Assumption 2 is satisfied. Therefore, we break the result into cases.
Scores Satisfying Assumption 2. Consider any fixed integer $k>1$. Recall $Q^{k}=\bigotimes_{j=1}^{k}Q$. Apply the Donsker-Varadhan Lemma with $h(\theta_{1:k})=k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}$, to see that, $P_{0}$-almost surely, $\forall Q\in\mathcal{P}(\Theta)$, we have
$\mathop{\text{\large$\int$}}\nolimits k\lambda_{n}\left\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\right\}\mathsf{d}Q^{k}(\theta_{1:k})\leq\text{KL}(Q^{k}\|\Pi^{k})+{\log}\mathop{\text{\large$\int$}}\nolimits e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k}).$
Rearranging terms, we have $P_{0}$-almost surely, $\forall Q\in\mathcal{P}(\Theta)$,
$\mathop{\text{\large$\int$}}\nolimits L(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})\leq\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}+\frac{{\log}\mathop{\text{\large$\int$}}\nolimits e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k})}{k\lambda_{n}},$
and $P_{0}$-almost surely, for $Q=Q_{n}$ in particular:
$\mathop{\text{\large$\int$}}\nolimits L(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})\leq\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})+\frac{\text{KL}(Q_{n}^{k}\|\Pi^{k})}{k\lambda_{n}}+\frac{{\log}\mathop{\text{\large$\int$}}\nolimits e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k})}{k\lambda_{n}}.$
Taking expectation with respect to $P_{0}$ on both sides and applying Jensen’s inequality to the log then leads to:
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})$ $\leq\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})+\frac{\mathbb{E}\text{KL}(Q_{n}^{k}\|\Pi^{k})}{k\lambda_{n}}+\frac{\mathbb{E}\log\mathop{\text{\large$\int$}}\nolimits e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k})}{k\lambda_{n}}$
$\leq\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})+\frac{\mathbb{E}\text{KL}(Q_{n}^{k}\|\Pi^{k})}{k\lambda_{n}}+\frac{\log\mathbb{E}\mathop{\text{\large$\int$}}\nolimits e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k})}{k\lambda_{n}}.$
The last term on the right-hand-side is controlled using Tonelli’s theorem and Assumption 2(ii), under which:
$\log\mathbb{E}\mathop{\text{\large$\int$}}\nolimits e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k})=\log\mathop{\text{\large$\int$}}\nolimits\mathbb{E}e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k})\leq\frac{k^{2}\lambda_{n}^{2}C^{2}}{n},$
which then gives
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})$ $\leq\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})+\frac{\mathbb{E}\text{KL}(Q_{n}^{k}\|\Pi^{k})}{k\lambda_{n}}+\frac{\lambda_{n}kC^{2}}{n}$
$=\mathbb{E}\left[\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})+\frac{\text{KL}(Q_{n}^{k}\|\Pi^{k})}{k\lambda_{n}}\right]+\frac{\lambda_{n}kC^{2}}{n}.$
Recall that $Q_{n}$ is the minimiser of the quantity inside the expectation on the right-hand-side to finally write:
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})\leq\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC^{2}}{n}.$
Now, we lower bound $\mathop{\text{\large$\int$}}\nolimits L(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})$ on the left-hand-side: using the convexity in Assumption 1 and the structure of the loss in Assumption 2(i), we have
$\mathcal{S}\left(\int P_{\theta}\mathsf{d}Q_{n}(\theta),P_{0}\right)$ $=\int\mathcal{S}(P_{\theta},P_{0})\mathsf{d}Q_{n}(\theta)-\mathbb{E}\{\Delta(Q_{n},X)\}$
$\leq\int\mathcal{S}(P_{\theta},P_{0})\mathsf{d}Q_{n}(\theta)-\int\delta(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})\equiv\int L(\theta_{1:k})\mathsf{d}Q_{n}(\theta_{1:k}),$
where $\delta(\theta_{1:k})=\mathbb{E}[\delta(\theta_{1:k},X)]$. Recalling that $P_{Q_{n}}=\int P_{\theta}\mathsf{d}Q_{n}(\theta)$, this gives us
$\mathbb{E}\left[\mathcal{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq\mathbb{E}\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}kC^{2}}{n}.$ (24)
Using Assumption 2(i), we can rewrite equation (24) as
$\mathbb{E}\left\{\mathcal{S}\left(P_{Q_{n}},P_{0}\right)\right\}$ $\leq\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC^{2}}{n}$
$=\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{S_{n}(P_{Q})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC^{2}}{n}$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{S}(P_{Q},P_{0})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}kC^{2}}{n}.$
Recalling that $\mathcal{S}(P_{0},P_{0})=\mathbb{E}S_{n}(P_{0})$, we subtract $\mathcal{S}(P_{0},P_{0})$ from both sides and write
$\mathbb{E}\{\mathcal{D}_{S}(P_{Q_{n}},P_{0})\}$ $=\mathbb{E}\left\{\mathcal{S}\left(P_{Q_{n}},P_{0}\right)-\mathcal{S}(P_{0},P_{0})\right\}$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{S}(P_{0},P_{0})-\mathcal{S}(P_{Q},P_{0})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}kC^{2}}{n}$
$=\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}(P_{Q},P_{0})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}kC^{2}}{n}$
$=\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}(P_{Q},P_{0})+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC^{2}}{n}.$
Scores Satisfying Assumption 2. For scores that do not satisfy Assumption 2, the above proof technique breaks down. We can, however, derive a similar bound using the variational definition of $Q_{n}$.
From the definition of $\mathcal{D}(P_{Q},P_{0})$, we have
$\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{0})]$ $=\mathbb{E}[\mathcal{S}(P_{Q_{n}},P_{0})]-\mathcal{S}(P_{0},P_{0})$ (25)
$=\mathbb{E}[{S}_{n}(P_{Q_{n}})]-\mathcal{S}(P_{0},P_{0})-\left\{\mathbb{E}[{S}_{n}(P_{Q_{n}})]-\mathbb{E}[\mathcal{S}(P_{Q_{n}},P_{0})]\right\}$
$\leq\mathbb{E}[{S}_{n}(P_{Q_{n}})]-\mathcal{S}(P_{0},P_{0})+\mathbb{E}\sup_{Q\in\mathcal{P}_{\Pi}(\Theta)}|S_{n}(P_{Q})-\mathcal{S}(P_{Q},P_{0})|$
$\leq\mathbb{E}[{S}_{n}(P_{Q_{n}})]-\mathcal{S}(P_{0},P_{0})+r_{n},$ (26)
where the last line follows from Assumption 2. Using the variational definition of $Q_{n}$, we have
$S_{n}(P_{Q_{n}})\leq S_{n}(P_{Q_{n}})+\frac{\mathrm{KL}(Q_{n}\|\Pi)}{\lambda_{n}}=\inf_{Q\in\mathcal{P}(\Theta)}\left\{S_{n}(P_{Q})+\frac{\mathrm{KL}(Q\|\Pi)}{\lambda_{n}}\right\}.$
Taking expectations then yields
$\mathbb{E}\left\{S_{n}(P_{Q_{n}})\right\}$ $\leq\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{S_{n}(P_{Q})+\frac{\mathrm{KL}(Q\|\Pi)}{\lambda_{n}}\right\}\right]\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{S}(P_{Q},P_{0})+\frac{\mathrm{KL}(Q\|\Pi)}{\lambda_{n}}\right\}.$ (27)
Plugging equation (27) into (26) yields
$\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{0})]$ $=\mathbb{E}[\mathcal{S}(P_{Q_{n}},P_{0})]-\mathcal{S}(P_{0},P_{0})$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{S}(P_{Q},P_{0})+\frac{\mathrm{KL}(Q\|\Pi)}{\lambda_{n}}\right\}-\mathcal{S}(P_{0},P_{0})+r_{n}$
$=\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}(P_{Q},P_{0})+\frac{\mathrm{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+r_{n}.$
The stated result now follows from the definition of $\nu_{n}$. ∎
Proof of Theorem 2.
From the convexity in Assumption 1,
$\mathcal{D}_{S}(P_{Q},P_{0})\leq\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}Q(\theta).$
Applying bound in the RHS of Lemma 1 delivers
$\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{0})]$ $\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}(P_{Q},P_{0})+\frac{\mathrm{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}Q(\theta)+\frac{\mathrm{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}.$ (28)
Recalling the definition of $\mathcal{B}_{r}$ in Assumption 3, define the candidate density
$\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta):=\frac{1(\theta\in\mathcal{B}_{r})\mathsf{d}\Pi(\theta)}{\Pi(\mathcal{B}_{r})}.$
Now, the RHS of the inequality in (28) is true for any $Q\in\mathcal{P}(\Theta)$, and remains true for $Q=\Pi_{\mathcal{B}_{n}}$. For $0\leq r\leq r_{0}$, over $\mathcal{B}_{r}$, $\mathcal{D}_{S}(P_{\theta},P_{0})\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+r$, so that we obtain
$\mathbb{E}\left[\mathcal{D}_{S}(P_{Q_{n}},P_{0})\right]$ $\leq\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\mathrm{KL}(\Pi_{\mathcal{B}_{r}}^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\nu_{n},$
$=\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\mathrm{KL}(\Pi_{\mathcal{B}_{r}}\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}$
$=\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\log\left\{1/\Pi(\mathcal{B}_{r})\right\}}{\lambda_{n}}\right\}+\nu_{n},$
$\leq\inf_{r_{0}\geq r>0}\left\{\int\left\{\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+r\right\}\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{d_{S}\log(c_{S}/r)}{\lambda_{n}}\right\}+\nu_{n}$
$=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\inf_{r_{0}\geq r>0}\left\{r+\frac{d_{S}\log(c_{S}/r)}{\lambda_{n}}\right\}+\nu_{n}\,.$
The remainder of the proof follows the same as in the proof of Lemma 3.
A similar argument holds for $Q_{n}^{\dagger}$.
Proof of Theorem 3.
Assumption 4 ensures the existence of some $Q^{\star}\in\arg\inf_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{S}\left(P_{Q},P_{0}\right)$ such that
${\mathsf{JenGap}}:=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\mathcal{D}_{S}\left(P_{Q^{\star}},P_{0}\right)>0.$
From Lemma 1, we can obtain:
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$ $\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}$
$=\inf_{Q\ll\Pi}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}$
$\leq\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q^{\star}(\theta),P_{0}\right)+\frac{\textnormal{KL}({Q}^{\star}\|\Pi)}{\lambda_{n}}+\nu_{n}.$ (29)
Adding and subtracting $\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})$ and using the definition of $\mathsf{JenGap}$, we can re-arrange (29) as
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$ $\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\frac{\textnormal{KL}({Q}^{\star}\|\Pi)}{\lambda_{n}}+\nu_{n}-\mathsf{JenGap}.$
Since, by Assumption 4, $Q^{\star}$ is such that $\mathrm{KL}(Q^{\star}\|\Pi)<\infty$, for $n$ large enough we have
$\frac{\textnormal{KL}({Q}^{\star}\|\Pi)}{\lambda_{n}}+\nu_{n}\leq\frac{1}{4}{\mathsf{JenGap}}\,$,
so that
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\frac{3}{4}{\mathsf{JenGap}}\,.$
Furthermore, from the continuity of $\theta\mapsto P_{\theta}$ in Assumption 4, and the concentration in Lemma 3, we have for $n$ large enough but finite,
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}^{\dagger}(\theta),P_{0}\right)\right]\geq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\frac{\mathsf{JenGap}}{4}\,.$
Therefore, for $n$ large enough but finite,
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$ $\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\frac{3}{4}{\mathsf{JenGap}}$
$\leq\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}^{\dagger}(\theta),P_{0}\right)\right]-\frac{1}{2}{\mathsf{JenGap}}$
$<\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}^{\dagger}(\theta),P_{0}\right)\right].$
To obtain the second result, return to equation (29), and use the fact that $Q^{\star}\in\operatornamewithlimits{arginf}_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{S}(P_{Q},P_{0})$ satisfies $\mathrm{KL}(Q^{\star}\|\Pi)<\infty$, by Assumption 4, to obtain
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$ $\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}$
$\leq\mathcal{D}_{S}(P_{Q^{\star}},P_{0})+\frac{\text{KL}(Q^{\star}\|\Pi)}{\lambda_{n}}+\nu_{n}$
$\leq\mathcal{D}_{S}(P_{Q^{\star}},P_{0})+\frac{C}{\lambda_{n}}+\nu_{n}.$
In the case of Assumption 2, $\nu_{n}=\log(n)/{n}^{1/2}$ so that choosing $\lambda_{n}\propto n^{1/2}/\log(n)$ yields the stated result. In the case of Assumption 2, $\nu_{n}=\min\{\log(n)/{n}^{1/2},r_{n}\}$, and again choosing $\lambda_{n}\propto{n}^{1/2}\log(n)$ also implies the stated result. ∎
Proof of Theorem 4.
We break the result up into cases depending on whether (i) or (ii) in Assumption 5 is satisfied.
Assumption 5(i). From equation (29) in the proof of Theorem 3, we have that
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]$ $\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}$
$\leq\inf_{Q\ll\Pi}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}$
$=\mathcal{D}_{S}(P_{Q^{\star}},P_{0})+\frac{C}{\lambda_{n}}+\nu_{n},$
where the equality follows from the fact that $\mathrm{KL}(Q^{\star}\|\Pi)<\infty$, and the last inequality follows by taking $\lambda_{n}\propto{n}^{1/2}/\log(n)$ so that $C/\lambda_{n}+\nu_{n}\leq C^{\prime}\cdot\nu_{n}$, for $n$ large enough. We remind the reader that $\mathcal{D}_{S}(P_{Q^{\star}},P_{0})=0$ under Assumption 5(i).
Since $\mathcal{D}_{S}(P_{Q^{\star}},P_{0})=0$ under convex recovery, we have
${\mathsf{JenGap}}=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})>0,$
and there exist an $n$ large enough so that
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq C^{\prime}\nu_{n}\leq\frac{\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})}{2}.$
Further, for $n$ large enough, as in the proof of Theorem 3, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}^{\dagger}},P_{0}\right)\right]>\frac{\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})}{2}.$ Putting these inequalities together, we see that, for $n$ large enough,
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq\frac{\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})}{2}<\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}^{\dagger}},P_{0}\right)\right].$
Assumption 5(ii). Recall $\mathcal{B}_{r,j}=\{\theta:\mathcal{D}_{S}(P_{\theta},P_{\theta_{j}^{\star}})\leq r\}$, denote $\mathsf{d}Q_{r}^{\star}(\theta)=\sum_{j=1}^{K}\omega_{j}\mathsf{d}\Pi^{\star}_{r,j}(\theta)$ where $\omega\in\Delta^{K-1}$, and
$\mathsf{d}\Pi^{\star}_{r,j}(\theta)=\begin{cases}\frac{\mathsf{d}\Pi(\theta)}{\Pi(\mathcal{B}_{r,j})}&\text{ if }\theta\in\mathcal{B}_{r,j}\\ 0&\text{ else }\end{cases}.$
Using $Q^{\star}_{r}$ within equation (29) in the proof of Theorem 3, yields
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\inf_{r>0}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q^{\star}_{r}(\theta),P_{0}\right)+\frac{\textnormal{KL}(Q^{\star}_{r}\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}$
$=\inf_{r>0}\left\{\mathcal{D}_{S}\left(\sum_{j=1}^{K}\omega_{j}\int P_{\theta}\,\mathsf{d}\Pi^{\star}_{r,j}(\theta),\sum_{j=1}^{K}\omega_{j}P_{\theta_{j}^{\star}}\right)+\frac{\textnormal{KL}\left(\sum_{j=1}^{K}\omega_{j}\Pi^{\star}_{r,j}\|\Pi\right)}{\lambda_{n}}\right\}+\nu_{n}$
$\leq\inf_{r>0}\left\{\sum_{j=1}^{K}\omega_{j}\,\int\,\mathcal{D}_{S}\left(P_{\theta},P_{\theta_{j}^{\star}}\right)\mathsf{d}\Pi^{\star}_{r,j}(\theta)+\frac{\sum_{j=1}^{K}\omega_{j}\textnormal{KL}\left(\Pi^{\star}_{r,j}\|\Pi\right)}{\lambda_{n}}\right\}+\nu_{n}$
$\leq\inf_{r>0}\left\{\sum_{j=1}^{K}\omega_{j}\,\int\,\mathcal{D}_{S}\left(P_{\theta},P_{\theta_{j}^{\star}}\right)\mathsf{d}\Pi^{\star}_{r,j}(\theta)+\frac{\sum_{j=1}^{K}\omega_{j}\log\{1/\Pi(\mathcal{B}_{r,j})\}}{\lambda_{n}}\right\}+\nu_{n}$
$\leq\inf_{r>0}\left\{\sum_{j=1}^{K}\omega_{j}\,\int\,\mathcal{D}_{S}\left(P_{\theta},P_{\theta_{j}^{\star}}\right)\mathsf{d}\Pi^{\star}_{r,j}(\theta)+\frac{d_{s}\log(c_{s}/r)}{\lambda_{n}}\right\}+\nu_{n}$
$\leq\inf_{r>0}\left\{r+\frac{d_{s}\log(c_{s}/r)}{\lambda_{n}}\right\}+\nu_{n}.$
Optimising $r$ using similar arguments to the end of Lemma 3 yields the stated result. ∎
Proof of Corollary 2.
The proof follows directly from Theorem 4; in particular,
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]-\mathbb{E}[\mathcal{D}_{S}(P_{Q^{\dagger}_{n}},P_{0})]\leq\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\lesssim\nu_{n}\,.$
Proof of Corollary 3 .
Since $\mathsf{d}_{2}(\cdot,\cdot)$ satisfies a weak triangle inequality, for some $C>0$,
                                                                                                    $\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}},P_{Q_{\star}})]$ ${}\leq{}$ $C\mathbb{E}\{\mathsf{d}_{2}(P_{Q_{n}},P_{0})+\mathsf{d}_{2}(P_{Q_{\star}},P_{0})\}$                                                                                                      (30)
=Cd_2(P_Q_n,P_0)(31) =Cd_2(P_Q_n,P_Q_⋆) , (32) where the equalities follows from the definition of convex recovery. Recall that, by Theorem 4, $\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{0})]\leq\nu_{n}$. Hence, from Assumption 6(iii), and concavity of the map $x\mapsto x^{1/\alpha}$ for $\alpha\geq 1$, $\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}},P_{Q_{\star}})]\leq$ $\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{Q_{\star}})^{1/\alpha}]\leq\left\{\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{Q_{\star}})]\right\}^{1/\alpha}\leq\nu_{n}^{1/\alpha}.$ (33) From equation (33), and Assumption 6(i) - identifiability and continuity of $\theta\mapsto P_{\theta}$ - conclude that $P_{Q_{n}}\rightarrow P_{0}$. Hence, by Assumption 6(ii) and the convergence in (33), there exists $n_{1}$ large enough so that, for $n\geq n_{1}$, E[d_1(Q_n,Q_⋆)]≤CE[d_2(P_Q_n,P_Q_⋆)]. Hence, for $n\geq n_{1}$, we can apply the above and equation (33) to obtain $\mathbb{E}[\mathsf{d}_{1}(Q_{n},Q_{\star})]$ $\leq C\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}},P_{Q_{\star}})]\leq C\cdot\nu_{n}^{1/\alpha}.$
Appendix D Further experimental details
D.1 Normal location illustrations
[…]
D.2 Palmer penguins example
The data containing $n=342$ measurements of the penguins’ bill lengths and depths in millimetres were imported from the palmerpenguins package in Python, and jointly centred to have zero-mean and scaled to have unit variance. To these data we fit a Gaussian kernel density estimator implemented in sklearn (Pedregosa et al., 2011), where the bandwidth was chosen with cross-validation across a logarithmic grid. We also fit for comparison a PrO posterior over the mean of an the isotropic bivariate Gaussian location model class, $\mathcal{M}_{\Theta}=\{\mathsf{N}(\theta,\sigma^{2}I):\,\theta\in\mathbb{R}^{2}\}$, where $\sigma^{2}=0.2$ is fixed, and which we fit with the squared MMD score. This same bandwidth was used for fitting the PrO posterior with the squared MMD as fitting the KDE. We estimated the PrO posterior under the squared MMD loss approximated with $m=n$ samples, $\lambda_{n}=10^{3}$, and the Wasserstein gradient flow was run with $p=32$ particles for $10,000$ iterations.
D.3 Regression with the MMD
Denote $d=\operatorname{dim}(\Theta)$. In Sections 25.2, and 5.3 we fit the linear models under the squared MMD loss. Denote the model class as $\{\mathsf{N}(y_{i};\,x_{i}^{\top}\theta,\sigma^{2}):\,\theta\in\mathbb{R}^{d}\}$, for $i=1,\ldots,n$ where $\sigma^{2}>0$ is treated as fixed and estimated from data: $\hat{\sigma}^{2}=\sum_{i=1}^{n}\frac{(y_{i}-\hat{y}_{i})}{n-d-1};\quad\hat{y}=X\hat{\beta};\quad\hat{\beta}=(X^{\top}X)^{-1}X^{\top}y.$ In the case of the Gibbs posterior, we use the efficient estimator proposed by Alquier and Gerber (2024, Equation 6): $S_{n}(P_{\theta})=\sum_{i=1}^{n}\ell^{\dagger}(\theta;\,X_{i},y_{i})$ where for all $\theta\in\Theta,\,x\in\mathbb{R}^{d},$ and $y\in\mathbb{R}$, $\ell^{\dagger}(\theta;\,X,y)=\mathbb{E}_{Y,Y^{\prime}\sim P_{\theta}(\cdot\mid x)}\{\kappa(Y,Y^{\prime})-2\kappa(Y,y)\}.$ In practice we estimate both of these expectations with $m$ Monte Carlo draws: $\ell^{\dagger}(\theta;\,X,y)\approx\frac{1}{m(m-1)}\sum_{i\neq j}\sum_{j=1}^{m}\kappa(Y_{i},Y^{\prime}_{j})-\frac{2}{m}\kappa(Y_{i},y),$ for $Y_{i},Y^{\prime}_{i}\sim P_{\theta}(\cdot\mid x).$ In the case of the PrO posterior, we instead use the tandem loss $L_{n}(\vartheta,\theta)=\sum_{i=1}^{n}\ell(\vartheta,\theta;\,X_{i},y_{i})$ where now $\ell(\vartheta,\theta;\,X,y)=\mathbb{E}_{\begin{subarray}{c}Y\sim P_{\vartheta}(\cdot\mid x)\\ Y^{\prime}\sim P_{\theta}(\cdot\mid x)\end{subarray}}\{\kappa(Y,Y^{\prime})-\kappa(Y,y)-\kappa(Y^{\prime},y)\}.$ And again, each of these expectations can be approximated by Monte Carlo: $\ell(\vartheta,\theta;\,X,y)\approx\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{j=1}^{m}\kappa(Y_{i},Y^{\prime}_{j})-\frac{1}{m}\kappa(Y_{i},y)-\frac{1}{m}\kappa(Y_{i}^{\prime},y),$ for $Y_{i}\sim P_{\vartheta}(\cdot\mid x)$ and $Y^{\prime}_{i}\sim P_{\theta}(\cdot\mid x)$. In both cases, we use a Gaussian kernel where the length scale is chosen according to the median heuristic (Garreau et al., 2018).
D.4 Linear regression example
[…]
D.4.1 Sensitivity of the Wasserstein gradient flow
In Figure 14 we show the sensitivity of the kernel gradient discrepancy (a diagnostic of the quality of the posterior samples) and two predictive scores to the choice of the number of particles and step size in a linear regression example. We sample $n=500$ data from the CR setting described in Section 2, and then fit the PrO posterior with $p=\{2^{1},2^{2},\ldots,2^{6}\}$ particles and $\mathsf{d}t$ between $0.1$ and $0.0001$. In each case, we set $\lambda_{n}=n^{1/2}$ and run the WGF for $10,000$ iterations and estimate the MMD with $m=32$ Monte Carlo samples. In Figure 14 we report the mean of the kernel gradient discrepancy (KGD), and predictive negative log likelihood (NLL) and MMD from $10$ fold cross-validation. All metrics are improved as $\mathsf{d}t$ is smaller, and as the number of particles grows. We observed here, and in other experiments, that as long as the number of particles is sufficiently large to capture multi-modal behaviour in the data, the step size becomes important to tune appropriately. And that the relationship relationship between predictive performance and step size is not always linear. This observation led us to choose $p$ of the order $\approx 2^{5}$, and optimise the step size in all experiments with the tuning-free FUSE schedule (Sharrock and Nemeth, 2025).
Refer to caption (a) Posterior diagnostic.
Refer to caption (b) Predictive metrics.
Figure 14: Synthetic linear regression example.
D.5 Binary classification example
We simulated $n=1,000$ in the manner described in Section 5.1, and fit all models with $\lambda_{n}=n^{1/2}$. The PrO posterior was run with $p=64$ particles for $10,000$ iterations. MALA was run for $4,000$ steps with $\mathsf{d}t=10^{-4}$.
D.6 River water flow example
We fit an $\operatorname{AR}(7)$ model with Student-$t$ prior over the intercept $\beta$, and Minnesota-style priors over the regression coefficients $\theta$:                                                                                                     $\Pi(\beta)$ ${}\sim{}$ $t_{3}(0,1^{2})$                                                                                                                                                                                                          $\Pi(\theta_{j})$ ${}\sim{}$ $\mathsf{N}\left(0,\,\frac{1}{j^{2}}\right),\,\quad j=1,\ldots 7.$                                                                                                      The PrO posterior was run with $p=50$ particles for $10,000$ iterations. MALA was run for $4,000$ steps after $46,000$ warmup iterations, with $\mathsf{d}t=10^{-6}$. The MMD was estimated with $m=32$ Monte Carlo samples, and both models were fit with $\lambda_{n}=n^{1/2}$.
D.7 Time series cross-validation
Following Bürkner et al. (2020), suppose we have $T$ observations from a time series, $y_{i},\ldots,y_{T}$, and let $L$ be the minimum number of observations from the series that we will require before making predictions for future data. Then we want to evaluate the predictive performance of $M$ one-step-ahead predictions made by the model, starting with its performance on the $L+1$-th datum: $\int P_{\theta}(y_{L+1})\mathsf{d}Q_{n}(\theta\mid y_{1:L})$. For instance, the leave-future-out cross-validated one-step-ahead log score is given as $\sum_{i=L}^{T-1}\log\int P_{\theta}(y_{i+1}\mid y_{1:i})\mathsf{d}Q_{n}(\theta\mid y_{1:i}),$ and we compute the $\operatorname{MMD}^{2}$ and $\operatorname{RMSE}$ similarly: fitting the model on $i$ data and evaluating on the the $i+1$-th datum because incrementing the index $i$ by one until we have no more observations on which to evaluate. This procedure is elegantly visualised in Figure 1 of Bürkner et al. (2020).
D.8 Conditional auto-regression example
The PrO posterior under the MMD was run with $p=16$ particles for $2,000$ iterations. MALA was run for $4,000$ steps after $96,000$ warmup iterations, with $\mathsf{d}t=10^{-6}$. The MMD was estimated with $m=32$ Monte Carlo samples, and both models were fit with $\lambda_{n}=n^{1/2}$. In the case of the log score, the Gibbs posterior is fit with the same hyper-parameters as described above, while the PrO posterior is now fit with $p=128$ particles with $4,000$ iterations. Refer to caption Figure 15: The empirical data distribution of the median housing value by census tract.
Appendix E Computation
E.1 Examples: MMD and logarithmic score
Our above development recovers the algorithm of Shen et al. (2025) for the special case of squared MMD. Here, elementary calculations show that for $\mathsf{L}_{\operatorname{MMD}}(\vartheta,\theta;x)=\langle\mu(P_{\vartheta})-\mu(\updelta_{x}),\mu(P_{\theta})-\mu(\updelta_{x})\rangle_{\mathcal{H}}$ and $\nabla_{1}\mathsf{L}_{\operatorname{MMD}}(\vartheta,\theta;x)$ denoting its gradient with respect to $\vartheta$,                                                                                                     $\mathcal{W}(Q)[\vartheta]$ ${}={}$ $\frac{1}{n}\sum_{i=1}^{n}\int\nabla_{1}\mathsf{L}_{\operatorname{MMD}}(\vartheta,\theta;x_{i})\;\mathsf{d}Q(\theta),$                                                                                                      (34) so that the evolution equations take the form                                                                                                     $\mathsf{d}\vartheta_{t}^{(j)}=-\left\{\frac{1}{(p-1)}\sum_{\begin{subarray}{c}\ell=1\\ \ell\neq j\end{subarray}}^{p}\frac{\lambda_{n}}{n}\sum_{i=1}^{n}\nabla_{1}\mathsf{L}_{\operatorname{MMD}}(\vartheta_{t}^{(j)},\vartheta_{t}^{(\ell)};x_{i})-\nabla_{\vartheta}\log\mathsf{d}\Pi(\vartheta_{t}^{(j)})\right\}\mathsf{d}t+\sqrt{2}\mathsf{d}B_{t}^{(j)}.$                                                                                                      (35) For the logarithmic scoring rule, one instead finds that                                                                                                     $\mathcal{W}(q)[\vartheta]$ ${}={}$ $\frac{1}{n}\sum_{i=1}^{n}\frac{\nabla_{\vartheta}\,\mathsf{d}P_{\vartheta}(x_{i})}{\int\mathsf{d}P_{\theta}(x_{i})\mathsf{d}Q(\theta)},$                                                                                                      (36) so that the evolution equations are recovered as                                                                                                     $\mathsf{d}\vartheta_{t}^{(j)}=-\left\{\frac{\lambda_{n}}{n}\sum_{i=1}^{n}\frac{\nabla_{\vartheta}\,\mathsf{d}P_{\vartheta_{t}^{(j)}}(x_{i})}{\frac{1}{(p-1)}\sum_{\begin{subarray}{c}\ell=1\\ \ell\neq j\end{subarray}}^{p}\mathsf{d}P_{\vartheta_{t}^{(\ell)}}(x_{i})}-\nabla_{\vartheta}\log\mathsf{d}\Pi(\vartheta_{t}^{(j)})\right\}\mathsf{d}t+\sqrt{2}\mathsf{d}B_{t}^{(j)}.$                                                                                                      (37)
E.2 Practicalities and Implementation
To provide intuition about how the sampling algorithm described in this section differs from more conventional Bayesian computation procedures, Figure 12 compares it to the Metropolis adjusted Langevin algorithm (MALA; Roberts and Tweedie, 1996; Roberts and Rosenthal, 1998) on several normal location examples. While the $p=32$ particles which target the PrO posterior are initialised randomly from the prior, they soon repel each other, and lead to a multi-modal distribution. In contrast, the single particle evolved according to MALA bounces around a singular mode. While the interacting particle approach has the distinct advantage of being able to sample from any PrO posterior, it introduces several hyper-parameters that may require careful tuning in practice. The most important of these are the step sizes for the discretisation of the underlying SDE, the number of particles $p$, as well as the exact choice of $\lambda_{n}$. Here, the choice of step size and the number of particles $p$ constitute a direct trade-off between approximation quality and computational efficiency: choosing the step size too small or the number of particles too large will in principle increase accuracy, but do so at the cost of exploding computational cost and substantively slowed convergence. Determining what constituted an appropriate step size in particular, is non-trivial and an active area of research. Concurrently, Sharrock and Nemeth (2025) developed an adaptive step size schedule for sampling form mean-field Langevin dynamics by adaptively tuning the step size, which they call FUSE. We set the step size adaptively according to FUSE in all experiments. While the theoretical result those authors showed do not extend to the PrO posterior, our empirical results give promise that similar results are attainable. We leave this as an area of future research. The relationship between the dimensionality of $\Theta$ and the number of particles required for consistent estimation has been investigated in the related family of Stein variational gradient descent algorithms by Banerjee et al. (2025, Theorems 1 and 4). While those authors found that the number of particles must scale polynomially or exponentially with the dimension of the problem, we observed that when the step size is chosen with FUSE that much fewer particles are required. While little is known about choosing these hyper-parameters in our setting, recent work by Chazal et al. (2025) may finally provide a principled way for diagnosing this (see, e.g., Appendix D.4.1 for a numerical demonstration). In contrast, we have clear ideas on how $\lambda_{n}$ should be set from the theory developed in Section 3: for the case covered in Theorem 1 this is means setting $\lambda_{n}\asymp n^{1/2}/\log(n)$.